### v1

# Height filtrations and base loci on flag bundles over a curve

Abstract.
Let $k$ be an algebraically closed field of characteristic zero. Let $C/k$ be a projective smooth curve with function field $K=k(C)$ and $G/k$ be a connected reductive group. Let $F$ be a principal $G$-bundle on $C$. Let $P\subseteq G$ be a parabolic subgroup and $\lambda:P\longrightarrow G$ be a strictly anti-dominant character. Then $F/P\longrightarrow C$ is a flag bundle and $\mathcal{L}_{\lambda}=F\times_{P}k_{\lambda}$ on $F/P$ is a relatively ample line bundle.
Let $h_{\mathcal{L}_{\lambda}}$ be the height function on $X=(F/P)_{K}$ induced by $\mathcal{L}_{\lambda}$. We compute its height filtration and successive minima: the height filtration is given by a Bruhat decomposition and successive minima are given by slopes.
An interesting application is that, in this case, Zhang’s inequality on successive minima can be enhanced into an equality. This is proved from the convex geometry point of view.
Let $f\in N^{1}(F/P)$ be the numerical class of a vertical fiber. We also compute the augmented base loci $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$ for any $t\in\mathbb{R}$ and it turns out that they are almost the same as the height filtration, which accords with the philosophy in Arakelov geometry that the positivity of the line bundle $\mathcal{L}_{\lambda}$ is equivalent to the positivity of its induced height function $h_{\mathcal{L}_{\lambda}}$. As a corollary, we compute the $k$-th movable cones of flag bundles over curves for all $k$.
1. Introduction
The main purpose of this paper is to give an explicit description of the height function and its related topics on a flag bundle over a curve. In this paper, all fields are of characteristic zero and all curves are projective smooth. By variety we mean an integral seperated scheme of finite type over a field.
1.1. Height filtration and successive minima
We start by giving a reminder on height functions in Arakelov theory and the relevant history. Let $K$ be either a number field or $K=k(C)$ where $C$ is a projective smooth curve over a field $k$. Let $X/K$ be a projective variety of dimension $d$ and $\overline{L}$ be an adelic line bundle on $X$. These data induce an Arakelov height function $h_{\overline{L}}$ on $X$ ([27], see also [25, §9] for a survey). A typical case is the geometric height, which is the one we concern in this article. Here we give a definition.
If $K=k(C)$, let $\mathcal{X}\longrightarrow C$ be a projective flat morphism with generic fiber $X\longrightarrow\operatorname{Spec}(K)$ and $\mathcal{L}$ be a line bundle on $\mathcal{X}$ with $\mathcal{L}_{K}\cong L$. The data $(\mathcal{X},\mathcal{L})$ define an adelic line bundle $\overline{L}$ and the height function $h_{\overline{L}}$ is given by
$h_{\overline{L}}:X(\overline{K})\longrightarrow\mathbb{Q},\;x\longmapsto\frac{\mathcal{L}\cdot\overline{\{x\}}}{{\mathrm{deg}}(x)}\;\;\text{where $\overline{\{x\}}$ is the closure of $x$ in $\mathcal{X}$}.$
We also denote this by $h_{\mathcal{L}}$ if there is no ambiguity. If $K$ is a number field, the height function can be defined similarly by arithmetic intersection theory.
Assume $L$ is ample. Let $Z_{t}(X,h_{\overline{L}})=\text{Zariski closure of}\;\big{\{}x\in X(\overline{K}):h_{\overline{L}}(x)<t\big{\}}$. Note that
  • $t\longmapsto Z_{t}(X,h_{\overline{L}})$ is an increasing filtration of Zariski closed subsets.
  • $Z_{t}(X,h_{\overline{L}})=X$ when $t\gg 0$ and $Z_{t}(X,h_{\overline{L}})=\emptyset$ when $t\ll 0$.
Since Zariski topology is Noetherian, the filtration $Z_{t}(X,h_{\overline{L}})$ gives a finite filtration $X_{0}=X\supsetneq X_{1}\supsetneq X_{2}\supsetneq\cdots\supsetneq X_{r}=\emptyset$, the height filtration. Its jumping points $\zeta_{i}(X,h_{\overline{L}})=\inf\big{\{}t:Z_{t}(X,h_{\overline{L}})=X_{i-1}\big{\}}$ are called the successive minima.
Note that our definition of successive minima are slightly different with Zhang [27]: Zhang considers only dimension jumps $e_{i}(X,h_{\overline{L}})=\inf\big{\{}t:\dim Z_{t}(X,h_{\overline{L}})\geq d-i+1\big{\}}$, $i=1,\dots,d+1$. In the following, $e_{i}(X,h_{\overline{L}})$ will be called Zhang’s successive minima to avoid ambiguity.
The first minimum $\zeta_{1}(X,h_{\overline{L}})=e_{1}(X,h_{\overline{L}})=\inf\big{\{}t:Z_{t}(X,h_{\overline{L}})=X\big{\}}$ is of particular interest, it is called the essential minimum and also denoted by $\zeta_{\text{ess}}(X,h_{\overline{L}})$. This invariant is the protagonist of the story of the Bogomolov conjecture, which in number field case says that a closed subvariety $X$ of an abelian variety $A$ has essential minimum zero with respect to the Néron-Tate height on $A$, if and only if it is the translation of an abelian subvariety by a torsion point, we refer the reader to [26, 24] for more details. The height filtration induced by the Néron-Tate height is $A\supsetneq\emptyset$ with the unique jumping point $\zeta=0$. This allows the equidistribution of small points to happen, and consequently implies the Bogomolov conjecture.
The height filtration (of toric height functions) on toric varieties is completely determined in [10]. An important example is on the projective space $\mathbb{P}^{n}_{\mathbb{Q}}$. Consider the height function induced by the universal bundle $\mathcal{O}_{\mathbb{P}^{n}_{\mathbb{Z}}}(1)$ equipped with the Fubini-Study metric $\|\cdot\|_{\text{FS}}$. The height filtration is given by successively deleting torus orbits, that is, $\mathbb{P}^{n}_{\mathbb{Q}}\supsetneq\bigcup_{i}\big{\{}x_{i}=0\big{\}}\supsetneq\bigcup_{i,j}\big{\{}x_{i}=0,x_{j}=0\big{\}}\supsetneq\cdots\supsetneq\bigcup_{i}\big{\{}x_{i}\neq 0\text{ and $x_{j}=0$ for all $j\neq i$}\big{\}}\supsetneq\emptyset$ with successive minima $\frac{1}{2}\log(k+1)$, $k=0,1,\dots,n$.
In this article, we consider another class of varieties with group actions: flag varieties. Let $k$ be an algebraic closed field, $G/k$ be a connected reductive group, $C/k$ be a curve, and $F$ be a principal $G$-bundle on $C$ with canonical reduction $F_{Q}$ to $Q$. This defines a linear functional ${\mathrm{deg}}(F_{Q})$ on $X(T)_{\mathbb{Q}}$. Let $P\subseteq G$ be a parabolic subgroup and $\lambda:P\longrightarrow G$ be a strictly anti-dominant character. Then $F/P\longrightarrow C$ is a flag bundle and $\mathcal{L}_{\lambda}=F\times_{P}k_{\lambda}$ on $F/P$ is a relatively ample line bundle. Let $X=(F/P)_{K}$ and $h_{\mathcal{L}_{\lambda}}$ be the height function on $X$ induced by $(F/P,\mathcal{L}_{\lambda})$. We compute its height filtration and successive minima.
Theorem A (Theorem 2.11).
The height filtration of $h_{\mathcal{L}_{\lambda}}$ on $X$ is given by successively deleting Schubert cells $C_{w}=(F_{Q}\times_{Q}QwP/P)_{K}$ for $w\in W_{Q}\backslash W/W_{P}$, i.e.
$Z_{t}=X\Bigg{\backslash}\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\geq t}C_{w}=\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle<t}C_{w}.$
In particular, successive minima are $\zeta_{w}=\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ and Zhang’s successive minima are $e_{i}=\min\big{\{}\zeta_{w}:\ell(w)=\dim G/P-i+1\big{\}}$ where $\ell(w)=\max\limits_{\sigma\in W_{Q}wW_{P}}\min\limits_{\tau\in\sigma W_{P}}\ell(\tau)$.
1.2. Zhang’s inequality
Let $X/K$ be a projective variety of dimension $d$, and $\overline{L}$ be a relative semipositive adelic line bundle with the underlying line bundle $L$ being ample. The height of $X$ is defined to be $h_{\overline{L}}(X)=\frac{\overline{L}^{d+1}}{(d+1)L^{n}}$. In [27], Zhang gives the following inequality:
$e_{1}\geq h_{\overline{L}}(X)\geq\frac{1}{d+1}\sum_{i=1}^{d+1}e_{i}.$
In the case of toric varieties, Zhang’s successive minima can be intepreted by convex geometry as shown in [9, 10]. Moreover, Zhang’s inequality can be proved directly [10, Theorem C] as an integral inequality of a convex function on a polytope. This can be partially generalized to arbitrary projective varieties by the theory of Okounkov bodies [19] and concave transforms [8, 2].
In this article, we carry on this spirit in the setting of flag bundles $F/P$. Given a strictly anti-dominant character $\lambda:P\longrightarrow\mathbb{G}_{m}$, representation theorists could associate to $\lambda$ a string polytope $\Delta_{\lambda}$, a weight polytope $\mathcal{P}_{\lambda}$ and a weight map $p:\Delta_{\lambda}\longrightarrow\mathcal{P}_{\lambda}$. Let $M_{\lambda}=G\times_{P}k_{\lambda}$ be the line bundle on $G/P$ induced by $\lambda$. Then integral points in $\Delta_{\lambda}$ parametrizes a weight basis of $H^{0}(G/P,M_{\lambda})$ and $p:\Delta_{\lambda}\longrightarrow\mathcal{P}_{\lambda}$ sends an integral point to the weight of the vector parametrized by it [18, §2.2].
Since the weight polytope is the convex hull of the Weyl orbit $W\lambda$, it follows immediately from Theorem A that successive minima of $h_{\mathcal{L}_{\lambda}}$ are the value of ${\mathrm{deg}}(F_{Q})$ at vertices of $\mathcal{P}_{\lambda}$. We will also prove
$h_{\mathcal{L}_{\lambda}}(X)=\frac{1}{\operatorname{Vol}(\Delta_{\lambda})}\int_{\Delta_{\lambda}}{\mathrm{deg}}(F_{Q})\circ p\;\mathrm{d}\mu.$
An interesting application is that, Zhang’s inequality in this case can be enhanced into an equality, in the sense that
Theorem B (Theorem 3.1).
Let $q:W/W_{P}\longrightarrow W_{Q}\backslash W/W_{P}$ be the canoniacl projection and for $w\in W_{Q}\backslash W/W_{P}$ define $m_{w}=\#q^{-1}(w)$. Then
$h_{\mathcal{L}_{\lambda}}(X)=\frac{1}{\#W/W_{P}}\sum_{w\in W_{Q}\backslash W/W_{P}}m_{w}\zeta_{w}.$
This follows from a simple computation of integration on the string polytopes: in fact for any linear functional $f:X(T)_{\mathbb{R}}\longrightarrow\mathbb{R}$, we have
$\frac{1}{\operatorname{Vol}(\Delta_{\lambda})}\int_{\Delta_{\lambda}}f\circ p\;\mathrm{d}\mu=\frac{1}{\#W/W_{P}}\sum_{w\in W/W_{P}}f(w\lambda).$
We remark that it is certainly not possible to enhance Zhang’s inequality into an equality in general. Consider the height function on $\mathbb{P}^{n}_{\mathbb{Q}}$ induced by the the Hermitian line bundle $\big{(}\mathcal{O}_{\mathbb{P}^{n}_{\mathbb{Z}}}(1),\|\cdot\|_{\text{FS}}\big{)}$. The variety height is $\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{i}\frac{1}{j}$ [7], which is a rational number while the successive minima are irrational numbers $\tfrac{1}{2}\log(k+1)$, $k=0,1,\dots,n$.
1.3. Base loci
It is a general philosophy in Arakelov geometry that the positivity of $\overline{L}$ are equivalent to the positivity of its induced height function $h_{\overline{L}}$. The study of essential minimum and pseudo-effectivity provides an evidence [10, 2, 22]. Note that if the underlying line bundle $L$ is big, then $\overline{L}$ is big $\Longleftrightarrow$ $\zeta_{\mathrm{ess}}(h_{\overline{L}})>0$.
On the other hand, the theory of augmented base loci $\mathrm{B}_{+}(\cdot)$ is well-developed [13, 14] in the study of positivity in algebraic geometry. Now back to the case $\mathcal{X}=F/P$, let $f$ be a line bundle given by a vertical fiber. Recall that it is a classical result that $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$ is the union of closed subvariety over which the restricted volume is 00 (see §4.1). In particular, if $\mathcal{Y}\subseteq\mathcal{X}$ is a closed subvariety such that $(\mathcal{L}_{\lambda}-tf)|_{\mathcal{Y}}$ is not big on $\mathcal{Y}$, then $\mathcal{Y}\subseteq\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$.
For $w\in W_{Q}\backslash W/W_{P}$ such that $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\leq t$, our calculation of $h_{\mathcal{L}_{\lambda}}$ gives that the essential minimum $\zeta_{\mathrm{ess}}(h_{(\mathcal{L}_{\lambda}-tf)|_{\mathcal{X}_{\omega}}})$ on the Schubert variety $\mathcal{X}_{w}=F_{Q}\times_{Q}\overline{QwP}/P$ is no bigger than 00, hence $(\mathcal{L}_{\lambda}-tf)|_{\mathcal{X}_{\omega}}$ is not big by the arithmetic bigness criterion mentioned above. We thus obtain that
$\bigcup_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\leq t}\mathcal{X}_{w}\subseteq\bigcup_{\mathcal{Y}\subseteq\mathcal{X},(\mathcal{L}_{\lambda}-tf)|_{\mathcal{Y}}\text{ not big}}\mathcal{Y}\subseteq\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$
where $\mathcal{Y}$ runs over all closed subvariety of $\mathcal{X}$. Moreover, we prove that both inclusions are actually equality:
Theorem C (Theorem 4.6).
The augmented base locus of $\mathcal{L}_{\lambda}-tf$ is given by
$\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)=\mathcal{X}\Bigg{\backslash}\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle>t}\mathcal{C}_{w}=\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\leq t}\mathcal{C}_{w}$
where $\mathcal{C}_{w}=F_{Q}\times_{Q}QwP/P$.
In particular, comparing this with Theorem A, we see that $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)_{K}=Z_{t}$ for all $t\in\mathbb{R}$, except for the critical points $t=\zeta_{w}$ at which $B_{+}(\mathcal{L}_{\lambda}-\zeta_{w}f)$ is upper continuous while $Z_{\zeta_{w}}$ is lower continuous.
1.4. Movable cones
Recall that for a projective variety $Y$ of dimension $d$, the k-th movable cones $\operatorname{Mov}^{k}(Y)$, $k=1,2,\dots,d$ are defined as
$\operatorname{Mov}^{k}(Y)=\big{\{}\alpha\in N^{1}(Y)_{\mathbb{R}}:\mathrm{B}_{+}(\alpha)\text{ has codimension $\geq k$}\big{\}}.$
$\operatorname{Mov}^{k}(Y)$ are open cones in $N^{1}(Y)_{\mathbb{R}}$ and we have obvious inclusions
$\operatorname{Mov}^{d}\subseteq\operatorname{Mov}^{d-1}\subseteq\cdots\subseteq\operatorname{Mov}^{1}.$
Note that $\operatorname{Mov}^{1}$ is the big cone and $\operatorname{Mov}^{d}$ is the ample cone.
Let $k$ be an algebraically closed field, $C/k$ be a curve and $E$ be a vector bundle on $C$. Note that $N^{1}(\mathbb{P}(E))_{\mathbb{R}}$ is two dimensional, generated by $f$ and $\xi=\mathcal{O}_{\mathbb{P}(E)}(1)$. The ample cone is given by extremal ray $f$ and $\xi-\mu_{\min(E)}f$, and the big cone is given by extremal ray $f$ and $\xi-\mu_{\max(E)}f$. The former is a theorem of Miyaoka [21] and the latter can be found in [15]. Fulger and Lehmann computed $\operatorname{Mov}^{2}(\mathbb{P}(E))$ in [16]. In [4], the big cone of the Grassmann bundle of $E$ is computed and in [6], the ample cone of flag bundles of $E$ is computed.
As a direct consequence of Theorem C, we compute the $k$-th movable cones of flag bundles $F/P$ over a curve $C$ for all $k$. Since any element in $N^{1}(F/P)_{\mathbb{R}}$ can be written as $\mathcal{L}_{\lambda}-tf$, we may define two functions $\langle\alpha^{\vee},\cdot\rangle$ and $\langle{\mathrm{deg}}(F_{Q}),w\cdot\rangle$ on $N^{1}(F/P)_{\mathbb{R}}$ as
  • $\langle\alpha^{\vee},\cdot\rangle$ sends $\mathcal{L}_{\lambda}-tf$ to $\langle\alpha^{\vee},\lambda\rangle$.
  • $\langle{\mathrm{deg}}(F_{Q}),w\cdot\rangle$ sends $\mathcal{L}_{\lambda}-tf$ to $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle-t$.
These functions are well-defined, and
Theorem D (Theorem 4.9).
The $k$-th movable cone $\operatorname{Mov}^{k}(F/P)$ is the cone defined by $\langle\alpha^{\vee},\cdot\rangle<0$ for any $\alpha\in\Delta\backslash\Delta_{P}$ and $\langle{\mathrm{deg}}(F_{Q}),w\cdot\rangle>0$ for any $w\in W/W_{P}$ with $\ell(w)\geq n-k+1$.
In particular, we find that $\mathcal{L}_{\lambda}$ is $k$-movable $\Longleftrightarrow$ the $k$-th minimum of Zhang of $h_{\mathcal{L}_{\lambda}}$ is positive.
1.5. Example: Grassmann bundles
We provide here the computation of essential minimum and big cone on Grassmann bundles. Let $k$ be an algebraically closed field and $C/k$ be a curve with function field $K=k(C)$. Let $E$ be a vector bundle of rank $n$ on $C$ with Harder-Narasimhan filtration $0\subsetneq E_{1}\subsetneq\cdots\subsetneq E_{r}$. Let $B$ be the subgroup of $\operatorname{GL}_{n}$ of all upper triangular matrices, and let $Q$ be the parabolic subgroup containing $B$ of type $(\operatorname{rank}(E_{1}),\operatorname{rank}(E_{2}),\dots,\operatorname{rank}(E_{r}))$. Here we say a parabolic subgroup containing $B$ is of type $(a_{1},\dots,a_{r})$ if it consists of upper block triguluar matrices of shape
$\left[\begin{array}[]{cccc}A_{1}&*&\cdots&*\\ &A_{2}&\cdots&*\\ &&\cdots&*\\ &&&A_{r}\end{array}\right],\quad A_{i}\in\operatorname{GL}_{a_{i}}.$
Let $F(E)$ be the frame bundle of $E$, which is a $\operatorname{GL}_{n}$-principal bundle. Then the canonical reduction of $F(E)$ is $F_{Q}$ to $Q$ where $F_{Q}$ is the $Q$-bundle of frames respecting the filtration $0\subsetneq E_{1}\subsetneq\cdots\subsetneq E_{r}$. Let $\mu_{i}$ ($i=1,\dots,n$) be the $i$-th number in the sequence
$\underbrace{\mu(E_{1}),\dots,\mu(E_{1})}_{\text{$\operatorname{rank}(E_{1})$ times}},\underbrace{\mu(E_{2}/E_{1}),\dots,\mu(E_{2}/E_{1})}_{\text{$\operatorname{rank}(E_{2}/E_{1})$ times}},\dots,\underbrace{\mu(E/E_{r-1}),\dots,\mu(E/E_{r-1})}_{\text{$\operatorname{rank}(E/E_{r-1})$ times}}.$
Note that $X(T)_{\mathbb{Q}}=\bigoplus_{i=1}^{n}\mathbb{Q}\lambda_{i}$ where $\lambda_{i}:T\longrightarrow\mathbb{G}_{m}$ is the character of taking the $i$-th entry, and $\langle{\mathrm{deg}}(F_{Q}),\lambda_{i}\rangle=\mu_{i}$.
Let $\operatorname{Gr}_{r}(E)=F(E)/P$ be the Grassmann bundle of $r$-dimension quotients, where $P$ is the parabolic subgroup containing $B$ of type $(n-r,r)$. Identify the Weyl group of $G$ with $S_{n}$ and of $P$ with $S_{n-r}\times S_{r}$ as usual. Then $W/W_{P}=S_{n}\big{/}(S_{n-r}\times S_{r})$ can be identified with the multi-index set $\big{\{}(i_{1},\dots,i_{r}):1\leq i_{1}<i_{2}<\cdots<i_{r}\leq n\big{\}}$ via
$\left[\begin{matrix}1&2&\cdots&n-r&n-r+1&\cdots&n\\ *&*&\cdots&*&i_{1}&\cdots&i_{r}\end{matrix}\right]\longleftrightarrow(i_{1},\dots,i_{r}).$
$N^{1}(\operatorname{Gr}_{r}(E))$ is two dimensional with basis $\mathcal{O}(1)$ (determinant bundle of the universal bundle on $\operatorname{Gr}_{r}(E)$) and $f$ (a vertical fiber). Let $\det_{2}$ be the character
$\det\nolimits_{2}:P\longrightarrow\mathbb{G}_{m},\quad\begin{bmatrix}A_{1}&*\\ 0&A_{2}\end{bmatrix}\longmapsto\det(A_{2}).$
One checks that $\mathcal{O}(1)=F(E)\times_{P}k_{\det_{2}}$, $\det_{2}=\lambda_{n-r+1}+\cdots+\lambda_{n}$ in $X(T)$ and $I=(i_{1},\dots,i_{r})\in W/W_{P}$ acts on $\det_{2}$ by $I\det_{2}=\lambda_{i_{1}}+\cdots+\lambda_{i_{r}}$.
Let $h_{\mathcal{O}(1)}$ be the height function on $\operatorname{Gr}_{r}(E_{K})$ induced by the model $(\operatorname{Gr}_{r}(E),\mathcal{O}(1))$. By Theorem A, the successive minima are $\langle{\mathrm{deg}}(F_{Q}),I\det_{2}\rangle=\sum_{i\in I}\mu_{i}$. (Theorem A says the successive minima are $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ for $w\in W_{Q}\backslash W/W_{P}$, but these numbers are really just $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ for $w\in W/W_{P}$.) In particular, the longest element $I_{0}=(1,2,\dots,r)$ gives the essential minimum
$\zeta_{\text{ess}}=\langle{\mathrm{deg}}(F_{Q}),I_{0}\det\nolimits_{2}\rangle=\mu_{1}+\cdots+\mu_{r}.$
The only element in $\Delta\backslash\Delta_{P}$ is $\alpha=\lambda_{n-r}-\lambda_{n-r+1}$ and the function $\langle\alpha^{\vee},\cdot\rangle$ on $N^{1}(\operatorname{Gr}_{r}(E))$ is given by $\mathcal{O}(1)\longmapsto\langle\alpha^{\vee},\det_{2}\rangle<0$ and $f\longmapsto 0$. Let $I_{0}=(1,2,\dots,r)\in W/W_{P}$ be the longest element. Then the function $\langle{\mathrm{deg}}(F_{Q}),I_{0}\cdot\rangle$ on $N^{1}(\operatorname{Gr}_{r}(E))$ is given by $\mathcal{O}(1)\longmapsto\mu_{1}+\cdots+\mu_{r}$ and $f\longmapsto 1$. By Theorem C, the big cone is given by $\langle\alpha^{\vee},\cdot\rangle<0$ and $\langle{\mathrm{deg}}(F_{Q}),I_{0}\cdot\rangle>0$, and one sees that it is the cone given by extremal rays $f$ and $\mathcal{O}(1)-(\mu_{1}+\dots+\mu_{r})f$.
1.6. Notations and conventions
In the following, let $k$ be an algebraically closed field of characteristic zero and $C/k$ be a projective smooth curve with function field $K=k(C)$. Let $G/k$ be a connected reductive group and denote the identity component of its center by $Z(G)$. Fix a Borel subgroup $B\subseteq G$ and a maximal torus $T\subseteq B$. Let $W$ be the Weyl group and $\Delta$ be the set of simple roots with respect to $(G,B,T)$. For any $\alpha\in\Delta$, we denote by $\alpha^{\vee}$ the corresponding simple coroot.
We shall consider only parabolic subgroup $P\subseteq G$ containing $B$. For any parabolic subgroup $P$, let $W_{P}\subseteq W$ be the Weyl group $W(L_{P})$ of the Levi factor $L_{P}\subseteq P$ and $\Delta_{P}\subseteq\Delta$ be the simple roots of $L_{P}$. For parabolic subgroups $P,Q\subseteq G$ and any double coset $w\in W_{Q}\backslash W/W_{P}$, set $\ell(w)=\max_{\sigma\in W_{Q}wW_{P}}\min_{\tau\in\sigma W_{P}}\ell(\tau)$, which equals to $\operatorname{dim}QwP/P$. Here by abuse of notations, $\ell(\tau)$ is the usual length of the Weyl element $\tau$.
For any linear algebraic group $\Gamma/k$, let $X(\Gamma)$ denote the character group $\mathrm{Hom}(\Gamma,\mathbb{G}_{m})$. For any $\mathbb{Z}$-algebra $A$, denote by $X(\Gamma)_{A}$ the tensor product $X(\Gamma)\otimes_{\mathbb{Z}}A$. For any linear functional $f$ on $X(\Gamma)$, we shall usually denote $f(\alpha)$ by $\langle f,\alpha\rangle$ for $\alpha\in X(\Gamma)$. For any $\alpha\in X(\Gamma)$, denote by $k_{\alpha}$ the one-dimensional algebraic representation on the vector space $k$ with $\Gamma$ acting by $\alpha$.
1.7. Organization of the paper
Let $F$ be a principal $G$-bundle on $C$ and $\lambda:P\longrightarrow\mathbb{G}_{m}$ be a strictly anti-dominant character. Then $F/P$ is a flag bundle over $C$ and $\mathcal{L}_{\lambda}=F\times_{P}k_{\lambda}$ on $F/P$ is a relatively ample line bundle. Let $X=(F/P)_{K}$ and $h_{\mathcal{L}_{\lambda}}$ be the height function on $X$ induced by $(F/P,\mathcal{L}_{\lambda})$.
In Chapter $2$, we compute the height filtration and successive minima of $h_{\mathcal{L}_{\lambda}}$. In Chapter $3$, we enhance Zhang’s inequality into an equality by convex geometry. In Chapter $4$, we compute the augmented base loci of $\mathcal{L}_{\lambda}-tf$ on $F/P$. As a corollary we get the movable cones $\operatorname{Mov}^{k}(F/P)$.
2. Height filtrations and successive minima
2.1. Vector bundles, slope stability and Harder-Narasimhan filtration
Let $E$ be a vector bundle on $C$. The degree of $E$ is ${\mathrm{deg}}(E)={\mathrm{deg}}(\det(E))$ and the slope of $E$ is $\mu(E)={\mathrm{deg}}(E)/\operatorname{rk}(E)$. It is called slope semistable if for every subbundle $F\subseteq E$, $\mu(F)\leq\mu(E)$. This is equivalent to $\mu(Q)\geq\mu(E)$ for every quotient bundle $Q$ of $E$.
In general, there exists uniquely a filtration $0=E_{0}\subseteq E_{1}\subseteq\cdots\subseteq E_{r}=E$ such that
  • $E_{i}/E_{i-1}$ is semistable;
  • $\mu(E_{i}/E_{i-1})>\mu(E_{i+1}/E_{i})$.
This filtration is called the Harder-Narasimhan filtration of $E$ and we denote $\mu_{i}(E)=\mu(E_{i}/E_{i-1})$. The first and the last term are of particular interests. The first one $\mu_{1}(E)=\mu(E_{1})$ is called maximal slope of $E$ and is also denoted by $\mu_{\max}(E)$. The last one $\mu_{r}(E)=\mu(E/E_{r-1})$ is called minimal slope of $E$ and is also denoted by $\mu_{\min}(E)$.
2.2. Principal bundles
For any linear algebraic group $\Gamma/k$, a principal $\Gamma$-bundle on $C$ is a variety $F$ equipped with a right action of $\Gamma$ and a $\Gamma$-equivariant smooth morphism $F\longrightarrow C$ such that the map
$F\times_{C}(C\times\Gamma)\longrightarrow F\times_{C}F,\quad(f,(x,g))\longmapsto(f,fg)$
is an isomorphism.
Attached to any principal $\Gamma$-bundle $F$, one has the degree map
${\mathrm{deg}}(F):X(\Gamma)\longrightarrow\mathbb{Z},\quad\lambda\longmapsto\langle{\mathrm{deg}}(F),\lambda\rangle={\mathrm{deg}}(F\times_{\Gamma}k_{\lambda}).$
Here ${\mathrm{deg}}(F\times_{\Gamma}k_{\lambda})$ is the degree of the line bundle $F\times_{\Gamma}k_{\lambda}$ on the curve $C$.
Let $H$ be a closed subgroup of a linear algebraic group $\Gamma/k$. A reduction of structure group of $F$ to $H$ is a pair $(F_{H},\varphi)$ where $F_{H}$ is a principal $H$-bundle and $\varphi:F_{H}\times_{H}\Gamma\cong F$ is an isomorphism. By the universal property of the quotient $F/H$, the assignment to any section $\sigma:C\longrightarrow F/H$ the reduction $\sigma^{*}F$ of $F$ to $H$ is a one-one correspondence between reductions of structure group of $F$ to $H$ and sections of $F/H\longrightarrow C$.
Let $F$ be a principal $G$-bundle on $C$ for the connected reductive group $G$. Note that for any parabolic subgroup $P\subseteq G$ with Levi subgroup $L_{P}$, the natural inclusion
$X(P)\longrightarrow X(L_{P})\longrightarrow X(Z(L_{P}))$
becomes an isomorphism after tensoring with $\mathbb{Q}$. Thus we have
$X(T)_{\mathbb{Q}}\longrightarrow X(Z(L_{P}))_{\mathbb{Q}}=X(P)_{\mathbb{Q}}$
and by taking duals, we get the so-called slope map $X(P)^{\vee}_{\mathbb{Q}}\longrightarrow X(T)^{\vee}_{\mathbb{Q}}$ introduced in [23, §2.1.3]. Let $F_{P}$ be a reduction of $F$ to $P$. By applying the slope map to ${\mathrm{deg}}(F_{P})$ we can define $\langle{\mathrm{deg}}(F_{P}),\lambda\rangle$ for any $\lambda\in X(T)$.
2.3. Semistability and canonical reduction
We mainly follow [5].
Definition 2.1 (semistable).
A principal $G$-bundle $F$ on $C$ is called semistable if for any parabolic subgroup $P$, any reduction $F_{P}$ of $F$ to $P$ and any dominant character $\lambda$ of $P$ which is trivial on $Z(G)$, we have $\langle{\mathrm{deg}}(F_{P}),\lambda\rangle\leq 0$.
Definition 2.2 (canonical reduction).
Let $F$ be a principal $G$-bundle on $C$. A reduction $F_{P}$ of $F$ to a parabolic subgroup $P$ is called canonical if the following two conditions hold:
  1. (1)
    The the principal $L_{P}$ bundle $F_{P}\times_{P}L_{P}$ is semistable.
  2. (2)
    For any non-trivial character $\lambda$ of $P$ which is non-negative linear combination of simple roots, $\langle{\mathrm{deg}}(F_{P}),\lambda\rangle>0$.
Behrend in [3] proved canonical reduction exists uniquely. When $G=\operatorname{GL}_{n}$, the above definition of semistability recovers Mumford’s slope stability of vector bundles and the above deinition of canonical reduction recovers the Harder-Narasimhan filtration of vector bundles.
2.4. Height function induced by $(F/P,\mathcal{L}_{\lambda}$)
Let $F$ be a principal $G$-bundle with canonical reduction $F_{Q}$ to some parabolic subgroup $Q\subseteq G$. Let $P\subseteq G$ be a parabolic subgroup. Set $\mathcal{X}=F/P$ and $X=\mathcal{X}_{K}$.
A character $\lambda:P\longrightarrow\mathbb{G}_{m}$ is called strictly anti-dominant if the natural pairing $\langle\alpha^{\vee},\lambda\rangle<0$ for any $\alpha\in\Delta\backslash\Delta_{P}$. Let $\lambda:P\longrightarrow\mathbb{G}$ be a stricly anti-dominant character. Then the line bundle $M_{\lambda}=G\times_{P}k_{\lambda}$ on $G/P$ is ample. Therefore $\mathcal{L}_{\lambda}=F\times_{G}M_{\lambda}$ is a relatively ample line bundle on $\mathcal{X}=F\times_{G}G/P$ and induces a height function
$\quad\quad\quad h_{\mathcal{L}_{\lambda}}:X(\overline{K})$ $\longrightarrow\mathbb{Q},\quad x\longmapsto\frac{\mathcal{L}_{\lambda}\cdot\overline{\{x\}}}{{\mathrm{deg}}(x)}$
where $\overline{\{x\}}$ is the Zariski closure of $x$ in $\mathcal{X}$.
For $x\in X(K)$, let $\sigma_{x}:C\longrightarrow\mathcal{X}$ be the section induced by $x:\operatorname{Spec}(K)\longrightarrow X$ via valuative criterion, and let $F_{P,x}=\sigma_{x}^{*}F$ be the corresponding reduction to $P$.
Lemma 2.3.
For $x\in X(K)$, $h_{\mathcal{L}_{\lambda}}(x)=\langle{\mathrm{deg}}(F_{P,x}),\lambda\rangle$.
Proof.
By definition, $h_{\mathcal{L}_{\lambda}}(x)$ is the degree of $\sigma_{x}^{*}(\mathcal{L}_{\lambda})$ and $\langle{\mathrm{deg}}(F_{P,x}),\lambda\rangle$ is the degree of $F_{P,x}\times_{P}k_{\lambda}$. We have equality of line bundles
$\sigma_{x}^{*}(\mathcal{L}_{\lambda})=\sigma_{x}^{*}(F\times_{P}k_{\lambda})=(\sigma_{x}^{*}F)\times_{P}k_{\lambda}=F_{P,x}\times_{P}k_{\lambda}$
and the lemma follows by taking degree. ∎
2.5. A height lower bound in Schubert cells
For $w\in W_{Q}\backslash W/W_{P}$, write $\mathcal{C}_{w}=F_{Q}\times_{Q}QwP/P$, $\mathcal{X}_{w}=F_{Q}\times_{Q}\overline{QwP}/P$, $C_{w}=\mathcal{C}_{w,K}$ and $X_{w}=\mathcal{X}_{w,K}$. We shall show that
Proposition 2.4.
For any $x\in C_{w}(\overline{K})$, $h_{\mathcal{L}_{\lambda}}(x)\geq\langle{\mathrm{deg}}F_{Q},w\lambda\rangle$.
Note that for any $w^{\prime}\in W_{Q}$ and $\lambda\in X(T)$, we have $w^{\prime}\lambda-\lambda\in\mathbb{Z}[\Delta_{Q}]$ consequently $\langle{\mathrm{deg}}(F_{Q}),w^{\prime}\lambda\rangle=\langle{\mathrm{deg}}(F_{Q}),\lambda\rangle$. Note also that for any $\lambda\in X(P)$ and $w\in W_{P}$, $w\lambda=\lambda$. So the number $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ is well-defined for any $\lambda\in X(P)$ and $w\in W_{Q}\backslash W/W_{P}$.
Definition 2.5 (relative position).
A reduction $F_{P}$ of $F$ to $P$ is called in relative position $w\in W_{Q}\backslash W/W_{P}$ with respect to $F_{Q}$ if the image of the natural map
$F_{Q}\times_{C}F_{P}\longrightarrow G_{C},\quad(a,b)\longmapsto a^{-1}b$
lies in $QwP_{C}$. Here we implicitly apply the injections
$F_{P}\hookrightarrow F_{P}\times_{P}G\cong F,\quad F_{Q}\hookrightarrow F_{Q}\times_{Q}G\cong F.$
Note that by the universal property of the quotient stacks $[G//Q\times P]$ and $[QwP//Q\times P]$, this definition coincides with the one in [23, §4.1].
Lemma 2.6.
When $x\in C_{w}(K)$, $F_{P,x}$ is of relative position $w$ with respect to $F_{Q}$.
Proof.
Note that the injection $F_{P,x}=\sigma_{x}^{*}F\longrightarrow F$ has image in $F_{Q}\times_{Q}QwP$ by the commutativity of the following diagram
Now the lemma follows since $F_{Q}\times_{C}F_{P,x}\longrightarrow C\times G$ factors as
$F_{Q}\times_{C}F_{P,x}\longrightarrow F_{Q}\times_{C}\Big{(}F_{Q}\times_{Q}QwP\Big{)}\longrightarrow$
$\Big{(}F_{Q}\times_{C}F_{Q}\Big{)}\times_{Q}QwP\longrightarrow(C\times Q)\times_{Q}QwP=C\times QwP.$
Proposition 2.7 ([23], Proposition $4.6$).
If $F_{P}$ is in relative position $w$ with respect to $F_{Q}$, then $\langle{\mathrm{deg}}(F_{P}),\lambda\rangle\geq\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ for any anti-dominant character $\lambda$.
Now we can prove Proposition 2.4.
Proof of Proposition 2.4.
Let $L/K$ be any finite field extension. Assume that $x\in C_{w}(L)$. Consider the following commutative diagram
[…]
where $C_{L}$ is the normalization of $C$ in $L$ and $F_{C_{L}}=F\times_{C}C_{L}$. Let $\mathcal{L}_{C_{L},\lambda}$ be the pullback of $\mathcal{L}_{\lambda}$ via the map $F_{C_{L}}/P\longrightarrow F/P$, which is also the line bundle associated to the character $\lambda$ of $P$ via the principal bundle $F_{C_{L}}$ on $C_{L}$. Then by Lemma 2.3,
$h_{\mathcal{L}_{\lambda}}(x)=\tfrac{1}{[L:K]}{\mathrm{deg}}(\sigma_{x}^{*}\mathcal{L}_{\lambda})=\tfrac{1}{[L:K]}{\mathrm{deg}}(\sigma_{\widetilde{x}}^{*}\mathcal{L}_{C_{L},\lambda})=\tfrac{1}{[L:K]}\langle{\mathrm{deg}}(F_{C_{L},P,\widetilde{x}}),\lambda\rangle$
where $F_{C_{L},P,\widetilde{x}}$ is the reduction of $F_{C_{L}}$ to $P$ corresponding to the rational point $\widetilde{x}\in X_{L}(L)$.
Note that the canonical reduction of $F_{C_{L}}$ is $F_{Q,C_{L}}=F_{Q}\times_{C}C_{L}$ and ${\mathrm{deg}}(F_{Q,C_{L}})=[L:K]{\mathrm{deg}}(F_{Q})$ in $X(T)_{\mathbb{Q}}^{\vee}$. By Lemma 2.6, $F_{Q,C_{L}}$ is of relative position $w$ with respect to $F_{C_{L},P,\widetilde{x}}$. We conclude by applying Proposition 2.7 to $F_{C_{L}}$. ∎
2.6. Height filtration and successive minima of $h_{\mathcal{L}_{\lambda}}$
In this subsection, we employ Ballaÿ’s theorem [2, Theorem 1.2] to compute essential minimum. We start by finding the maximal slope of $\pi_{*}\mathcal{L}_{\lambda}|_{\mathcal{X}_{w}}$.
Lemma 2.8 ([17], Part I, §5.18).
On $\mathcal{X}_{w}=F_{Q}\times_{Q}\overline{QwP}/P\subseteq\mathcal{X}$, we have $\pi_{*}\big{(}\mathcal{L}_{\lambda}|_{\mathcal{X}_{w}}\big{)}=F_{Q}\times_{Q}H^{0}(\overline{QwP}/P,M_{\lambda})$.
Schieder computed the Harder-Narasimhan filtration of $F_{Q}\times_{Q}H^{0}(\overline{QwP}/P,M_{\lambda})$ in [23, §5.1]. Let $V/k$ be a finite dimensional $Q$-representation with weight decomposition $V=\bigoplus_{\nu\in X(T)}V[\nu]$. The $Q$-filtration $V_{\bullet,{\mathrm{deg}}(F_{Q})}$ of $V$ is defined by
$V_{t,{\mathrm{deg}}(F_{Q})}=\bigoplus_{\langle{\mathrm{deg}}(F_{Q}),\nu\rangle\geq t}V[\nu].$
Proposition 2.9.
The Harder-Narasimhan filtration of $F_{Q}\times_{Q}H^{0}(\overline{QwP}/P,M_{\lambda})$ is $F_{Q}\times_{Q}H^{0}(\overline{QwP}/P,M_{\lambda})_{\bullet,{\mathrm{deg}}(F_{Q})}$. Moreover, the maximal slope is $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$.
Proof.
The first assertion is a slight generalization of [23, Proposition 5.1], whose original texts only consider $G$-representations restricted to $Q$, but the proof actually works for arbitrary $Q$-representations. For the second assertion, we only need to show the highest weights in $H^{0}(\overline{QwP}/P,M_{\lambda})$ belong to $w\lambda+\mathbb{Z}[\Delta_{Q}]$.
When $Q=B$, $H^{0}(\overline{BwP}/P,M_{\lambda})$ has a single highest weight $w\lambda$ by the theory of Demazure modules [17, Chapter 14]. For general $Q$, let $\pi:W/W_{P}\longrightarrow W_{Q}\backslash W/W_{P}$ be the canoniacl projection. Note that
$\overline{QwP}/P=\bigcup_{w^{\prime}\in\pi^{-1}(w)}\overline{Bw^{\prime}P}/P.$
Since the restriction map
$H^{0}(\overline{QwP}/P,M_{\lambda})\longrightarrow H^{0}(\overline{Bw^{\prime}P}/P,M_{\lambda})$
is surjective for each $w^{\prime}\in\pi^{-1}(w)$ and the diagonal map
$H^{0}(\overline{QwP}/P,M_{\lambda})\longrightarrow\prod_{w^{\prime}\in\pi^{-1}(w)}H^{0}(\overline{Bw^{\prime}P}/P,M_{\lambda})$
is injective, the highest weights in $H^{0}(\overline{QwP}/P,M_{\lambda})$ belong to $w\lambda+\mathbb{Z}[\Delta_{Q}]$. ∎
Corollary 2.10.
The essential minimum $\zeta_{1}(h_{\mathcal{L}_{\lambda}},X_{w})$ of $h_{\mathcal{L}_{\lambda}}$ on $X_{w}$ is $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$.
Proof.
By Ballaÿ’s theorem [2, Theorem 1.2],
$\quad\quad\quad\zeta_{1}(h_{\mathcal{L}_{\lambda}},X_{w})$ $=\displaystyle\lim_{n\rightarrow\infty}\frac{\mu_{\max}(\pi_{*}\mathcal{L}_{n\lambda}|_{\mathcal{X}_{w}})}{n}$
$=\lim_{n\rightarrow\infty}\frac{n\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle}{n}=\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle.$
Theorem 2.11.
The height filtration of $h_{\mathcal{L}_{\lambda}}$ on $X$ is given by successively deleting Schubert cells $C_{w}$ for $w\in W_{Q}\backslash W/W_{P}$, i.e.
$Z_{t}=X\Bigg{\backslash}\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\geq t}C_{w}=\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle<t}C_{w}.$
In particular, successive minima are $\zeta_{w}=\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ and Zhang’s successive minima are $e_{i}=\min\big{\{}\zeta_{w}:\ell(w)=d-i+1\big{\}}$.
Proof.
On $X_{w}$, $\zeta_{1}(h_{\mathcal{L}_{\lambda}},X_{w})=\zeta_{w}$ and for any $x\in C_{w}$, $h(x)\geq\zeta_{w}$. Thus $Z_{\zeta_{w}}(X_{w})\subseteq X_{w}\backslash C_{w}=\bigcup_{w^{\prime}<w}X_{w^{\prime}}$. For a closed subvariety $Y$ of $X$, let $Z_{t}(Y)$ be the Zariski closure in $Y$ of $\big{\{}y\in Y:h(y)<t\big{\}}$. We claim that $Z_{\zeta_{w}}(X_{w})=X_{w}\backslash C_{w}=\bigcup_{w^{\prime}<w}X_{w^{\prime}}$.
In fact, suppose conversely that $Z_{\zeta_{w}}(X_{w})\subsetneq X_{w}\backslash C_{w}$. Then there exists $X_{w^{\prime}}\subsetneq X_{w}$ such that $Z_{\zeta_{w}}(X_{w})\cap X_{w^{\prime}}\subsetneq X_{w^{\prime}}$. This forces $Z_{\zeta_{w}}(X_{w^{\prime}})\subseteq Z_{\zeta_{w}}(X_{w})\cap X_{w^{\prime}}\subsetneq X_{w^{\prime}}$. But on $X_{w^{\prime}}$, the essential minimum is $\zeta_{w^{\prime}}<\zeta_{w}$, so $Z_{\zeta_{w}}(X_{w^{\prime}})=X_{w^{\prime}}$. We get contradiction and we have thus proved that on $X_{w}$, the set $Z_{t}(X_{w})=X_{w}$ when $t>\zeta_{w}$ and $Z_{t}(X_{w})=X_{w}\backslash C_{w}$ when $t\leq\zeta_{w}$.
Let $w_{0}\in W_{Q}\backslash W/W_{P}$ be the longest element. On $X=X_{w_{0}}$, when $t>\zeta_{w_{0}}$, $Z_{t}=X_{w_{0}}$ and when $t\leq\zeta_{w_{0}}$, $Z_{t}=X_{w_{0}}\backslash C_{w_{0}}=\bigcup_{w^{\prime}}X_{w^{\prime}}$ with $w^{\prime}<w_{0}$ and $\ell(w^{\prime})=\ell(w_{0})-1$. Repeating this procedure on each smaller $X_{w^{\prime}}$, we can complete the proof by induction. ∎
3. Reinforcement of Zhang’s inequality into an equality
In this section, we shall establish the following Zhang-type equality on flag varieties over function fields: the height of the variety in fact equals to a weighted average of the successive minima.
Theorem 3.1.
Let $q:W/W_{P}\longrightarrow W_{Q}\backslash W/W_{P}$ be the canoniacl projection. For $w\in W_{Q}\backslash W/W_{P}$, define $m_{w}=\#q^{-1}(w)$. Then
$h_{\mathcal{L}_{\lambda}}(X)=\frac{1}{\#W/W_{P}}\sum_{w\in W_{Q}\backslash W/W_{P}}m_{w}\zeta_{w}.$
We shall establish this theorem by convex geometry. Let $E$ be a vector bundle on $C$ with HN filtration $0=E_{0}\subseteq E_{1}\subseteq\cdots\subseteq E_{r}=E$ with successive slopes $\mu_{i}=\mu(E_{i}/E_{i-1})$. Chen defines a measure $\nu_{E}$ on $\mathbb{R}$ by
$\nu_{E}=\frac{1}{\operatorname{rank}(E)}\sum_{i=1}^{r}\operatorname{rank}(E_{i}/E_{i-1})\delta_{\mu_{i}},$
where $\delta_{a}$ is the Dirac measure on $\mathbb{R}$ supported at $a\in\mathbb{R}$. For any $\varepsilon>0$ and any measure $\mu$, let $T_{\varepsilon}\mu$ be the unique measure such that $\int_{\mathbb{R}}f(x)\mathrm{d}T_{\varepsilon}\mu(x)=\int_{\mathbb{R}}f(\varepsilon x)\mathrm{d}\mu(x)$ for any compactly-supported continuous function $f:\mathbb{R}\longrightarrow\mathbb{R}$.
Theorem 3.2 (Chen; [11], Theorem 1.1).
Let $\pi:\mathcal{X}\longrightarrow C$ be a projective flat morphism and $\mathcal{L}$ be a line bundle on $\mathcal{X}$. Assume $\mathcal{L}$ is generically big. Then
  1. (1)
    $\big{\{}T_{\frac{1}{n}}\nu_{\pi_{*}\mathcal{L}^{n}}\big{\}}_{n}$ converges weakly. We denote the limit measure by $\nu^{\pi}_{\mathcal{L}}$.
  2. (2)
    $\operatorname{Vol}(\mathcal{L})=\dim\mathcal{X}\cdot\operatorname{Vol}(\mathcal{L}_{K})\cdot\int_{\mathbb{R}}x_{+}\;\mathrm{d}\nu^{\pi}_{\mathcal{L}}$, where $x_{+}=\max\{0,x\}$.
Let $w_{0}\in W$ be the longest element and fix a reduced decomposition $w_{0}=s_{1}\cdot s_{2}\cdots s_{{\ell(w_{0})}}$. Recall that $\lambda:P\longrightarrow\mathbb{G}_{m}$ is a strictly anti-dominant character and $N=\dim G/P$. Let $\Delta_{\lambda}=\Delta_{\underline{w}_{0},\lambda}\subseteq\mathbb{R}^{N}$ be the string polytope associated to $\lambda$ with respect to $\underline{w}_{0}=(s_{1},\cdots,s_{\ell(w_{0})})$. It is equipped with a weight map $p:\Delta_{\lambda}\longrightarrow\mathcal{P}_{\lambda}$ where $\mathcal{P}_{\lambda}\subseteq X(T)_{\mathbb{R}}$ is the weight polytope associated to $\lambda$ (see [1, §1] for details). The following properties will be useful:
  1. (1)
    The volume $\operatorname{Vol}(\Delta_{\lambda})$ of the convex polytope equals to the volume $\operatorname{Vol}(M_{\lambda})$ of the line bundle.
  2. (2)
    Integral points $\Delta_{\lambda}\cap\mathbb{Z}^{N}$ parametrize a basis of $H^{0}(G/P,M_{\lambda})$. In particular, #$\#\Delta_{\lambda}\cap\mathbb{Z}^{N}=\dim H^{0}(G/P,M_{\lambda})$.
  3. (3)
    Lattice points $\mathcal{P}_{\lambda}\cap X(T)$ parameterize weights occuring in $H^{0}(G/P,M_{\lambda})$ and $p$ sends an integral point to the underlying weight of the corresponding basis vector.
  4. (4)
    $\Delta_{\lambda_{1}}+\Delta_{\lambda_{2}}=\Delta_{\lambda_{1}+\lambda_{2}}$ and $p_{\lambda_{1}}+p_{\lambda_{2}}=p_{\lambda_{1}+\lambda_{2}}$. Here the sum is the Minkowski sum. In particular, the diagram
    ${\Delta_{\lambda}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p_{\lambda}}$$\scriptstyle{\cdot n}$${\Delta_{n\lambda}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p_{n\lambda}}$${\mathcal{P}_{\lambda}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\cdot n}$${\mathcal{P}_{n\lambda}}$
    is commutative for any $n\in\mathbb{N}$, where $\cdot n$ is the scalar multiplication by $n$.
Let $\mu$ be the Lebesgue measure on $\mathbb{R}^{N}$ normalized with respect to $\mathbb{Z}^{N}$.
Lemma 3.3.
The following equalities hold:
$\nu_{\pi_{*}\mathcal{L}^{n}_{\lambda}}=\frac{1}{\#(\Delta_{n\lambda}\cap\mathbb{Z}^{N})}\sum_{x\in\Delta_{n\lambda}\cap\mathbb{Z}^{N}}\delta_{{\mathrm{deg}}F_{Q}\circ p(x)},$
$T_{\frac{1}{n}}\nu_{\pi_{*}\mathcal{L}_{\lambda}^{n}}=\frac{1}{\#(\Delta_{\lambda}\cap\frac{1}{n}\mathbb{Z}^{N})}\sum_{x\in\Delta_{\lambda}\cap\frac{1}{n}\mathbb{Z}^{N}}\delta_{{\mathrm{deg}}F_{Q}\circ p(x)},$
$\nu^{\pi}_{\mathcal{L}_{\lambda}}=\frac{1}{\operatorname{Vol}(\Delta_{\lambda})}\cdot({\mathrm{deg}}F_{Q}\circ p)_{*}\mu.$
Proof.
Bearing Lemma 2.8, Proposition 2.9 and the aforementioned properties of string polytopes at hand, we can carry over the argument in [11, Proposition 3.5] to show the first equality. Appling $T_{1/n}$ and taking limits, we obtain the second and third equalities. ∎
Theorem 3.4.
The variety height $h_{\mathcal{L}_{\lambda}}(X)$ of $X$ is $\frac{1}{\operatorname{Vol}(\Delta_{\lambda})}\int_{\Delta_{\lambda}}{\mathrm{deg}}(F_{Q})\circ p\;\mathrm{d}\mu$.
Proof.
We may consider $\mathcal{L}_{\lambda}+tf$ for large $t$ so that $\mathcal{L}_{\lambda}+tf$ is ample, where $f$ is the line bundle given by a vertical fiber. The measure $\mu^{\pi}_{\mathcal{L}_{\lambda}+tf}$ is supported on $\mathbb{R}_{+}$ and is a right shift of $\mu^{\pi}_{\mathcal{L}_{\lambda}}$ by $t$, so Theorem 3.2 says
$\quad\quad\quad(\mathcal{L}_{\lambda}+tf)^{N+1}$ $=(N+1)\operatorname{Vol}(\mathcal{L}_{\lambda,K})\int_{\mathbb{R}}x\;\mathrm{d}\nu^{\pi}_{\mathcal{L}_{\lambda}+tf}$
$=(N+1)\int_{\Delta_{\lambda}}{\mathrm{deg}}(F_{Q})\circ p+t\;\mathrm{d}\mu$
where the second equality is due to Lemma 3.3 and $\operatorname{Vol}(\mathcal{L}_{\lambda,K})=\operatorname{Vol}(\Delta_{\lambda})$. Thus
$h_{\mathcal{L}_{\lambda}+tf}(X)=\frac{1}{\operatorname{Vol}(\Delta_{\lambda})}\int_{\Delta_{\lambda}}{\mathrm{deg}}(F_{Q})\circ p\;\mathrm{d}\mu+t.$
Note that $h_{\mathcal{L}_{\lambda}+tf}(X)=h_{\mathcal{L}_{\lambda}}(X)+t$. So substracting $t$, we get
$h_{\mathcal{L}_{\lambda}}(X)=\frac{1}{\operatorname{Vol}(\Delta_{\lambda})}\int_{\Delta_{\lambda}}{\mathrm{deg}}(F_{Q})\circ p\;\mathrm{d}\mu.$
To prove Theorem 3.1, we need the following general result.
Proposition 3.5.
For any linear function $f:X(T)_{\mathbb{R}}\longrightarrow\mathbb{R}$,
$\frac{1}{\operatorname{Vol}(\Delta_{\lambda})}\int_{\Delta_{\lambda}}f\circ p\;\mathrm{d}\mu=\frac{1}{\#W/W_{P}}\sum_{w\in W/W_{P}}f(w\lambda).$
We start with two lemmata.
Lemma 3.6.
The pushforward measure $p_{*}\mu$ on $\mathcal{P}_{\lambda}$ is $W$-invariant.
Proof.
Note that the Lebesgue measure $\mu$ on $\Delta_{\lambda}$ is the limit of the Dirac measure $\delta_{\Delta_{\lambda}\cap\frac{1}{n}\mathbb{Z}^{N}}$, i.e. $\mu=\lim_{n}\delta_{\Delta_{\lambda}\cap\frac{1}{n}\mathbb{Z}^{N}}$. Consider the map $\Delta_{n\lambda}\longrightarrow\mathcal{P}_{n\lambda}$ and note that it sends integral points to its corresponding weights. Dividing by n, we see that
$p_{*}\delta_{\Delta_{\lambda}\cap\frac{1}{n}\mathbb{Z}^{N}}=\operatorname{mult}_{n}\delta_{\mathcal{P}_{\lambda}\cap\frac{1}{n}X(T)},$
where $\operatorname{mult}_{n}$ is the multiplicity function
$\operatorname{mult}_{n}:\mathcal{P}_{\lambda}\cap\tfrac{1}{n}X(T)\longrightarrow\mathbb{N},\quad\mu\longmapsto\dim H^{0}(G/P,L_{n\lambda})[n\mu].$
Since both $\operatorname{mult}_{n}$ and $\delta_{\mathcal{P}_{\lambda}\cap\tfrac{1}{n}X(T)}$ are $W$-invariant, we deduce that
$p_{*}\mu=\lim_{n}p_{*}\delta_{\Delta_{\lambda}\cap\frac{1}{n}\mathbb{Z}^{N}}$
is also $W$-invariant. ∎
Lemma 3.7.
The map $\mathcal{P}_{\lambda}\longrightarrow\mathcal{P}_{\lambda}$, $\mu\longmapsto\sum_{w\in W}w\mu$ is constant.
Proof.
We claim that for any $\mu_{1},\mu_{2}\in X(T)$ such that $\mu_{1}|_{Z(G)}=\mu_{2}|_{Z(G)}$,
$\sum_{w\in W}w\mu_{1}=\sum_{w\in W}w\mu_{2}.$
In fact, for any simply connected cover $G^{sc}\longrightarrow G$, the preimage $T^{sc}$ of $T$ is a maximal torus in $G^{sc}$ and $W=W(G,T)=W(G^{sc},T^{sc})$. Via the natural map $X(T)\longrightarrow X(T^{sc})$, we are reduced to the case $G$ is simply connected. In this case, $T^{W}=Z(G)$ and the desired equality follows directly.
$H^{0}(G/P,L_{\lambda})$ is a simple $G$-module so all weights $\mu$ occuring in $H^{0}(G/P,L_{\lambda})$ sharing the same restriction to $Z(G)$. So by the above claim, the lemma is true for lattice points $\mathcal{P}_{\lambda}\cap X(T)$. Note that the map $\mu\longmapsto\sum_{w\in W}w\mu$ is linear, we are done. ∎
Proof of Proposition 3.5.
Let $\mathcal{P}_{\lambda,+}$ be the intersection of $\mathcal{P}_{\lambda}$ with the anti-dominant cone $\big{\{}\langle\alpha^{\vee},\lambda\rangle<0:\alpha\in\Delta\backslash\Delta_{P}\big{\}}$. Note that the subgroup $W_{P}\subseteq W$ acts trivially on $\mathcal{P}_{\lambda}$. By Lemma 3.6, one has
$\int_{\Delta_{\lambda}}f\circ p\;\mathrm{d}\mu=\int_{\mathcal{P}_{\lambda}}f\;\mathrm{d}p_{*}\mu=\int_{\mathcal{P}_{\lambda,+}}\Big{(}\sum_{w\in W/W_{P}}f(wx)\Big{)}\;\mathrm{d}p_{*}\mu.$
By the linearity of $f$ and Lemma 3.7, it moreover equals to
$\int_{\mathcal{P}_{\lambda,+}}f\Big{(}\sum_{w\in W/W_{P}}wx\Big{)}\;\mathrm{d}p_{*}\mu=\Big{(}\sum_{w\in W/W_{P}}f(w\lambda)\Big{)}\int_{\mathcal{P}_{\lambda,+}}1\;\mathrm{d}p_{*}\mu.$
On the other hand,
$\operatorname{Vol}(\Delta_{\lambda})=\int_{\mathcal{P}_{\lambda}}1\;\mathrm{d}p_{*}\mu=\#{W/W_{P}}\int_{\mathcal{P}_{\lambda,+}}1\;\mathrm{d}p_{*}\mu.$
Taking ratios, we obtain the desired equality. ∎
Proof of Theorem 3.1.
Take $f={\mathrm{deg}}(F_{Q})$ in Proposition 3.5 and apply Theorem 3.4. ∎
Remark 3.8.
The fact that the variety height and successive minima are given by
$\displaystyle h_{\mathcal{L}_{\lambda}}(X)=\frac{1}{\operatorname{Vol(\Delta_{\lambda})}}\int_{\mathcal{P}_{\lambda}}{\mathrm{deg}}(F_{Q})\;\mathrm{d}p_{*}\mu,\quad\zeta_{w}=\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$
shows that the Arakelov geometry of $(F/P,\mathcal{L}_{\lambda})$ depends only on the function ${\mathrm{deg}}(F_{Q})$ on $X(T)_{\mathbb{Q}}$, which might be surprising at the first glance.
This is in fact because the Arakelov geometry of $(F/P,\mathcal{L}_{\lambda})$ is determined by the Harder-Narasimhan filtration of $\pi_{*}\mathcal{L}_{n\lambda}=F_{Q}\times_{Q}H^{0}(G/P,M_{n\lambda})$ (Lemma 2.8) for all $n\geq 1$, and the unipotent part of $Q$ acts trivially [23, Lemma 5.1].
Example 3.9.
Let $G=\operatorname{GL}_{4}$, $E$ be a vector bundle on $C$ of rank $4$ and consider the $\operatorname{Gr}(2,4)$-bundle $\operatorname{Gr}_{2}(E)$. In this case the string polytopes (with respect to a certain reduced decomposition) is just the Gelfand-Zetlin polytope [18, Remark 2.4]. Identify $X(T)_{\mathbb{Q}}$ with $\bigoplus_{i}\mathbb{Q}\lambda_{i}$ as before. The line bundle $\mathcal{O}(1)$ correspond to the character $\lambda=(0,0,1,1)$.
The Gelfand-Zetlin polytope $\operatorname{GZ}_{\lambda}$ is defined by $0\leq x_{2}\leq x_{1}\leq x_{3}\leq 1$ and $x_{2}\leq x_{4}\leq x_{3}$ in $\mathbb{R}^{4}$, with vertices $(0,0,0,0)$, $(0,0,1,0)$, $(0,0,1,1)$, $(1,0,1,0)$, $(1,0,1,1)$ and $(1,1,1,1)$. The weight polytope $\mathcal{P}_{\lambda}$ is the convex hull of $(0,0,1,1)$, $(0,1,0,1)$, $(0,1,1,0)$, $(1,1,0,0)$, $(1,0,1,0)$ and $(1,0,0,1)$, the weight map $p:\operatorname{GZ}_{\lambda}\longrightarrow\mathcal{P}_{\lambda}$ is
$(x_{1},x_{2},x_{3},x_{4})\longmapsto(x_{1},-x_{1}+x_{2}+x_{3},-x_{2}-x_{3}+x_{4},-x_{4})+(0,0,1,1)$
and ${\mathrm{deg}}(F_{Q})$ is the function on $X(T)_{\mathbb{Q}}$ given by $\lambda_{i}\longmapsto\mu_{i}$ as in §1.5.
A direct computation shows that
$\operatorname{Vol}(\operatorname{GZ}_{\lambda})=\frac{1}{12},\quad\int_{\operatorname{GZ}_{\lambda}}{\mathrm{deg}}(F_{Q})\circ p\;\mathrm{d}\mu=\frac{1}{24}(\mu_{1}+\mu_{2}+\mu_{3}+\mu_{4}).$
Thus $h(X)=\frac{1}{2}(\mu_{1}+\mu_{2}+\mu_{3}+\mu_{4})$.
On the other hand, $W/W_{P}=S_{4}\big{/}(S_{2}\times S_{2})$, #$\#W/W_{P}=6$ and
$\sum_{w\in W/W_{P}}\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle=\sum_{1\leq i<j\leq 4}\mu_{i}+\mu_{j}.$
We see that $\frac{1}{\#W/W_{P}}\sum_{w\in W/W_{P}}\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ is also $\frac{1}{2}(\mu_{1}+\mu_{2}+\mu_{3}+\mu_{4})$.
4. Base loci and movable cones
4.1. Preliminaries on algebraic geometry
We briefly recall some general theory of base loci. Let $\mathbf{k}$ be a field. In our case, it would be either the base field $k$ or the function field $K=K(C)$. Let $Y\rightarrow\operatorname{Spec}\mathbf{k}$ be a normal projective variety over $\mathbf{k}$, and $L$ be a line bundle over $X$. Let $V$ be a $\mathbf{k}$-subspace of $H^{0}(Y,L)$. Consider the evaluation map
$H^{0}(Y,L)\otimes_{\mathbf{k}}L^{\vee}\rightarrow\mathcal{O}_{Y}.$
The base locus $\mathrm{Bs}(V)$ of $V$ is the reduced scheme associated to the closed subscheme defined by the ideal sheaf given by the image of $\pi^{*}V\otimes{L^{\vee}}\rightarrow\mathcal{O}_{Y}$.
Definition 4.1.
The stable base locus of $L$ is defined as $\mathrm{B}(L)=\bigcap_{n\geq 1}\mathrm{Bs}(H^{0}(Y,nL))$. Take any ample line bundle $A$. The augmented base locus of $L$ is defined as $\mathrm{B}_{+}(L)=\bigcap_{n\geq 1}\mathrm{B}(nL-A)$, which is independent of the choice of $A$.
The following characterization of augmented base loci in terms of restricted volumes is well known, which we include here for the reader’s convenience.
Theorem 4.2 ([14], Theorem C).
Let $Z$ be a closed subvariety of $Y$. The restricted volume is defined as
$\operatorname{Vol}(Y|Z,L)=\limsup_{n\rightarrow+\infty}\frac{\dim_{\mathbf{k}}\mathrm{Im}(H^{0}(Y,nL)\rightarrow H^{0}(Z,nL|_{Z}))}{n^{\dim Z}/(\dim Z)!}$
where Im()normal-$\mathrm{Im}(\cdot)$ denotes the image. Then $\mathrm{B}_{+}(L)=\bigcup_{\operatorname{Vol}(Y|Z,L)=0}Z$.
Let $\mathrm{Pic}(Y)$ be the picard group of $Y$. Since $\mathrm{B}(L)=\mathrm{B}(mL)$ for any $m>1$, we can extend the definitions of $\mathrm{B}(\cdot)$ and $\mathrm{B}_{+}(\cdot)$ to $\mathrm{Pic}(Y)_{\mathbb{Q}}:=\mathrm{Pic}(Y)\otimes\mathbb{Q}$. Moreover, let $N^{1}(Y):=\mathrm{Div}(X)/\equiv_{num}$ be the numerical group of divisors. It is proved in [13] that the augmented base loci are well defined on $N^{1}(Y)_{\mathbb{R}}:=N^{1}(Y)\otimes_{\mathbb{Z}}\mathbb{R}$.
Let $\mathcal{Y}$ be a projective variety with a surjective morphism $\mathcal{Y}\rightarrow C$, and $\mathcal{L}$ be a line bundle over $\mathcal{Y}$. For any scheme-theoretical point $p\in C$, we denote by $\kappa(p)$ the residue field of $p$. Let $\mathcal{Z}$ be a subvariety of $\mathcal{Y}$. The fiber $\mathcal{Z}\times_{C}\mathrm{Spec}(\kappa(p))$ of $\mathcal{Z}$ over $p$, is denoted as $\mathcal{Z}_{p}$. We also denote by $\mathcal{L}_{p}$ the restriction $\mathcal{L}|_{\mathcal{Y}_{p}}$ of $\mathcal{L}$ to $\mathcal{Y}_{p}$.
We have the following lemma, which will be useful for our computation of augmented base loci in the next subsection.
Lemma 4.3.
The restriction of $\mathrm{B}(\mathcal{L})$ to $\mathcal{Y}_{p}$ can be computed by
$B(\mathcal{L})_{p}=\bigcap\nolimits_{n\geq 1}\mathrm{Bs}(\mathrm{Im}(H^{0}(\mathcal{X},n\mathcal{L})\otimes_{k}\kappa(p)\rightarrow H^{0}(\mathcal{Y}_{p},n\mathcal{L}_{p}))).$
4.2. Augmented base loci on the flag bundle $\mathcal{X}$
Let $\lambda\in X(P)$ be a strictly anti-dominant character. We fix a closed point $p_{0}\in C$, and set $f=\pi^{*}(\mathcal{O}_{C}(p_{0}))$. In this subsection, we are going to compute the augmented base locus $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$. Remind that we have the decomposition
$H^{0}(G/P,M_{\lambda})=\bigoplus_{\nu\in X(T)}H^{0}(G/P,M_{\lambda})[\nu]$
and the filtration
$H^{0}(G/P,M_{\lambda})_{t,{\mathrm{deg}}(F_{Q})}=\bigoplus_{\langle{\mathrm{deg}}(F_{Q}),\nu\rangle\geq t}H^{0}(G/P,M_{\lambda})[\nu].$
We start with a lemma describing the vanishing behaviour of sections in $H^{0}(G/P,M_{\lambda})[\nu]$.
Lemma 4.4.
Take $w\in W/W_{P}$. Then
  1. (1)
    $s\neq 0\in H^{0}(G/P,M_{\lambda})[w\lambda]$Longrightarrow $s(x)\neq 0$ for any $x\in BwP/P$.
  2. (2)
    $s\in H^{0}(G/P,M_{\lambda})[\nu]$ for $\nu\not\leq w\lambda$Longrightarrow $s(x)=0$ for any $x\in\overline{BwP}/P$.
Proof.
Consider the restriction map
$\mathrm{Res}:\ H^{0}(G/P,M_{\lambda})\longrightarrow H^{0}(\overline{BwP}/P,M_{\lambda}),$
which is $B$-equivariant and surjective. Since $H^{0}(G/P,M_{\lambda})[w\lambda]$ and $H^{0}(\overline{BwP}/P,M_{\lambda})[w\lambda]$ are both one-dimensional, one has $\mathrm{Res}(s)\neq 0$ for any $s\neq 0\in H^{0}(G/P,M_{\lambda})[w\lambda]$.
By the closedness of the vanishing locus of $\mathrm{Res}(s)$, there exists $x\in BwP/P$ such that $\mathrm{Res}(s)(x)\neq 0$. Note that by the highest weight theory, $B$ acts on $\mathrm{Res}(s)$ via $w\lambda$ and $BwP/P$ is a single $B$-orbit under translation. Consequently, $\mathrm{Res}(s)(x)\neq 0$ for any $x\in BwP/P$.
As for $(2)$, for any $\nu\not\leq w\lambda$, $H^{0}(\overline{BwP}/P,M_{\lambda})[\nu]=0$. Consequently $\mathrm{Res}(s)=0$ for any $s\in H^{0}(G/P,M_{\lambda})[\nu]$.
Lemma 4.5.
Let $p$ be a scheme-theoretical point of $C$. Consider the filtration $H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})_{\bullet,{\mathrm{deg}}(F_{Q})}$ of $H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})$ given by
$H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})_{t,{\mathrm{deg}}(F_{Q})}=F_{Q,\kappa(p)}\times_{Q}H^{0}(G/P,M_{\lambda})_{t,{\mathrm{deg}}(F_{Q})}$
for $t\in\mathbb{R}$, we have $\operatorname{Bs}\Big{(}H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})_{t,{\mathrm{deg}}(F_{Q})}\Big{)}=\mathcal{X}_{p}\backslash\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\geq t}\mathcal{C}_{w,p}$.
Proof.
Let $F_{T,\kappa(p)}$ be a reduction of $F_{Q,\kappa(p)}$ to $T$. For any $\nu\in X(T)$, consider the $\kappa(p)$-subspace $H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})[\nu]=F_{T,\kappa(p)}\times_{T}H^{0}(G/P,M_{\lambda})[\nu]$ of $H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})=F_{T,\kappa(p)}\times_{T}H^{0}(G/P,M_{\lambda})$. For $w^{\prime}\in W/W_{P}$, set $C_{w^{\prime}}=F_{T,\kappa(p)}\times_{T}Bw^{\prime}P/P$ and $X_{w^{\prime}}=F_{T,\kappa(p)}\times_{T}\overline{Bw^{\prime}P}/P$. Then $\mathcal{C}_{w,p}=\coprod_{\pi(w^{\prime})=w}C_{w^{\prime}}$ and $\mathcal{X}_{w,p}=\bigcup_{\pi(w^{\prime})=w}X_{w^{\prime}}$ where $\pi:W/W_{P}\longrightarrow W_{Q}\backslash W/W_{P}$ is the natural projection.
Since the principal $T$-bundle $F_{T,\kappa(p)}$ is trivial, we deduce that $s(x)\neq 0$ for any $x\in C_{w^{\prime}}$, any $w^{\prime}\in\pi^{-1}(w)$, and any $s\in H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})[w^{\prime}\lambda]$ from Lemma 4.4(1). Thus if $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\geq t$, $H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})[w^{\prime}\lambda]\subseteq H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})_{t,{\mathrm{deg}}(F_{Q})}$ for any $w^{\prime}\in\pi^{-1}(w)$ and consequently, $x\not\in\operatorname{Bs}(H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})_{\geq t})$ for any $x\in C_{w^{\prime}}$. This proves $\operatorname{Bs}(H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})_{t,{\mathrm{deg}}(F_{Q})})\subseteq\mathcal{X}_{p}\backslash\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\geq t}\mathcal{C}_{w,p}$.
On the other hand, we claim that if $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle<t$, then for any $s\in H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})_{t,{\mathrm{deg}}(F_{Q})}$, the restriction of $s$ to $X_{w}$ is zero. In fact, we may assume $s\in H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})[\nu]$ for some $\nu$ and $\langle{\mathrm{deg}}(F_{Q}),\nu\rangle\geq t$ implies that $\nu\not\leq w^{\prime}\lambda$ for any $w^{\prime}\in\pi^{-1}(w)$ (because if $\nu\leq w^{\prime}\lambda$ for some $w^{\prime}$, then $\langle{\mathrm{deg}}(F_{Q}),\nu\rangle\leq\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle<t$). By Lemma 4.4(2), the restriction of $s$ to $X_{w^{\prime}}$ is zero for each $w^{\prime}$. We win. ∎
Now we are ready to compute the augmented base locus of $\mathcal{L}_{\lambda}-tf$.
Theorem 4.6.
The augmented base locus of $\mathcal{L}_{\lambda}-tf$ is given by
$\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)=\mathcal{X}\Bigg{\backslash}\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle>t}\mathcal{C}_{w}=\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\leq t}\mathcal{C}_{w}.$
Proof.
We may assume that $t\in\mathbb{N}$ by taking a sufficiently large tensor power of $\mathcal{L}_{\lambda}-tf$. Let $t_{0}=\min\{\langle{\mathrm{deg}}(F_{Q}),\nu\rangle:H^{0}(G/P,M_{\lambda})[\nu]\not=0\text{ and }\langle{\mathrm{deg}}(F_{Q}),\nu\rangle>t\}$. Take $E=F_{Q}\times_{Q}H^{0}(G/P,M_{\lambda})_{t_{0},{\mathrm{deg}}(F_{Q})}\otimes\mathcal{O}_{C}(-tx)$, which is a subbundle of $\pi_{*}(\mathcal{L}_{\lambda}-tf).$ Moreover, $\mu_{\min}(E)=t_{0}-t>0$, i.e. $E$ is ample.
For each scheme-theoretical point $p\in C$, we have $E(p):=E\otimes_{C}\kappa(p)=F_{Q,\kappa(p)}\times H^{0}(G/P,M_{\lambda})_{t_{0},{\mathrm{deg}}(F_{Q})}.$ The ampleness of $E$ gives that
$H^{0}(C,\mathrm{Sym}^{m}E)\otimes_{k}\kappa(p)\longrightarrow\mathrm{Sym}^{m}(E(p))$
is surjective for every $m\gg 0$. Since $\pi_{*}(\mathcal{L})\otimes_{C}\kappa(p)=H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})$, we have the commutative diagram
which implies that $V:=\mathrm{Im}(\mathrm{Sym}^{m}E(p)\rightarrow H^{0}(\mathcal{X}_{p},m\mathcal{L}_{\lambda,p}))$ is contained in the image $\mathrm{Im}(H^{0}(C,\pi_{*}(m\mathcal{L}_{\lambda}-tmf))\otimes_{k}\kappa(p)\rightarrow H^{0}(\mathcal{X}_{p},m\mathcal{L}_{\lambda,p}))=\mathrm{Im}(H^{0}(\mathcal{X},m\mathcal{L}_{\lambda}-tmf)\otimes_{k}\kappa(p)\rightarrow H^{0}(\mathcal{X},m\mathcal{L}_{\lambda,p})).$ Hence
$\mathrm{B}(\mathcal{L}_{\lambda}-tf)_{p}\subseteq\mathrm{Bs}(V)=\mathrm{Bs}(E(p))=\mathcal{X}_{p}\backslash\coprod\nolimits_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle>t}\mathcal{C}_{w,p}$
by applying Lemma 4.5 to $H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})_{t_{0},{\mathrm{deg}}(F_{Q})}$. Consequently,
(4.1) $\mathrm{B}(\mathcal{L}_{\lambda}-tf)\subseteq\mathcal{X}\backslash\coprod\nolimits_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle>t}\mathcal{C}_{w}.$
Now we take an ample line bundle $\mathcal{A}$ over $\mathcal{X}$ of form $\mathcal{L}_{\lambda^{\prime}}-t^{\prime}f$ where $\lambda^{\prime}\in X(T)$ is a strictly anti-dominant character and $\langle{\mathrm{deg}}(F_{Q}),\nu\rangle-t^{\prime}>0$ for any $\nu\in X(T)$ such that $H^{0}(G/P,M_{\lambda^{\prime}})[\nu]\not=0.$
Take sufficiently large $n$ such that $\langle{\mathrm{deg}}(F_{Q}),w\lambda^{\prime}\rangle-t^{\prime}<n(t_{0}-t)$ for any $w\in W_{Q}\backslash W/W_{P}$. Thus $\langle{\mathrm{deg}}(F_{Q}),w(n\lambda-\lambda^{\prime})\rangle>nt-t^{\prime}$ if and only if $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\geq t_{0}.$ Similarly as in (4.1), we obtain that
$\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)\subset\mathrm{B}(n(\mathcal{L}_{\lambda}-tf)-\mathcal{L}_{\lambda^{\prime}}-t^{\prime}f)\subseteq\mathcal{X}\Bigg{\backslash}\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle>t}\mathcal{C}_{w}.$
On the other hand, for any $w\in W_{Q}\backslash W/W_{P}$ such that $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\leq t$, since
$\limsup_{m\rightarrow+\infty}\frac{\mu_{\max}(\pi_{*}(m(\mathcal{L}_{\lambda}-tf)|_{\mathcal{X}_{w}})}{n}=\zeta_{1}(\mathcal{L}_{\lambda}-tf|_{\mathcal{X}_{w}})=\zeta_{w}-t\leq 0,$
we have $\operatorname{Vol}(\mathcal{L}_{\lambda}-tf|_{\mathcal{X}_{w}})=0$ by a comparison result [12, Theorem 2.4] deduced from the Riemann-Roch theorem. Therefore $\operatorname{Vol}(\mathcal{X}|\mathcal{X}_{w},\mathcal{L}_{\lambda}-tf)=0$, which implies that $\mathcal{X}_{w}\subseteq B_{+}(\mathcal{L}_{\lambda}-tf)$ due to Theorem 4.2. In conclusion, $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)=\mathcal{X}\backslash\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle>t}\mathcal{C}_{w}.$
Remark 4.7.
Comparing Theorem 2.11 and Theorem 4.6, we find that $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)_{K}$ agrees with the height filtration $Z_{t}$ for every $t\not=\zeta_{w}$. At critical points, $\mathrm{B}_{+}$ is upper continuous while $Z_{t}$ is lower continuous.
4.3. Positive cones of flag bundles over curves
We first give a reminder on the definitions of various positive cones of a projective variety. Let $Y$ be a smooth projective variety over a field $\mathbf{k}$ (of characteristic zero). We recall the following:
  1. (1)
    The big cone $\mathrm{Big}(Y)\subseteq N^{1}(Y)_{\mathbb{R}}$ is the cone of numerical classes of big divisors. Its closure is called Pseudo-effective cone, denoted as $\mathrm{Psef}(Y)$.
  2. (2)
    The ample cone $\mathrm{Amp}(Y)\subseteq N^{1}(Y)_{\mathbb{R}}$ is the cone of numerical classes of ample divisors. Its closure is called nef cone, denoted as $\mathrm{Nef}(Y)$.
  3. (3)
    For each $k=1,\cdots,\dim Y$, the $k$-th movable cone is defined as $\operatorname{Mov}^{k}(Y)=\big{\{}D\in N^{1}(Y)_{\mathbb{R}}:\mathrm{codim}(\mathrm{B}_{+}(D))\geq k\big{\}}$ which is open in $N^{1}(Y)_{\mathbb{R}}$. They are open cones in $N^{1}(Y)_{\mathbb{R}}$ and we have obvious inclusions $\operatorname{Mov}^{d}\subseteq\cdots\subseteq\operatorname{Mov}^{1}$. Note that $\operatorname{Mov}^{1}$ is the big cone and $\operatorname{Mov}^{d}$ is the ample cone.
Now we come back to the case of flag varieties. Denote by
$c:X(P)\longrightarrow\mathrm{Pic}(G/P),\quad\lambda\longmapsto M_{\lambda}$
the character map. We have an exact sequence [20, Theorem 18.32]:
$0\longrightarrow X(G)\longrightarrow X(P)\longrightarrow\mathrm{Pic}(G/P)\longrightarrow\mathrm{Pic}(G)\longrightarrow\mathrm{Pic}(P)\longrightarrow 0$
Since $G$ and $P$ are connected linear algebraic groups over a characteristic zero field, both $\mathrm{Pic}(G)$ and $\operatorname{Pic}(P)$ are finite. Notice that the numerical equivalence and the linear equivalence coincide on $G/P$ since $G/P$ is smooth fano. Therefore the linear map $X(P)_{\mathbb{R}}\rightarrow N^{1}(G/P)_{\mathbb{R}}$ is surjective. It is well-known that $M_{\lambda}$ is ample $\Longleftrightarrow$ $M_{\lambda}$ is big $\Longleftrightarrow$ $\lambda$ is strictly anti-dominant, thus
$\mathrm{Amp}(G/P)=\mathrm{Big}(G/P)=\big{\{}c(\lambda):\langle\alpha^{\vee},\lambda\rangle<0\text{ for all $\alpha\in\Delta\backslash\Delta_{P}$}\big{\}}$
[…]
We now compute $\mathrm{Mov}^{i}(F/P)$.
Lemma 4.8.
Let $f\in N^{1}(F/P)$ be the class of a vertical fiber. Then every element in $N^{1}(F/P)$ can be writen as $\mathcal{L}_{\lambda}-tf$.
Proof.
Apply [20, Theorem 18.32] to the $G$-bundle $F\longrightarrow C$. We get an exact sequence
$X(P)\longrightarrow\operatorname{Pic}(F/P)\longrightarrow\operatorname{Pic}(F)\longrightarrow\operatorname{Pic}(P).$
Apply it to the $P$-bundle $F\longrightarrow F/P$. We get another exact sequence
$X(G)\longrightarrow\operatorname{Pic}(C)\longrightarrow\operatorname{Pic}(F)\longrightarrow\operatorname{Pic}(G).$
Note that after tensoring $\mathbb{R}$, $\operatorname{Pic}(P)_{\mathbb{R}}=\operatorname{Pic}(G)_{\mathbb{R}}=0$.
Consider the diagram
where $g$, $\phi$ and $\psi$ are just pullbacks. Note that $g$ and $\psi$ are surjective. Let $x\in\operatorname{Pic}(F/P)$ such that $g(x)=\psi(a)$ for some $a\in\operatorname{Pic}(C)_{\mathbb{R}}$. Then $g(x-\phi(a))=0$ and the first line is exact at the middle, so $x-\phi(a)=f(b)$. ∎
By the above description, we define two functions $\langle\alpha^{\vee},\cdot\rangle$ and $\langle{\mathrm{deg}}(F_{Q}),w\cdot\rangle$ on $N^{1}(F/P)_{\mathbb{R}}$ as
  1. (1)
    $\langle\alpha^{\vee},\cdot\rangle$ sends $\mathcal{L}_{\lambda}-tf$ to $\langle\alpha^{\vee},\lambda\rangle$.
  2. (2)
    $\langle{\mathrm{deg}}(F_{Q}),w\cdot\rangle$ sends $\mathcal{L}_{\lambda}-tf$ to $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle-t$.
To see they are well-defined, it suffices to deal with the case $\mathcal{L}_{\lambda}$ equals to a multiple of $f$. In this case, we see that $M_{\lambda}$ is a trivial line bundle on $G/P$ by restricting $\mathcal{L}_{\lambda}$ to a fiber. This implies $\lambda\in X(G)$ by the exact sequence $0\longrightarrow X(G)\longrightarrow X(P)\longrightarrow\operatorname{Pic}(G/P)$. Then $\mathcal{L}_{\lambda}=F\times_{P}k_{\lambda}=\pi^{*}(F\times_{G}k_{\lambda})$ where $\pi$ is the structure map $F/P\longrightarrow C$. In particular, $\mathcal{L}_{\lambda}=\langle{\mathrm{deg}}(F),\lambda\rangle f$ in $N^{1}(F/P)_{\mathbb{R}}$.
  1. (1)
    $\langle\alpha^{\vee},\cdot\rangle$ sends $f$ to zero and sends $\lambda\in X(G)$ also to zero. This proves $\langle\alpha^{\vee},\cdot\rangle$ is well-defined.
  2. (2)
    $\langle{\mathrm{deg}}(F_{Q}),w\cdot\rangle$ sends $\langle{\mathrm{deg}}(F),\lambda\rangle f$ to $\langle{\mathrm{deg}}(F),\lambda\rangle$ since $\lambda\in X(G)$ is invariant under Weyl group action and sends $\lambda\in X(G)$ also to $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle=\langle{\mathrm{deg}}(F_{Q}),\lambda\rangle=\langle{\mathrm{deg}}(F),\lambda\rangle$ since $F\times_{G}k_{\lambda}=F_{Q}\times_{Q}G\times_{G}k_{\lambda}=F_{Q}\times_{Q}k_{\lambda}$. This proves $\langle{\mathrm{deg}}(F_{Q}),w\cdot\rangle$ is well-defined.
Theorem 4.9.
The $k$-th movable cone $\operatorname{Mov}^{k}(F/P)$ is the cone defined by
  1. (1)
    $\langle\alpha^{\vee},\cdot\rangle<0$ for any $\alpha\in\Delta\backslash\Delta_{P}$, and
  2. (2)
    $\langle{\mathrm{deg}}(F_{Q}),w\cdot\rangle>0$ for any $w\in W/W_{P}$ with $\ell(w)\geq n-k+1$.
Proof.
The sufficiency follows immediately after Theorem 4.6. For the necessity, let $\mathcal{L}_{\lambda}-tf$ be a line bundle on $F/P$ such that $\mathrm{codim}(\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf))\geq k$. Then in particular $\mathcal{L}_{\lambda}-tf$ is big. This impies $M_{\lambda}$ is big on $G/P$, which is equivalent to $\langle\alpha^{\vee},\lambda\rangle<0$ for all $\alpha\in\Delta\backslash\Delta_{P}$. Then $\lambda$ is strictly anti-dominant and Theorem 4.6 applies, saying $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)=\mathcal{X}\backslash\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle-t>0}\mathcal{C}_{w}$. It has $\operatorname{codim}\geq k$Longleftrightarrow $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle-t>0$ for all $w$ with $\ell(w)\geq n-k+1$. ∎
Corollary 4.10.
$\mathcal{L}_{\lambda}-tf\in\operatorname{Mov}^{k}(F/P)$ normal-$\Longleftrightarrow$ $e_{k}(h_{\mathcal{L}_{\lambda}})>t$ where $e_{k}$ is the $k$-th minimum of Zhang.

### v2

@@
-Height filtrations and base loci on flag bundles over a curve
+Arakelov geometry on flag varieties over function fields and related topics
 Abstract.
 Let $k$ be an algebraically closed field of characteristic zero.
-Let $C/k$ be a projective smooth curve with function field $K=k(C)$ and $G/k$ be a connected reductive group.
-Let $F$ be a principal $G$-bundle on $C$.
-Let $P\subseteq G$ be a parabolic subgroup and $\lambda:P\longrightarrow G$ be a strictly anti-dominant character.
+Let $G$ be a connected reductive group over $k$, $P\subseteq G$ be a parabolic subgroup and $\lambda:P\longrightarrow G$ be a strictly anti-dominant character.
+Let $C$ be a projective smooth curve over $k$ with function field $K=k(C)$ and $F$ be a principal $G$-bundle on $C$.
 Then $F/P\longrightarrow C$ is a flag bundle and $\mathcal{L}_{\lambda}=F\times_{P}k_{\lambda}$ on $F/P$ is a relatively ample line bundle.
-Let $h_{\mathcal{L}_{\lambda}}$ be the height function on $X=(F/P)_{K}$ induced by $\mathcal{L}_{\lambda}$.
-We compute its height filtration and successive minima: the height filtration is given by a Bruhat decomposition and successive minima are given by slopes.
-An interesting application is that, in this case, Zhang’s inequality on successive minima can be enhanced into an equality. This is proved from the convex geometry point of view.
+We compute the height filtration, successive minima, and the Boucksom-Chen concave transform of the height function $h_{\mathcal{L}_{\lambda}}:X(\overline{K})\longrightarrow\mathbb{R}$ over the flag variety $X=(F/P)_{K}$.
+An interesting application is that the height of $X$ equals to a weighted average of successive minima, and one may view this as a refinement of Zhang’s inequality of successive minima.
 Let $f\in N^{1}(F/P)$ be the numerical class of a vertical fiber.
-We also compute the augmented base loci $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$ for any $t\in\mathbb{R}$ and it turns out that they are almost the same as the height filtration, which accords with the philosophy in Arakelov geometry that the positivity of the line bundle $\mathcal{L}_{\lambda}$ is equivalent to the positivity of its induced height function $h_{\mathcal{L}_{\lambda}}$.
+We compute the augmented base loci $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$ for any $t\in\mathbb{R}$, and it turns out that they are almost the same as the height filtration.
 As a corollary, we compute the $k$-th movable cones of flag bundles over curves for all $k$.
 1. Introduction
-The main purpose of this paper is to give an explicit description of the height function and its related topics on a flag bundle over a curve.
-In this paper, all fields are of characteristic zero and all curves are projective smooth.
-By variety we mean an integral seperated scheme of finite type over a field.
 1.1. Height filtration and successive minima
-We start by giving a reminder on height functions in Arakelov theory and the relevant history.
+We start by giving a reminder on height functions in Arakelov theory.
 Let $K$ be either a number field or $K=k(C)$ where $C$ is a projective smooth curve over a field $k$.
-Let $X/K$ be a projective variety of dimension $d$ and $\overline{L}$ be an adelic line bundle on $X$.
-These data induce an Arakelov height function $h_{\overline{L}}$ on $X$ ([27], see also [25, §9] for a survey).
+Let $X$ be a projective variety of dimension $d$ over $K$ and $\overline{L}$ be an adelic line bundle on $X$.
+These data induce an Arakelov height function $h_{\overline{L}}$ on $X$ (see [24, §9] for a survey).
 A typical case is the geometric height, which is the one we concern in this article.
-Here we give a definition.
-If $K=k(C)$, let $\mathcal{X}\longrightarrow C$ be a projective flat morphism with generic fiber $X\longrightarrow\operatorname{Spec}(K)$ and $\mathcal{L}$ be a line bundle on $\mathcal{X}$ with $\mathcal{L}_{K}\cong L$.
+Here we give the definition.
+If $K=k(C)$, consider a projective flat morphism $\mathcal{X}\longrightarrow C$ with the generic fiber $X\longrightarrow\operatorname{Spec}(K)$ and a line bundle $\mathcal{L}$ on $\mathcal{X}$ with $\mathcal{L}_{K}\simeq L$.
 The data $(\mathcal{X},\mathcal{L})$ define an adelic line bundle $\overline{L}$ and the height function $h_{\overline{L}}$ is given by
 $h_{\overline{L}}:X(\overline{K})\longrightarrow\mathbb{Q},\;x\longmapsto\frac{\mathcal{L}\cdot\overline{\{x\}}}{{\mathrm{deg}}(x)}\;\;\text{where $\overline{\{x\}}$ is the closure of $x$ in $\mathcal{X}$}.$
-We also denote this by $h_{\mathcal{L}}$ if there is no ambiguity. If $K$ is a number field, the height function can be defined similarly by arithmetic intersection theory.
-Assume $L$ is ample. Let $Z_{t}(X,h_{\overline{L}})=\text{Zariski closure of}\;\big{\{}x\in X(\overline{K}):h_{\overline{L}}(x)<t\big{\}}$. Note that
-  • $t\longmapsto Z_{t}(X,h_{\overline{L}})$ is an increasing filtration of Zariski closed subsets.
-  • $Z_{t}(X,h_{\overline{L}})=X$ when $t\gg 0$ and $Z_{t}(X,h_{\overline{L}})=\emptyset$ when $t\ll 0$.
-Since Zariski topology is Noetherian, the filtration $Z_{t}(X,h_{\overline{L}})$ gives a finite filtration $X_{0}=X\supsetneq X_{1}\supsetneq X_{2}\supsetneq\cdots\supsetneq X_{r}=\emptyset$, the height filtration.
-Its jumping points $\zeta_{i}(X,h_{\overline{L}})=\inf\big{\{}t:Z_{t}(X,h_{\overline{L}})=X_{i-1}\big{\}}$ are called the successive minima.
-Note that our definition of successive minima are slightly different with Zhang [27]: Zhang considers only dimension jumps $e_{i}(X,h_{\overline{L}})=\inf\big{\{}t:\dim Z_{t}(X,h_{\overline{L}})\geq d-i+1\big{\}}$, $i=1,\dots,d+1$.
-In the following, $e_{i}(X,h_{\overline{L}})$ will be called Zhang’s successive minima to avoid ambiguity.
-The first minimum $\zeta_{1}(X,h_{\overline{L}})=e_{1}(X,h_{\overline{L}})=\inf\big{\{}t:Z_{t}(X,h_{\overline{L}})=X\big{\}}$ is of particular interest, it is called the essential minimum and also denoted by $\zeta_{\text{ess}}(X,h_{\overline{L}})$.
-This invariant is the protagonist of the story of the Bogomolov conjecture, which in number field case says that a closed subvariety $X$ of an abelian variety $A$ has essential minimum zero with respect to the Néron-Tate height on $A$, if and only if it is the translation of an abelian subvariety by a torsion point, we refer the reader to [26, 24] for more details.
-The height filtration induced by the Néron-Tate height is $A\supsetneq\emptyset$ with the unique jumping point $\zeta=0$.
-This allows the equidistribution of small points to happen, and consequently implies the Bogomolov conjecture.
-The height filtration (of toric height functions) on toric varieties is completely determined in [10].
-An important example is on the projective space $\mathbb{P}^{n}_{\mathbb{Q}}$.
-Consider the height function induced by the universal bundle $\mathcal{O}_{\mathbb{P}^{n}_{\mathbb{Z}}}(1)$ equipped with the Fubini-Study metric $\|\cdot\|_{\text{FS}}$.
-The height filtration is given by successively deleting torus orbits, that is, $\mathbb{P}^{n}_{\mathbb{Q}}\supsetneq\bigcup_{i}\big{\{}x_{i}=0\big{\}}\supsetneq\bigcup_{i,j}\big{\{}x_{i}=0,x_{j}=0\big{\}}\supsetneq\cdots\supsetneq\bigcup_{i}\big{\{}x_{i}\neq 0\text{ and $x_{j}=0$ for all $j\neq i$}\big{\}}\supsetneq\emptyset$ with successive minima $\frac{1}{2}\log(k+1)$, $k=0,1,\dots,n$.
-In this article, we consider another class of varieties with group actions: flag varieties.
-Let $k$ be an algebraic closed field, $G/k$ be a connected reductive group, $C/k$ be a curve, and $F$ be a principal $G$-bundle on $C$ with canonical reduction $F_{Q}$ to $Q$.
-This defines a linear functional ${\mathrm{deg}}(F_{Q})$ on $X(T)_{\mathbb{Q}}$.
-Let $P\subseteq G$ be a parabolic subgroup and $\lambda:P\longrightarrow G$ be a strictly anti-dominant character.
-Then $F/P\longrightarrow C$ is a flag bundle and $\mathcal{L}_{\lambda}=F\times_{P}k_{\lambda}$ on $F/P$ is a relatively ample line bundle.
-Let $X=(F/P)_{K}$ and $h_{\mathcal{L}_{\lambda}}$ be the height function on $X$ induced by $(F/P,\mathcal{L}_{\lambda})$.
-We compute its height filtration and successive minima.
+We also denote this by $h_{\mathcal{L}}$ if there is no ambiguity. If $K$ is a number field, the height function can be defined similarly via arithmetic intersection theory.
+For any $t\in\mathbb{R}$, let $Z_{t}\subseteq X$ be the Zariski closure of the set $\big{\{}x\in X(\overline{K}):h_{\overline{L}}(x)<t\big{\}}$.
+Let’s call $\big{\{}Z_{t}:t\in\mathbb{R}\big{\}}$ the height filtration, and call its jumping points the successive minima.
+Note that our definition of successive minima are slightly different with Zhang [25].
+Zhang considers only dimension jumps $e_{i}=\inf\big{\{}t:\dim Z_{t}\geq d-i+1\big{\}}$, $i=1,\dots,d+1$.
+In this article, $e_{i}$ will be called the successive minima of Zhang to avoid ambiguity.
+Determining the height filtration is generally difficult and only two cases are known in the literature:
+  • The Néron-Tate height function on abelian varieties $A$ over $K$.
+Let $L$ be an ample symmetric line bundle over $A$ endowed with the unique adelic metric such that $[2]^{*}\overline{L}=2^{\dim A}\overline{L}$.
+The height filtration of the induced Néron-Tate height function $\widehat{h}_{L}$ is $A\supsetneq\emptyset$ and the jumping point is $\zeta=0$;
+  • The toric height function on toric varieties [8].
+They are given by successively deleting torus orbits.
+An important example is the projective space $\mathbb{P}^{n}_{\mathbb{Q}}$ and the height function induced by the universal bundle $\mathcal{O}_{\mathbb{P}^{n}_{\mathbb{Z}}}(1)$ equipped with the Fubini-Study metric $\|\cdot\|_{\text{FS}}$.
+The height filtration on $\mathbb{P}^{n}_{\mathbb{Q}}$ is given by
+    $\quad\quad\quad\mathbb{P}^{n}_{\mathbb{Q}}\supsetneq\bigcup_{i}\big{\{}x_{i}=0\big{\}}\supsetneq\bigcup_{i,j}\big{\{}x_{i}=0,x_{j}=0\big{\}}\supsetneq\cdots\supsetneq\emptyset$
+    with successive minima $\frac{1}{2}\log(k+1)$, $k=0,1,\dots,n$.
+Our first result (Theorem A below) provides a new case where height filtrations can be explicitly computed.
+Let $k$ be an algebraically closed field of characteristic zero and $C$ be a projective smooth curve over $k$ with function field $K=k(C)$.
+Let $G$ be a connected reductive group over $k$, $P\subseteq G$ be a parabolic subgroup and $\lambda:P\longrightarrow\mathbb{G}_{m}$ be a strictly antidominant character.
+Let $F$ be a principal $G$-bundle over $C$ and $\mathcal{X}=F/P$ with generic fiber $X=(F/P)_{K}$.
+Let $\mathcal{L}_{\lambda}=F\times_{P}k_{\lambda}$ and $h_{\mathcal{L}_{\lambda}}:X(\overline{K})\longrightarrow\mathbb{R}$ be the induced height function.
+${X=(F/P)_{K}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{X}=F/P\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Spek}(K)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C}$
+Let $F_{Q}$ be the canonical reduction of $F$ to $Q$, and ${\mathrm{deg}}(F_{Q})\in X(T)^{\vee}_{\mathbb{Q}}$ be the induced cocharacter.
+Let $W,W_{P}$ and $W_{Q}$ be the Weyl group of $G,L_{P}$ and $L_{Q}$.
+For $w\in W_{Q}\backslash W/W_{P}$, let $C_{w}=(F_{Q}\times_{Q}QwP/P)_{K}\subseteq X$ be the corresponding Schubert cell.
+The numbers $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ are well-defined.
 Theorem A (Theorem 2.1).
-The height filtration of $h_{\mathcal{L}_{\lambda}}$ on $X$ is given by successively deleting Schubert cells $C_{w}=(F_{Q}\times_{Q}QwP/P)_{K}$ for $w\in W_{Q}\backslash W/W_{P}$, i.e.
+For any $t\in\mathbb{R}$, let $Z_{t}\subseteq X$ be the Zariski closure of the set $\big{\{}x\in X(\overline{K}):h_{\mathcal{L}_{\lambda}}(x)<t\big{\}}$. Then
 $Z_{t}=X\Bigg{\backslash}\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\geq t}C_{w}=\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle<t}C_{w}.$
-In particular, successive minima are $\zeta_{w}=\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ and Zhang’s successive minima are $e_{i}=\min\big{\{}\zeta_{w}:\ell(w)=\dim G/P-i+1\big{\}}$ where $\ell(w)=\max\limits_{\sigma\in W_{Q}wW_{P}}\min\limits_{\tau\in\sigma W_{P}}\ell(\tau)$.
-1.2. Zhang’s inequality
-Let $X/K$ be a projective variety of dimension $d$, and $\overline{L}$ be a relative semipositive adelic line bundle with the underlying line bundle $L$ being ample.
-The height of $X$ is defined to be $h_{\overline{L}}(X)=\frac{\overline{L}^{d+1}}{(d+1)L^{n}}$.
-In [27], Zhang gives the following inequality:
-$e_{1}\geq h_{\overline{L}}(X)\geq\frac{1}{d+1}\sum_{i=1}^{d+1}e_{i}.$
-In the case of toric varieties, Zhang’s successive minima can be intepreted by convex geometry as shown in [9, 10].
-Moreover, Zhang’s inequality can be proved directly [10, Theorem C] as an integral inequality of a convex function on a polytope.
-This can be partially generalized to arbitrary projective varieties by the theory of Okounkov bodies [19] and concave transforms [8, 2].
-In this article, we carry on this spirit in the setting of flag bundles $F/P$.
-Given a strictly anti-dominant character $\lambda:P\longrightarrow\mathbb{G}_{m}$, representation theorists could associate to $\lambda$ a string polytope $\Delta_{\lambda}$, a weight polytope $\mathcal{P}_{\lambda}$ and a weight map $p:\Delta_{\lambda}\longrightarrow\mathcal{P}_{\lambda}$.
-Let $M_{\lambda}=G\times_{P}k_{\lambda}$ be the line bundle on $G/P$ induced by $\lambda$.
-Then integral points in $\Delta_{\lambda}$ parametrizes a weight basis of $H^{0}(G/P,M_{\lambda})$ and $p:\Delta_{\lambda}\longrightarrow\mathcal{P}_{\lambda}$ sends an integral point to the weight of the vector parametrized by it [18, §2.2].
-Since the weight polytope is the convex hull of the Weyl orbit $W\lambda$, it follows immediately from Theorem A that successive minima of $h_{\mathcal{L}_{\lambda}}$ are the value of ${\mathrm{deg}}(F_{Q})$ at vertices of $\mathcal{P}_{\lambda}$.
-We will also prove
-$h_{\mathcal{L}_{\lambda}}(X)=\frac{1}{\operatorname{Vol}(\Delta_{\lambda})}\int_{\Delta_{\lambda}}{\mathrm{deg}}(F_{Q})\circ p\;\mathrm{d}\mu.$
-An interesting application is that, Zhang’s inequality in this case can be enhanced into an equality, in the sense that
+In particular, the filtration $\big{\{}Z_{t}:t\in\mathbb{R}\big{\}}$ is finite (while in general height filtrations could be infinite) and is given by successively deleting Schubert cells $C_{w}$ with jumping points $\zeta_{w}=\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$, $w\in W_{Q}\backslash W/W_{P}$.
+1.2. String polytopes and the Boucksom-Chen concave transform
+Attached to a projective variety $X$ over $K$ and an adelic line bundle $\overline{L}$ on $X$ with $L$ ample and $\overline{L}$ semipositive, we have the Okounkov body $\Delta$ and the Boucksom-Chen concave transform $\varphi:\Delta\longrightarrow\mathbb{R}$ thanks to [17, 18, 6, 1].
+The maximum of $\varphi$ equals to the essential minimum of $h_{\overline{L}}$, the minimum of $\varphi$ equals to the absolute minimum of $h_{\overline{L}}$ and the integral mean of $\varphi$ equals to the height of $X$.
+In toric cases, they can be expressed as summation of local roof functions [7].
+But in general, the Boucksom-Chen concave transform is difficult to compute.
+Our second result (Theorem B) provides a new case where the Boucksom-Chen concave transform can be explicitly computed.
+To state the result, keep the notations in Theorem A.
+Let $w_{0}\in W$ be the longest element and fix a reduced decomposition $\underline{w_{0}}=s_{i_{1}}s_{i_{2}}\cdots s_{i_{N}}$, $N=\dim G/B$.
+Then we have the string valuation $\nu$ and the string polytope $\Delta_{\lambda}\subseteq\mathbb{R}^{N}$ [19, 3], which is a polytope of dimension $d=\dim G/P$.
+Let $\mathcal{P}_{\lambda}$ be the weight polytope and $p:\Delta_{\lambda}\longrightarrow\mathcal{P}_{\lambda}$ be the weight map.
+The Okounkov body of $(X,L)$ (with respect to a specific valuation) is the string polytope $\Delta_{\lambda}$.
 Theorem B (Theorem 3.1).
-Let $q:W/W_{P}\longrightarrow W_{Q}\backslash W/W_{P}$ be the canoniacl projection and for $w\in W_{Q}\backslash W/W_{P}$ define $m_{w}=\#q^{-1}(w)$. Then
-$h_{\mathcal{L}_{\lambda}}(X)=\frac{1}{\#W/W_{P}}\sum_{w\in W_{Q}\backslash W/W_{P}}m_{w}\zeta_{w}.$
-This follows from a simple computation of integration on the string polytopes: in fact for any linear functional $f:X(T)_{\mathbb{R}}\longrightarrow\mathbb{R}$, we have
-$\frac{1}{\operatorname{Vol}(\Delta_{\lambda})}\int_{\Delta_{\lambda}}f\circ p\;\mathrm{d}\mu=\frac{1}{\#W/W_{P}}\sum_{w\in W/W_{P}}f(w\lambda).$
-We remark that it is certainly not possible to enhance Zhang’s inequality into an equality in general.
-Consider the height function on $\mathbb{P}^{n}_{\mathbb{Q}}$ induced by the the Hermitian line bundle $\big{(}\mathcal{O}_{\mathbb{P}^{n}_{\mathbb{Z}}}(1),\|\cdot\|_{\text{FS}}\big{)}$.
-The variety height is $\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{i}\frac{1}{j}$ [7], which is a rational number while the successive minima are irrational numbers $\tfrac{1}{2}\log(k+1)$, $k=0,1,\dots,n$.
-1.3. Base loci
-It is a general philosophy in Arakelov geometry that the positivity of $\overline{L}$ are equivalent to the positivity of its induced height function $h_{\overline{L}}$.
-The study of essential minimum and pseudo-effectivity provides an evidence [10, 2, 22].
-Note that if the underlying line bundle $L$ is big, then $\overline{L}$ is big $\Longleftrightarrow$ $\zeta_{\mathrm{ess}}(h_{\overline{L}})>0$.
-On the other hand, the theory of augmented base loci $\mathrm{B}_{+}(\cdot)$ is well-developed [13, 14] in the study of positivity in algebraic geometry.
-Now back to the case $\mathcal{X}=F/P$, let $f$ be a line bundle given by a vertical fiber.
-Recall that it is a classical result that $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$ is the union of closed subvariety over which the restricted volume is 00 (see §4.1).
-In particular, if $\mathcal{Y}\subseteq\mathcal{X}$ is a closed subvariety such that $(\mathcal{L}_{\lambda}-tf)|_{\mathcal{Y}}$ is not big on $\mathcal{Y}$, then $\mathcal{Y}\subseteq\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$.
-For $w\in W_{Q}\backslash W/W_{P}$ such that $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\leq t$, our calculation of $h_{\mathcal{L}_{\lambda}}$ gives that the essential minimum $\zeta_{\mathrm{ess}}(h_{(\mathcal{L}_{\lambda}-tf)|_{\mathcal{X}_{\omega}}})$ on the Schubert variety $\mathcal{X}_{w}=F_{Q}\times_{Q}\overline{QwP}/P$ is no bigger than 00, hence $(\mathcal{L}_{\lambda}-tf)|_{\mathcal{X}_{\omega}}$ is not big by the arithmetic bigness criterion mentioned above.
-We thus obtain that
-$\bigcup_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\leq t}\mathcal{X}_{w}\subseteq\bigcup_{\mathcal{Y}\subseteq\mathcal{X},(\mathcal{L}_{\lambda}-tf)|_{\mathcal{Y}}\text{ not big}}\mathcal{Y}\subseteq\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$
-where $\mathcal{Y}$ runs over all closed subvariety of $\mathcal{X}$. Moreover, we prove that both inclusions are actually equality:
+The Boucksom-Chen concave transform of ${\mathcal{L}_{\lambda}}$ is $\Delta_{\lambda}\longrightarrow\mathbb{R},x\longmapsto\langle{\mathrm{deg}}(F_{Q}),p(x)\rangle$.
+Let $\mu$ be the Lebesgue measure on the real span of $\Delta_{\lambda}$, normalized with respect to $\mathbb{Z}^{N}$.
+As we mentioned above, the variety height is given by the integral mean of the Boucksom-Chen concave transform [6].
+Thus a direct corollary is
+$h_{\mathcal{L}_{\lambda}}(X)=\frac{\mathcal{L}_{\lambda}^{\dim\mathcal{X}}}{\dim\mathcal{X}\cdot L^{\dim X}}=\frac{1}{\operatorname{Vol}(\Delta_{\lambda})}\int_{\Delta_{\lambda}}\langle{\mathrm{deg}}(F_{Q}),p(x)\rangle\,\mathrm{d}\mu(x).$
+The integration can be computed via Weyl symmetry.
 Theorem C (Theorem 3.2).
-The augmented base locus of $\mathcal{L}_{\lambda}-tf$ is given by
-$\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)=\mathcal{X}\Bigg{\backslash}\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle>t}\mathcal{C}_{w}=\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\leq t}\mathcal{C}_{w}$
-where $\mathcal{C}_{w}=F_{Q}\times_{Q}QwP/P$.
-In particular, comparing this with Theorem A, we see that $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)_{K}=Z_{t}$ for all $t\in\mathbb{R}$, except for the critical points $t=\zeta_{w}$ at which $B_{+}(\mathcal{L}_{\lambda}-\zeta_{w}f)$ is upper continuous while $Z_{\zeta_{w}}$ is lower continuous.
-1.4. Movable cones
-Recall that for a projective variety $Y$ of dimension $d$, the k-th movable cones $\operatorname{Mov}^{k}(Y)$, $k=1,2,\dots,d$ are defined as
-$\operatorname{Mov}^{k}(Y)=\big{\{}\alpha\in N^{1}(Y)_{\mathbb{R}}:\mathrm{B}_{+}(\alpha)\text{ has codimension $\geq k$}\big{\}}.$
-$\operatorname{Mov}^{k}(Y)$ are open cones in $N^{1}(Y)_{\mathbb{R}}$ and we have obvious inclusions
-$\operatorname{Mov}^{d}\subseteq\operatorname{Mov}^{d-1}\subseteq\cdots\subseteq\operatorname{Mov}^{1}.$
-Note that $\operatorname{Mov}^{1}$ is the big cone and $\operatorname{Mov}^{d}$ is the ample cone.
-Let $k$ be an algebraically closed field, $C/k$ be a curve and $E$ be a vector bundle on $C$.
-Note that $N^{1}(\mathbb{P}(E))_{\mathbb{R}}$ is two dimensional, generated by $f$ and $\xi=\mathcal{O}_{\mathbb{P}(E)}(1)$.
-The ample cone is given by extremal ray $f$ and $\xi-\mu_{\min(E)}f$, and the big cone is given by extremal ray $f$ and $\xi-\mu_{\max(E)}f$.
-The former is a theorem of Miyaoka [21] and the latter can be found in [15].
+Let $q:W/W_{P}\longrightarrow W_{Q}\backslash W/W_{P}$ be the canoniacl projection. For $w\in W_{Q}\backslash W/W_{P}$, let $m_{w}=\#q^{-1}(w)$. Then
+$h_{\mathcal{L}_{\lambda}}(X)=\frac{1}{\#W/W_{P}}\sum_{w\in W_{Q}\backslash W/W_{P}}m_{w}\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle.$
+Note that $\zeta_{w}=\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ in Theorem C are successive minima of the height function $h_{\mathcal{L}_{\lambda}}$ on $X$.
+Thus Theorem C builds an equality between the variety height and a weighted average of the successve mimina.
+One may recognize this as a refinement of Zhang’s inequality of successive minima [25]
+$e_{1}\geq h_{\mathcal{L}_{\lambda}}(X)\geq\frac{1}{d+1}\sum_{i=1}^{d+1}e_{i}.$
+into an equality. Here $e_{i}$ is the the successive minima of Zhang.
+We remark that it is not possible to build such an equality in general.
+Consider the height function on $\mathbb{P}^{n}_{\mathbb{Q}}$ induced by the Hermitian line bundle $\big{(}\mathcal{O}_{\mathbb{P}^{n}_{\mathbb{Z}}}(1),\|\cdot\|_{\text{FS}}\big{)}$.
+The variety height is a rational number $\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{i}\frac{1}{j}$ while the successive minima are irrational numbers $\tfrac{1}{2}\log(k+1)$, $k=0,1,\dots,n$.
+1.3. Base loci and movable cones
+It is a general philosophy in Arakelov geometry that the positivity of $\overline{L}$ is equivalent to the positivity of the induced height function $h_{\overline{L}}$.
+The study of essential minimum and pseudo-effectivity [8, 1, 22] provides an evidence.
+Our third result (Theorem D below) provides a new evidence to support the above mentioned philosophy.
+Keep the notations in Theorem A.
+Let $f\in N^{1}(F/P)$ be the class of a vertical fiber and $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$ be the augmented base locus of $\mathcal{L}_{\lambda}-tf$ for $t\in\mathbb{R}$.
+They measure the positivity of the line bundle $\mathcal{L}_{\lambda}$.
+Theorem D (Theorem 4.1).
+$\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)=\mathcal{X}\Bigg{\backslash}\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle>t}\mathcal{C}_{w}=\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\leq t}\mathcal{C}_{w}.$
+The height filtration $Z_{t}$ in Theorem A measures the positivity of the height function $h_{\mathcal{L}_{\lambda}}$, and we see that $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$ is just the Zariski closure of $Z_{t}$ in $\mathcal{X}$ for every $t\not=\zeta_{w}$.
+At critical points, $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$ is upper continuous while $Z_{t}$ is lower continuous.
+A class $\alpha\in N^{1}(F/P)_{\mathbb{R}}$ is called $k$-movable if $\mathrm{B}_{+}(\alpha)$ has codimension $\geq k$.
+As a direct corollary of Theorem D, we compute the movable cones on flag bundles.
+(Please see §4.3 for the notations.)
+Theorem E (Theorem 4.2).
+The $k$-th movable cone $\operatorname{Mov}^{k}(F/P)$ coincides with the cone defined by
+  1. (1)
+    $\langle\alpha^{\vee},\cdot\rangle<0$ for any $\alpha\in\Delta\backslash\Delta_{P}$, and
+  2. (2)
+    $\langle{\mathrm{deg}}(F_{Q}),w\cdot\rangle>0$ for any $w\in W/W_{P}$ with $\ell(w)\geq n-k+1$.
+Let $E$ be a vector bundle on a curve $C$.
+The ample cone of $\mathbb{P}(E)$ is computed in [21] and the big cone of $\mathbb{P}(E)$ is computed in [12].
 Fulger and Lehmann computed $\operatorname{Mov}^{2}(\mathbb{P}(E))$ in [13].
 In [4], the big cone of the Grassmann bundle of $E$ is computed and in [5], the ample cone of flag bundles of $E$ is computed.
-As a direct consequence of Theorem C, we compute the $k$-th movable cones of flag bundles $F/P$ over a curve $C$ for all $k$.
-Since any element in $N^{1}(F/P)_{\mathbb{R}}$ can be written as $\mathcal{L}_{\lambda}-tf$, we may define two functions $\langle\alpha^{\vee},\cdot\rangle$ and $\langle{\mathrm{deg}}(F_{Q}),w\cdot\rangle$ on $N^{1}(F/P)_{\mathbb{R}}$ as
-  • $\langle\alpha^{\vee},\cdot\rangle$ sends $\mathcal{L}_{\lambda}-tf$ to $\langle\alpha^{\vee},\lambda\rangle$.
-  • $\langle{\mathrm{deg}}(F_{Q}),w\cdot\rangle$ sends $\mathcal{L}_{\lambda}-tf$ to $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle-t$.
-These functions are well-defined, and
-Theorem D (Theorem 4.9).
-The $k$-th movable cone $\operatorname{Mov}^{k}(F/P)$ is the cone defined by $\langle\alpha^{\vee},\cdot\rangle<0$ for any $\alpha\in\Delta\backslash\Delta_{P}$ and $\langle{\mathrm{deg}}(F_{Q}),w\cdot\rangle>0$ for any $w\in W/W_{P}$ with $\ell(w)\geq n-k+1$.
-In particular, we find that $\mathcal{L}_{\lambda}$ is $k$-movable $\Longleftrightarrow$ the $k$-th minimum of Zhang of $h_{\mathcal{L}_{\lambda}}$ is positive.
+Our Theorem E is a simultaneous generalization of all these results.
+Theorem F (Corollary 4.3).
+Let $\mathcal{L}$ be a line bundle on $F/P$.
+Then $\mathcal{L}$ is $k$-movable if and only if $e_{k}(h_{\mathcal{L}})>0$.
+Here $e_{k}(h_{\mathcal{L}})$ is the $k$-th minimum of Zhang of the height function $h_{\mathcal{L}}$.
+Take $k=1$, we recover [1, Theorem 1.1] (in this special case) that $\mathcal{L}_{\lambda}$ big $\Longleftrightarrow$ $\zeta_{\operatorname{ess}}(h_{\mathcal{L}_{\lambda}})>0$.
 1.4. Example: Grassmann bundles
-We provide here the computation of essential minimum and big cone on Grassmann bundles.
-Let $k$ be an algebraically closed field and $C/k$ be a curve with function field $K=k(C)$.
+We provide here the computation of the essential minimum and the big cone on Grassmann bundles.
+Let $k$ be an algebraically closed field and $C$ be a curve over $k$ with function field $K=k(C)$.
 Let $E$ be a vector bundle of rank $n$ on $C$ with Harder-Narasimhan filtration $0\subsetneq E_{1}\subsetneq\cdots\subsetneq E_{r}$.
 Let $B$ be the subgroup of $\operatorname{GL}_{n}$ of all upper triangular matrices, and let $Q$ be the parabolic subgroup containing $B$ of type $(\operatorname{rank}(E_{1}),\operatorname{rank}(E_{2}),\dots,\operatorname{rank}(E_{r}))$.
-Here we say a parabolic subgroup containing $B$ is of type $(a_{1},\dots,a_{r})$ if it consists of upper block triguluar matrices of shape
+Here we say a parabolic subgroup containing $B$ is of type $(a_{1},\dots,a_{r})$ if it consists of upper block trianguluar matrices of the shape
 $\left[\begin{array}[]{cccc}A_{1}&*&\cdots&*\\ &A_{2}&\cdots&*\\ &&\cdots&*\\ &&&A_{r}\end{array}\right],\quad A_{i}\in\operatorname{GL}_{a_{i}}.$
 Let $F(E)$ be the frame bundle of $E$, which is a $\operatorname{GL}_{n}$-principal bundle.
-Then the canonical reduction of $F(E)$ is $F_{Q}$ to $Q$ where $F_{Q}$ is the $Q$-bundle of frames respecting the filtration $0\subsetneq E_{1}\subsetneq\cdots\subsetneq E_{r}$.
-Let $\mu_{i}$ ($i=1,\dots,n$) be the $i$-th number in the sequence
+Then the canonical reduction of $F(E)$ to $Q$ is $F_{Q}$, the $Q$-bundle of frames respecting the filtration $0\subsetneq E_{1}\subsetneq\cdots\subsetneq E_{r}$.
+Let $\mu_{i}$ ($1\leq i\leq n$) be the $i$-th number in the sequence
 $\underbrace{\mu(E_{1}),\dots,\mu(E_{1})}_{\text{$\operatorname{rank}(E_{1})$ times}},\underbrace{\mu(E_{2}/E_{1}),\dots,\mu(E_{2}/E_{1})}_{\text{$\operatorname{rank}(E_{2}/E_{1})$ times}},\dots,\underbrace{\mu(E/E_{r-1}),\dots,\mu(E/E_{r-1})}_{\text{$\operatorname{rank}(E/E_{r-1})$ times}}.$
@@
 Let $\operatorname{Gr}_{r}(E)=F(E)/P$ be the Grassmann bundle of $r$-dimension quotients, where $P$ is the parabolic subgroup containing $B$ of type $(n-r,r)$.
-Identify the Weyl group of $G$ with $S_{n}$ and of $P$ with $S_{n-r}\times S_{r}$ as usual.
-Then $W/W_{P}=S_{n}\big{/}(S_{n-r}\times S_{r})$ can be identified with the multi-index set $\big{\{}(i_{1},\dots,i_{r}):1\leq i_{1}<i_{2}<\cdots<i_{r}\leq n\big{\}}$ via
+Identify the Weyl group of $G$ with $S_{n}$ and of $P$ with $S_{n-r}\times S_{r}$ as usual and $W/W_{P}=S_{n}\big{/}(S_{n-r}\times S_{r})$ with the multi-index set $\big{\{}(i_{1},\dots,i_{r}):1\leq i_{1}<i_{2}<\cdots<i_{r}\leq n\big{\}}$ via
 $\left[\begin{matrix}1&2&\cdots&n-r&n-r+1&\cdots&n\\ *&*&\cdots&*&i_{1}&\cdots&i_{r}\end{matrix}\right]\longleftrightarrow(i_{1},\dots,i_{r}).$
@@
 Let $h_{\mathcal{O}(1)}$ be the height function on $\operatorname{Gr}_{r}(E_{K})$ induced by the model $(\operatorname{Gr}_{r}(E),\mathcal{O}(1))$.
-By Theorem A, the successive minima are $\langle{\mathrm{deg}}(F_{Q}),I\det_{2}\rangle=\sum_{i\in I}\mu_{i}$.
-(Theorem A says the successive minima are $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ for $w\in W_{Q}\backslash W/W_{P}$, but these numbers are really just $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ for $w\in W/W_{P}$.) In particular, the longest element $I_{0}=(1,2,\dots,r)$ gives the essential minimum
-$\zeta_{\text{ess}}=\langle{\mathrm{deg}}(F_{Q}),I_{0}\det\nolimits_{2}\rangle=\mu_{1}+\cdots+\mu_{r}.$
+Let $I_{0}=(1,2,\dots,r)\in W/W_{P}$ be the longest Weyl element.
+By Theorem A, the essential minimum of $h_{\mathcal{O}(1)}$ is $\zeta_{\operatorname{ess}}=\langle{\mathrm{deg}}(F_{Q}),I_{0}\det_{2}\rangle=\mu_{1}+\cdots+\mu_{r}$.
 The only element in $\Delta\backslash\Delta_{P}$ is $\alpha=\lambda_{n-r}-\lambda_{n-r+1}$ and the function $\langle\alpha^{\vee},\cdot\rangle$ on $N^{1}(\operatorname{Gr}_{r}(E))$ is given by $\mathcal{O}(1)\longmapsto\langle\alpha^{\vee},\det_{2}\rangle<0$ and $f\longmapsto 0$.
-Let $I_{0}=(1,2,\dots,r)\in W/W_{P}$ be the longest element.
-Then the function $\langle{\mathrm{deg}}(F_{Q}),I_{0}\cdot\rangle$ on $N^{1}(\operatorname{Gr}_{r}(E))$ is given by $\mathcal{O}(1)\longmapsto\mu_{1}+\cdots+\mu_{r}$ and $f\longmapsto 1$.
-By Theorem C, the big cone is given by $\langle\alpha^{\vee},\cdot\rangle<0$ and $\langle{\mathrm{deg}}(F_{Q}),I_{0}\cdot\rangle>0$, and one sees that it is the cone given by extremal rays $f$ and $\mathcal{O}(1)-(\mu_{1}+\dots+\mu_{r})f$.
-1.6. Notations and conventions
-In the following, let $k$ be an algebraically closed field of characteristic zero and $C/k$ be a projective smooth curve with function field $K=k(C)$.
-Let $G/k$ be a connected reductive group and denote the identity component of its center by $Z(G)$.
+$\langle{\mathrm{deg}}(F_{Q}),I_{0}\cdot\rangle$ on $N^{1}(\operatorname{Gr}_{r}(E))$ is given by $\mathcal{O}(1)\longmapsto\mu_{1}+\cdots+\mu_{r}$ and $f\longmapsto 1$.
+By Theorem F, the big cone is given by $\langle\alpha^{\vee},\cdot\rangle<0$ and $\langle{\mathrm{deg}}(F_{Q}),I_{0}\cdot\rangle>0$.
+One sees that it is the cone given by extremal rays $f$ and $\mathcal{O}(1)-(\mu_{1}+\dots+\mu_{r})f$.
+1.5. Acknowledgements
+We thank Xinyi Yuan and Ye Tian for their consistent supports. We thank Jie Liu, Marc Besson, Hongsheng Hu, Mihai Fulger and Brian Lehmann for discussions and email correspondences.
+The idea of comparing height filtrations and base loci was proposed by Huayi Chen in a private conversation.
+The idea of computing concave transform was inspired from discussions with François Ballaÿ.
+The idea of proving Proposition 3.4 by Weyl symmetry was inspired from discussions with Tao Gui.
+We wish to thank them especially.
+2. Height filtrations and successive minima
+Let $k$ be an algebraically closed field of characteristic zero and $C$ be a projective smooth curve over $k$ with function field $K=k(C)$.
+Let $G$ be a connected reductive group over $k$.
+Let $P\subseteq G$ be a parabolic subgroup and $\lambda:P\longrightarrow\mathbb{G}_{m}$ be a strictly antidominant character.
+Let $F$ be a principal $G$-bundle over $C$.
+Let $\mathcal{X}=F/P$ and $X=(F/P)_{K}$.
+Let $\mathcal{L}_{\lambda}=F\times_{P}k_{\lambda}$ and $h_{\mathcal{L}_{\lambda}}:X(\overline{K})\longrightarrow\mathbb{R}$ be the induced height function.
+${X=(F/P)_{K}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{X}=F/P\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Spek}(K)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C}$
+Let $F_{Q}$ be the canonical reduction of $F$ to $Q$, and ${\mathrm{deg}}(F_{Q})\in X(T)^{\vee}_{\mathbb{Q}}$ be the induced cocharacter.
+Let $W,W_{P}$ and $W_{Q}$ be the Weyl group of $G,L_{P}$ and $L_{Q}$.
+For $w\in W_{Q}\backslash W/W_{P}$, let $C_{w}=(F_{Q}\times_{Q}QwP/P)_{K}\subseteq X$ be the corresponding Schubert cell.
+The main result in this section is the following theorem:
+Theorem 2.1.
+For any $t\in\mathbb{R}$, let $Z_{t}\subseteq X$ be the Zariski closure of the set $\big{\{}x\in X(\overline{K}):h_{\mathcal{L}_{\lambda}}(x)<t\big{\}}$. Then
+$Z_{t}=X\Bigg{\backslash}\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\geq t}C_{w}=\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle<t}C_{w}.$
+Corollary 2.2.
+The filtration $\big{\{}Z_{t}:t\in\mathbb{R}\big{\}}$ is finite, and is given by successively deleting Schubert cells $C_{w}$ with jumping points $\zeta_{w}=\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ for $w\in W_{Q}\backslash W/W_{P}$.
+Let’s call $\big{\{}Z_{t}:t\in\mathbb{R}\big{\}}$ the height filtration, and call its jumping points $\zeta_{w}$ the successive minima.
+Note that our definition of successive minima are slightly different with Zhang [25].
+Let $d=\dim G/P$, Zhang considers only dimension jumps $e_{i}=\inf\big{\{}t:\dim Z_{t}\geq d-i+1\big{\}}$, $i=1,\dots,d+1$.
+In the following, $e_{i}$ will be called Zhang’s successive minima to avoid ambiguity.
+For $w\in W_{Q}\backslash W/W_{P}$, let
+$\ell(w)=\max_{\sigma\in W_{Q}wW_{P}}\min_{\tau\in\sigma W_{P}}\ell(\tau)$
+be the length function, which is nothing but the dimension of the cell $C_{w}$.
+Corollary 2.3.
+Zhang’s successive minima are $e_{i}=\min\big{\{}\zeta_{w}:\ell(w)\geq d-i+1\big{\}}$.
+2.1. Vector bundles
+Let $E$ be a vector bundle on $C$.
+The degree of $E$ is ${\mathrm{deg}}(E):={\mathrm{deg}}(\det(E))$ and the slope of $E$ is $\mu(E):={\mathrm{deg}}(E)/\operatorname{rk}(E)$.
+It is called slope semistable if for every subbundle $F$ of $E$, $\mu(F)\leq\mu(E)$.
+This is equivalent to $\mu(Q)\geq\mu(E)$ for every quotient bundle $Q$ of $E$.
+There exists uniquely a filtration $0=E_{0}\subseteq E_{1}\subseteq\cdots\subseteq E_{r}=E$ such that
+  1. (1)
+    $E_{i}/E_{i-1}$ is semistable;
+  2. (2)
+    $\mu(E_{i}/E_{i-1})>\mu(E_{i+1}/E_{i})$.
+This filtration is called the Harder-Narasimhan filtration of $E$.
+$\mu(E_{1})$ is called maximal slope of $E$ and is also denoted by $\mu_{\max}(E)$.
+$\mu(E/E_{r-1})$ is called minimal slope of $E$ and is also denoted by $\mu_{\min}(E)$.
+2.2. Principal bundles
+For any linear algebraic group $\Gamma/k$, let $X(\Gamma)=\mathrm{Hom}(\Gamma,\mathbb{G}_{m})$ denote the character group of $\Gamma$.
+For a cocharacter $f\in X(\Gamma)^{\vee}$ and a character $\lambda\in X(\Gamma)$, we shall denote the pairing by $\langle f,\lambda\rangle$.
+For any $\lambda\in X(\Gamma)$, denote by $k_{\lambda}$ the one-dimensional representation on the vector space $k$ with $\Gamma$ acting by $\lambda$.
+A principal $\Gamma$-bundle on $C$ is a variety $F$ equipped with a right action of $\Gamma$ and a $\Gamma$-equivariant smooth morphism $F\longrightarrow C$ such that the map
+$F\times_{C}(C\times\Gamma)\longrightarrow F\times_{C}F,\quad(f,(x,g))\longmapsto(f,fg)$
+is an isomorphism.
+Attached to a principal $\Gamma$-bundle $F$, one has an associated cocharacter
+${\mathrm{deg}}(F):X(\Gamma)\longrightarrow\mathbb{Z},\quad\lambda\longmapsto\langle{\mathrm{deg}}(F),\lambda\rangle={\mathrm{deg}}(F\times_{\Gamma}k_{\lambda}).$
+Here ${\mathrm{deg}}(F\times_{\Gamma}k_{\lambda})$ is the degree of the line bundle $F\times_{\Gamma}k_{\lambda}$ on the curve $C$.
+Let $H$ be a closed subgroup of a linear algebraic group $\Gamma$ over $k$.
+A reduction of structure group of $F$ to $H$ is a pair $(F_{H},\phi)$ where $F_{H}$ is a principal $H$-bundle and $\phi:F_{H}\times_{H}\Gamma\simeq F$ is an isomorphism.
+By the universal property of the quotient $F/H$, the assignment to a section $\sigma:C\longrightarrow F/H$ the reduction $\sigma^{*}F$ of $F$ to $H$ is a one-one correspondence between reductions of structure group of $F$ to $H$ and sections of $F/H\longrightarrow C$.
+2.3. Reductive groups, characters and cocharacters
+Let $G$ be a connected reductive group over $k$.
 Fix a Borel subgroup $B\subseteq G$ and a maximal torus $T\subseteq B$.
@@
 For any $\alpha\in\Delta$, we denote by $\alpha^{\vee}$ the corresponding simple coroot.
-We shall consider only parabolic subgroup $P\subseteq G$ containing $B$.
-For any parabolic subgroup $P$, let $W_{P}\subseteq W$ be the Weyl group $W(L_{P})$ of the Levi factor $L_{P}\subseteq P$ and $\Delta_{P}\subseteq\Delta$ be the simple roots of $L_{P}$.
-For parabolic subgroups $P,Q\subseteq G$ and any double coset $w\in W_{Q}\backslash W/W_{P}$, set $\ell(w)=\max_{\sigma\in W_{Q}wW_{P}}\min_{\tau\in\sigma W_{P}}\ell(\tau)$, which equals to $\operatorname{dim}QwP/P$.
-Here by abuse of notations, $\ell(\tau)$ is the usual length of the Weyl element $\tau$.
-For any linear algebraic group $\Gamma/k$, let $X(\Gamma)$ denote the character group $\mathrm{Hom}(\Gamma,\mathbb{G}_{m})$.
-For any $\mathbb{Z}$-algebra $A$, denote by $X(\Gamma)_{A}$ the tensor product $X(\Gamma)\otimes_{\mathbb{Z}}A$.
-For any linear functional $f$ on $X(\Gamma)$, we shall usually denote $f(\alpha)$ by $\langle f,\alpha\rangle$ for $\alpha\in X(\Gamma)$.
-For any $\alpha\in X(\Gamma)$, denote by $k_{\alpha}$ the one-dimensional algebraic representation on the vector space $k$ with $\Gamma$ acting by $\alpha$.
-1.7. Organization of the paper
-Let $F$ be a principal $G$-bundle on $C$ and $\lambda:P\longrightarrow\mathbb{G}_{m}$ be a strictly anti-dominant character.
-Then $F/P$ is a flag bundle over $C$ and $\mathcal{L}_{\lambda}=F\times_{P}k_{\lambda}$ on $F/P$ is a relatively ample line bundle.
-Let $X=(F/P)_{K}$ and $h_{\mathcal{L}_{\lambda}}$ be the height function on $X$ induced by $(F/P,\mathcal{L}_{\lambda})$.
-In Chapter $2$, we compute the height filtration and successive minima of $h_{\mathcal{L}_{\lambda}}$.
-In Chapter $3$, we enhance Zhang’s inequality into an equality by convex geometry.
-In Chapter $4$, we compute the augmented base loci of $\mathcal{L}_{\lambda}-tf$ on $F/P$.
-As a corollary we get the movable cones $\operatorname{Mov}^{k}(F/P)$.
-2. Height filtrations and successive minima
-2.1. Vector bundles, slope stability and Harder-Narasimhan filtration
-Let $E$ be a vector bundle on $C$.
-The degree of $E$ is ${\mathrm{deg}}(E)={\mathrm{deg}}(\det(E))$ and the slope of $E$ is $\mu(E)={\mathrm{deg}}(E)/\operatorname{rk}(E)$.
-It is called slope semistable if for every subbundle $F\subseteq E$, $\mu(F)\leq\mu(E)$.
-This is equivalent to $\mu(Q)\geq\mu(E)$ for every quotient bundle $Q$ of $E$.
-In general, there exists uniquely a filtration $0=E_{0}\subseteq E_{1}\subseteq\cdots\subseteq E_{r}=E$ such that
-  • $E_{i}/E_{i-1}$ is semistable;
-  • $\mu(E_{i}/E_{i-1})>\mu(E_{i+1}/E_{i})$.
-This filtration is called the Harder-Narasimhan filtration of $E$ and we denote $\mu_{i}(E)=\mu(E_{i}/E_{i-1})$.
-The first and the last term are of particular interests.
-The first one $\mu_{1}(E)=\mu(E_{1})$ is called maximal slope of $E$ and is also denoted by $\mu_{\max}(E)$.
-The last one $\mu_{r}(E)=\mu(E/E_{r-1})$ is called minimal slope of $E$ and is also denoted by $\mu_{\min}(E)$.
-2.2. Principal bundles
-For any linear algebraic group $\Gamma/k$, a principal $\Gamma$-bundle on $C$ is a variety $F$ equipped with a right action of $\Gamma$ and a $\Gamma$-equivariant smooth morphism $F\longrightarrow C$ such that the map
-$F\times_{C}(C\times\Gamma)\longrightarrow F\times_{C}F,\quad(f,(x,g))\longmapsto(f,fg)$
-is an isomorphism.
-Attached to any principal $\Gamma$-bundle $F$, one has the degree map
-${\mathrm{deg}}(F):X(\Gamma)\longrightarrow\mathbb{Z},\quad\lambda\longmapsto\langle{\mathrm{deg}}(F),\lambda\rangle={\mathrm{deg}}(F\times_{\Gamma}k_{\lambda}).$
-Here ${\mathrm{deg}}(F\times_{\Gamma}k_{\lambda})$ is the degree of the line bundle $F\times_{\Gamma}k_{\lambda}$ on the curve $C$.
-Let $H$ be a closed subgroup of a linear algebraic group $\Gamma/k$.
-A reduction of structure group of $F$ to $H$ is a pair $(F_{H},\varphi)$ where $F_{H}$ is a principal $H$-bundle and $\varphi:F_{H}\times_{H}\Gamma\cong F$ is an isomorphism.
-By the universal property of the quotient $F/H$, the assignment to any section $\sigma:C\longrightarrow F/H$ the reduction $\sigma^{*}F$ of $F$ to $H$ is a one-one correspondence between reductions of structure group of $F$ to $H$ and sections of $F/H\longrightarrow C$.
-Let $F$ be a principal $G$-bundle on $C$ for the connected reductive group $G$. Note that for any parabolic subgroup $P\subseteq G$ with Levi subgroup $L_{P}$, the natural inclusion
+We shall consider only parabolic subgroups containing $B$.
+For such a parabolic subgroup $P$, let $W_{P}\subseteq W$ be the Weyl group $W(L_{P})$ of the Levi factor $L_{P}\subseteq P$ and $\Delta_{P}\subseteq\Delta$ be the simple roots of $L_{P}$.
+Note that the natural inclusion
 $X(P)\longrightarrow X(L_{P})\longrightarrow X(Z(L_{P}))$
@@
 and by taking duals, we get the so-called slope map $X(P)^{\vee}_{\mathbb{Q}}\longrightarrow X(T)^{\vee}_{\mathbb{Q}}$ introduced in [23, §2.1.3].
-Let $F_{P}$ be a reduction of $F$ to $P$.
-By applying the slope map to ${\mathrm{deg}}(F_{P})$ we can define $\langle{\mathrm{deg}}(F_{P}),\lambda\rangle$ for any $\lambda\in X(T)$.
-2.3. Semistability and canonical reduction
-We mainly follow [5].
-Definition 2.1 (semistable).
-A principal $G$-bundle $F$ on $C$ is called semistable if for any parabolic subgroup $P$, any reduction $F_{P}$ of $F$ to $P$ and any dominant character $\lambda$ of $P$ which is trivial on $Z(G)$, we have $\langle{\mathrm{deg}}(F_{P}),\lambda\rangle\leq 0$.
-Definition 2.2 (canonical reduction).
-Let $F$ be a principal $G$-bundle on $C$. A reduction $F_{P}$ of $F$ to a parabolic subgroup $P$ is called canonical if the following two conditions hold:
+In other words, a cocharacter on $X(P)_{\mathbb{Q}}$ can be extended canonically to $X(T)_{\mathbb{Q}}$.
+2.4. The canonical reduction to a parabolic subgroup
+Let $G$ be a connected reductive group over $k$ and $F$ be a principal $G$-bundle over $C$.
+Let $F_{P}$ be a reduction of $F$ to a parabolic subgoup $P$.
+Let ${\mathrm{deg}}(F_{P})\in X(T)_{\mathbb{Q}}^{\vee}$ be the induced (rational) cocharacter.
+The $G$-bundle $F$ is called semistable if for any parabolic subgroup $P$, any reduction $F_{P}$ of $F$ to $P$ and any dominant character $\lambda$ of $P$ which is trivial on $Z(G)$, we have $\langle{\mathrm{deg}}(F_{P}),\lambda\rangle\leq 0$.
+Among all filtrations of a vector bundle, there is a canonical one (the Harder-Narasimhan filtration).
+Similarly, among all reduction to parabolic subgroups, there is a canonical one.
+A reduction $F_{P}$ of $F$ to a parabolic subgroup $P$ is called canonical if the following two conditions hold:
   1. (1)
@@
 Behrend in [2] proved canonical reduction exists uniquely.
-When $G=\operatorname{GL}_{n}$, the above definition of semistability recovers Mumford’s slope stability of vector bundles and the above deinition of canonical reduction recovers the Harder-Narasimhan filtration of vector bundles.
-2.4. Height function induced by $(F/P,\mathcal{L}_{\lambda}$)
-Let $F$ be a principal $G$-bundle with canonical reduction $F_{Q}$ to some parabolic subgroup $Q\subseteq G$. Let $P\subseteq G$ be a parabolic subgroup. Set $\mathcal{X}=F/P$ and $X=\mathcal{X}_{K}$.
-A character $\lambda:P\longrightarrow\mathbb{G}_{m}$ is called strictly anti-dominant if the natural pairing $\langle\alpha^{\vee},\lambda\rangle<0$ for any $\alpha\in\Delta\backslash\Delta_{P}$.
-Let $\lambda:P\longrightarrow\mathbb{G}$ be a stricly anti-dominant character.
-Then the line bundle $M_{\lambda}=G\times_{P}k_{\lambda}$ on $G/P$ is ample.
-Therefore $\mathcal{L}_{\lambda}=F\times_{G}M_{\lambda}$ is a relatively ample line bundle on $\mathcal{X}=F\times_{G}G/P$ and induces a height function
-$\quad\quad\quad h_{\mathcal{L}_{\lambda}}:X(\overline{K})$ $\longrightarrow\mathbb{Q},\quad x\longmapsto\frac{\mathcal{L}_{\lambda}\cdot\overline{\{x\}}}{{\mathrm{deg}}(x)}$
-where $\overline{\{x\}}$ is the Zariski closure of $x$ in $\mathcal{X}$.
+Note that when $G=\operatorname{GL}_{n}$, the above definition of semistability recovers Mumford’s slope stability of vector bundles and the above deinition of canonical reduction recovers the Harder-Narasimhan filtration of vector bundles.
+2.5. Arakelov geometry on flag varieties over function fields
+Let $G$ be a connected reductive group over $k$ and $F$ be a principal $G$-bundle over $C$.
+Let $P\subseteq G$ be a parabolic subgroup and $\lambda:P\longrightarrow\mathbb{G}_{m}$ be a strictly antidominant character.
+Let $\mathcal{X}=F/P$ and $X=(F/P)_{K}$.
+${X=(F/P)_{K},L\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{X}=F/P,\mathcal{L}_{\lambda}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Spek}(K)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C}$
+Let $\mathcal{L}_{\lambda}=F\times_{P}k_{\lambda}$, $L=\mathcal{L}_{\lambda,K}$ and $h_{\mathcal{L}_{\lambda}}:X(\overline{K})\longrightarrow\mathbb{R},x\longmapsto\frac{\overline{\{x\}}\cdot\mathcal{L}_{\lambda}}{{\mathrm{deg}}(x)}$ be the induced height function.
+Here $\overline{\{x\}}$ is the Zariski closure of $x$ in $\mathcal{X}$.
+For any $t\in\mathbb{R}$, let $Z_{t}\subseteq X$ be the Zariski closure of the set $\big{\{}x\in X(\overline{K}):h_{\mathcal{L}_{\lambda}}(x)<t\big{\}}$.
+Let $\zeta_{\operatorname{ess}}(X):=\inf\big{\{}t:Z_{t}=X\big{\}}$ be the essential minimum of $h_{\mathcal{L}_{\lambda}}$ on $X$.
+Let $F_{Q}$ be the canonical reduction of $F$ to $Q$, and ${\mathrm{deg}}(F_{Q})\in X(T)^{\vee}_{\mathbb{Q}}$ be the induced cocharacter.
+Let $W,W_{P}$ and $W_{Q}$ be the Weyl group of $G,L_{P}$ and $L_{Q}$.
+For $w\in W_{Q}\backslash W/W_{P}$, write $\mathcal{C}_{w}=F_{Q}\times_{Q}QwP/P$, $\mathcal{X}_{w}=F_{Q}\times_{Q}\overline{QwP}/P$, $C_{w}=\mathcal{C}_{w,K}$ and $X_{w}=\mathcal{X}_{w,K}$.
+Note that for any $w^{\prime}\in W_{Q}$ and $\lambda\in X(T)$, we have $w^{\prime}\lambda-\lambda\in\mathbb{Z}[\Delta_{Q}]$ consequently $\langle{\mathrm{deg}}(F_{Q}),w^{\prime}\lambda\rangle=\langle{\mathrm{deg}}(F_{Q}),\lambda\rangle$.
+Note also that for any $\lambda\in X(P)$ and $w\in W_{P}$, $w\lambda=\lambda$.
+So the number $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ is well-defined for any $\lambda\in X(P)$ and $w\in W_{Q}\backslash W/W_{P}$.
+$\mathcal{X}_{w}$ is a closed subscheme of $\mathcal{X}$, so $\mathcal{L}_{\lambda}|_{\mathcal{X}_{w}}$ induces a height function $h_{\mathcal{L}_{\lambda}}:X_{w}(\overline{K})\longrightarrow\mathbb{R}$, which is nothing but the restriction of $h_{\mathcal{L}_{\lambda}}:X(\overline{K})\longrightarrow\mathbb{R}$ to $X_{w}(\overline{K})$.
+Let $\zeta_{\operatorname{ess}}(X_{w})$ be the essential minimum of $h_{\mathcal{L}_{\lambda}}$ on $X_{w}$.
+Theorem 2.1 follows from the following two propositions.
+Proposition 2.4.
+$h_{\mathcal{L}_{\lambda}}(x)\geq\langle{\mathrm{deg}}F_{Q},w\lambda\rangle$ for any $x\in C_{w}(\overline{K})$.
+Proposition 2.5.
+$\zeta_{\operatorname{ess}}(X_{w})=\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$.
+Theorem 2.1.
+For any $t\in\mathbb{R}$, let $Z_{t}\subseteq X$ be the Zariski closure of the set $\big{\{}x\in X(\overline{K}):h_{\mathcal{L}_{\lambda}}(x)<t\big{\}}$. Then
+$Z_{t}=X\Bigg{\backslash}\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\geq t}C_{w}=\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle<t}C_{w}.$
+Proof.
+We write $\zeta_{w}=\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ and $h=h_{\mathcal{L}_{\lambda}}$ to simplify notations.
+For a closed subvariety $Y$ of $X$, let $Z_{t}(Y)$ be the Zariski closure in $Y$ of $\big{\{}y\in Y(\overline{K}):h(y)<t\big{\}}$.
+By Proposition 2.4, $Z_{\zeta_{w}}(X_{w})\subseteq X_{w}\backslash C_{w}$, and we claim that $Z_{\zeta_{w}}(X_{w})=X_{w}\backslash C_{w}=\bigcup_{w^{\prime}<w}X_{w^{\prime}}$.
+In fact, suppose conversely that $Z_{\zeta_{w}}(X_{w})\subsetneq X_{w}\backslash C_{w}$.
+Then there exists $X_{w^{\prime}}\subsetneq X_{w}$ such that $Z_{\zeta_{w}}(X_{w})\cap X_{w^{\prime}}\subsetneq X_{w^{\prime}}$.
+This forces $Z_{\zeta_{w}}(X_{w^{\prime}})\subseteq Z_{\zeta_{w}}(X_{w})\cap X_{w^{\prime}}\subsetneq X_{w^{\prime}}$.
+But on $X_{w^{\prime}}$, Proposition 2.5 says the essential minimum is $\zeta_{w^{\prime}}<\zeta_{w}$, so $Z_{\zeta_{w}}(X_{w^{\prime}})=X_{w^{\prime}}$.
+Contradiction!
+Thus on $X_{w}$, the set $Z_{t}(X_{w})=X_{w}$ when $t>\zeta_{w}$ and $Z_{t}(X_{w})=X_{w}\backslash C_{w}$ when $t\leq\zeta_{w}$.
+Let $w_{0}\in W_{Q}\backslash W/W_{P}$ be the longest element.
+On $X=X_{w_{0}}$, when $t>\zeta_{w_{0}}$, $Z_{t}=X_{w_{0}}$ and when $t\leq\zeta_{w_{0}}$, $Z_{t}=X_{w_{0}}\backslash C_{w_{0}}=\bigcup_{w^{\prime}}X_{w^{\prime}}$ with $w^{\prime}<w_{0}$ and $\ell(w^{\prime})=\ell(w_{0})-1$.
+Repeating this procedure on each smaller $X_{w^{\prime}}$, we can complete the proof by induction.
+∎
+2.6. Proof of Proposition 2.4
 For $x\in X(K)$, let $\sigma_{x}:C\longrightarrow\mathcal{X}$ be the section induced by $x:\operatorname{Spec}(K)\longrightarrow X$ via valuative criterion, and let $F_{P,x}=\sigma_{x}^{*}F$ be the corresponding reduction to $P$.
@@
 and the lemma follows by taking degree. ∎
-2.5. A height lower bound in Schubert cells
-For $w\in W_{Q}\backslash W/W_{P}$, write $\mathcal{C}_{w}=F_{Q}\times_{Q}QwP/P$, $\mathcal{X}_{w}=F_{Q}\times_{Q}\overline{QwP}/P$, $C_{w}=\mathcal{C}_{w,K}$ and $X_{w}=\mathcal{X}_{w,K}$.
-We shall show that
-Proposition 2.4.
-For any $x\in C_{w}(\overline{K})$, $h_{\mathcal{L}_{\lambda}}(x)\geq\langle{\mathrm{deg}}F_{Q},w\lambda\rangle$.
-Note that for any $w^{\prime}\in W_{Q}$ and $\lambda\in X(T)$, we have $w^{\prime}\lambda-\lambda\in\mathbb{Z}[\Delta_{Q}]$ consequently $\langle{\mathrm{deg}}(F_{Q}),w^{\prime}\lambda\rangle=\langle{\mathrm{deg}}(F_{Q}),\lambda\rangle$.
-Note also that for any $\lambda\in X(P)$ and $w\in W_{P}$, $w\lambda=\lambda$.
-So the number $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ is well-defined for any $\lambda\in X(P)$ and $w\in W_{Q}\backslash W/W_{P}$.
 Definition 2.7 (relative position).
@@
 lies in $QwP_{C}$. Here we implicitly apply the injections
-$F_{P}\hookrightarrow F_{P}\times_{P}G\cong F,\quad F_{Q}\hookrightarrow F_{Q}\times_{Q}G\cong F.$
+$F_{P}\hookrightarrow F_{P}\times_{P}G\simeq F,\quad F_{Q}\hookrightarrow F_{Q}\times_{Q}G\simeq F.$
 Note that by the universal property of the quotient stacks $[G//Q\times P]$ and $[QwP//Q\times P]$, this definition coincides with the one in [23, §4.1].
@@
 $\Big{(}F_{Q}\times_{C}F_{Q}\Big{)}\times_{Q}QwP\longrightarrow(C\times Q)\times_{Q}QwP=C\times QwP.$
-Proposition 2.7 ([23], Proposition $4.6$).
+Proposition 2.9.
 If $F_{P}$ is in relative position $w$ with respect to $F_{Q}$, then $\langle{\mathrm{deg}}(F_{P}),\lambda\rangle\geq\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ for any anti-dominant character $\lambda$.
-Now we can prove Proposition 2.4.
+Proof.
+This is written in the proof of [23, Theorem 4.1]. See [23, §4.5.3]. ∎
 Proof of Proposition 2.4.
-Let $L/K$ be any finite field extension. Assume that $x\in C_{w}(L)$. Consider the following commutative diagram
+Let $L$ be any finite field extension of $K$. Assume that $x\in C_{w}(L)$. Consider the following commutative diagram
 […]
@@
 ∎
-2.6. Height filtration and successive minima of $h_{\mathcal{L}_{\lambda}}$
-In this subsection, we employ Ballaÿ’s theorem [2, Theorem 1.2] to compute essential minimum. We start by finding the maximal slope of $\pi_{*}\mathcal{L}_{\lambda}|_{\mathcal{X}_{w}}$.
+2.7. Proof of Proposition 2.5
+In this subsection, we employ Ballaÿ’s theorem [1, Theorem 1.2] to compute the essential minimum.
+Let $\pi:F/P\longrightarrow C$ be the structure morphism.
+We start by finding the maximal slope of $\pi_{*}\mathcal{L}_{\lambda}|_{\mathcal{X}_{w}}$.
 Lemma 2.10 ([14], Part I, §5.18).
@@
 Schieder computed the Harder-Narasimhan filtration of $F_{Q}\times_{Q}H^{0}(\overline{QwP}/P,M_{\lambda})$ in [23, §5.1].
-Let $V/k$ be a finite dimensional $Q$-representation with weight decomposition $V=\bigoplus_{\nu\in X(T)}V[\nu]$.
-The $Q$-filtration $V_{\bullet,{\mathrm{deg}}(F_{Q})}$ of $V$ is defined by
-$V_{t,{\mathrm{deg}}(F_{Q})}=\bigoplus_{\langle{\mathrm{deg}}(F_{Q}),\nu\rangle\geq t}V[\nu].$
+To state his result, we fix the following notations which will be used frequently.
+Definition 2.11.
+Let $V$ be a finite dimensional $Q$-representation over $k$ with weight decomposition $V=\bigoplus_{\mu\in X(T)}V[\mu]$. The $Q$-filtration $V_{\bullet,{\mathrm{deg}}(F_{Q})}$ of $V$ is defined by
+$V_{t,{\mathrm{deg}}(F_{Q})}=\bigoplus_{\langle{\mathrm{deg}}(F_{Q}),\mu\rangle\geq t}V[\mu].$
 Proposition 2.12.
@@
 For the second assertion, we only need to show the highest weights in $H^{0}(\overline{QwP}/P,M_{\lambda})$ belong to $w\lambda+\mathbb{Z}[\Delta_{Q}]$.
-When $Q=B$, $H^{0}(\overline{BwP}/P,M_{\lambda})$ has a single highest weight $w\lambda$ by the theory of Demazure modules [17, Chapter 14].
+When $Q=B$, $H^{0}(\overline{BwP}/P,M_{\lambda})$ has a single highest weight $w\lambda$ by the theory of Demazure modules [14, Chapter 14].
 For general $Q$, let $\pi:W/W_{P}\longrightarrow W_{Q}\backslash W/W_{P}$ be the canoniacl projection.
@@
 is injective, the highest weights in $H^{0}(\overline{QwP}/P,M_{\lambda})$ belong to $w\lambda+\mathbb{Z}[\Delta_{Q}]$. ∎
-Corollary 2.10.
-The essential minimum $\zeta_{1}(h_{\mathcal{L}_{\lambda}},X_{w})$ of $h_{\mathcal{L}_{\lambda}}$ on $X_{w}$ is $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$.
+Proof of Proposition 2.5.
+By Ballaÿ’s theorem [1, Theorem 1.2],
+$\quad\quad\quad\zeta_{\operatorname{ess}}(X_{w})$ $=\displaystyle\lim_{n\rightarrow\infty}\frac{\mu_{\max}(\pi_{*}\mathcal{L}_{n\lambda}|_{\mathcal{X}_{w}})}{n}$
+$=\lim_{n\rightarrow\infty}\frac{n\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle}{n}=\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle.$
+3. String polytopes and the Boucksom-Chen concave transform
+Keep the notations in §2.
+Let $w_{0}\in W$ be the longest Weyl element and fix a reduced decomposition $\underline{w_{0}}=s_{i_{1}}s_{i_{2}}\cdots s_{i_{N}}$, $N=\dim G/B$.
+Then for the antidominant weight $\lambda:P\longrightarrow\mathbb{G}_{m}$, we have the string valuation $\nu$ and the string polytope $\Delta_{\lambda}\subseteq\mathbb{R}^{N}$, which is a polytope of dimension $d=\dim G/P$.
+Let $\mathcal{P}_{\lambda}$ be the weight polytope and $p:\Delta_{\lambda}\longrightarrow\mathcal{P}_{\lambda}$ be the weight map.
+The main theorem of this section is the following:
+Theorem 3.1.
+The Boucksom-Chen concave transform of ${\mathcal{L}_{\lambda}}$ (with respect to a specific valuation) is $\Delta_{\lambda}\longrightarrow\mathbb{R},x\longmapsto\langle{\mathrm{deg}}(F_{Q}),p(x)\rangle$.
+Let $\mu$ be the Lebesgue measure on the real span of $\Delta_{\lambda}$, normalized with respect to $\mathbb{Z}^{N}$. As a direct corollary, we have by [6] that
+$h_{\mathcal{L}_{\lambda}}(X)=\frac{\mathcal{L}_{\lambda}^{\dim\mathcal{X}}}{\dim\mathcal{X}\cdot L^{\dim X}}=\frac{1}{\operatorname{Vol}(\Delta_{\lambda})}\int_{\Delta_{\lambda}}\langle{\mathrm{deg}}(F_{Q}),p(x)\rangle\,\mathrm{d}\mu(x).$
+The integration can be computed via Weyl symmetry:
+Theorem 3.2.
+Let $q:W/W_{P}\longrightarrow W_{Q}\backslash W/W_{P}$ be the canoniacl projection. For $w\in W_{Q}\backslash W/W_{P}$, let $m_{w}=\#q^{-1}(w)$. Then
+$h_{\mathcal{L}_{\lambda}}(X)=\frac{1}{\#W/W_{P}}\sum_{w\in W_{Q}\backslash W/W_{P}}m_{w}\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle.$
+Note that $\zeta_{w}=\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ are successive minima of the height function $h_{\mathcal{L}_{\lambda}}$ on $X$ by Theorem 2.1.
+Thus this theorem builds an equality between the variety height and a weighted average of the successve mimina.
+One may recognize this as a refinement of Zhang’s inequality of successive minima.
+3.1. Okounkov bodies and the Boucksom-Chen concave transform
+Let $\mathbf{k}$ be a field and $Y$ be a normal projective variety of dimension $d$ over $\mathbf{k}$.
+Let $L$ be a big line bundle on $Y$, $V_{n}=H^{0}(Y,nL)$ and $V_{\bullet}=\bigoplus_{n\geq 0}V_{n}$.
+Let $\nu:V_{\bullet}\backslash\{0\}\longrightarrow\mathbb{Z}^{r}$ be a valuation with one dimensional leaves, i.e.
+  • for $a\in\mathbf{k}^{\times}$ and $s\in V_{\bullet}\backslash\{0\}$, we have $\nu(as)=\nu(s)$.
+  • for $s_{1},s_{2}\in V_{\bullet}\backslash\{0\}$, we have $\nu(s_{1}s_{2})=\nu(s_{1})+\nu(s_{2})$ and $\nu(s_{1}+s_{2})\geq\min\big{\{}\nu(s_{1}),\nu(s_{2})\big{\}}$.
+  • the quotient space $\big{\{}s\in H^{0}(Y,nL):\nu(s)\geq a\big{\}}\Big{/}\big{\{}s\in H^{0}(Y,nL):\nu(s)>a\big{\}}$ is at most one dimensional.
+Let $W_{\bullet}\subseteq V_{\bullet}$ be a graded subalgebra.
+We define $\Delta_{\nu}(W_{\bullet})\subseteq\mathbb{R}^{r}$ to be the closure of the set $\bigcup_{n\geq 1}\tfrac{1}{n}\nu(W_{n}\backslash\{0\})$.
+In particular, we define the Okounkov body $\Delta_{\nu}(Y,L):=\Delta_{\nu}(V_{\bullet})$.
+A fundamental theorem is $\operatorname{Vol}(L)=d!\operatorname{Vol}(\Delta_{\nu}(Y,L))$ [16, Theorem 1.15], where the left hand side is the volume of the line bundle $L$ and the right hand side is the volume of the polytope $\Delta_{\nu}(Y,L)$.
+Let $\big{\{}V_{\bullet}^{t}:t\in\mathbb{R}\big{\}}$ be a decreasing $\mathbb{R}$-filtration of $V_{\bullet}$ that is multiplicative, i.e.
+$V_{n_{1}}^{t_{1}}\cdot V_{n_{2}}^{t_{2}}\subseteq V_{n_{1}+n_{2}}^{t_{1}+t_{2}}$.
+The Boucksom-Chen concave transform associated to $\left(V_{\bullet},\nu,\big{\{}V_{\bullet}^{t}\big{\}}\right)$ is the function $G:\Delta_{\nu}(V_{\bullet})\longrightarrow\mathbb{R}\cup\{-\infty\}$, $x\longmapsto\sup\big{\{}t\in\mathbb{R}:x\in\Delta_{\nu}(V_{\bullet}^{t})\big{\}}$ [6].
+$G$ is concave, as indicated in the name.
+In particular, $G$ is continuous in the interior.
+3.2. String polytopes and weight polytopes
+Let $w_{0}\in W$ be the longest Weyl element and fix a reduced decomposition $\underline{w_{0}}=s_{i_{1}}s_{i_{2}}\cdots s_{i_{N}}$, $N=\dim G/B$.
+In [19, 3], the authors constructed the string valuation $\nu=\nu_{\underline{w_{0}}}:H^{0}(G/P,nM_{\lambda})\backslash\{0\}\longrightarrow\mathbb{Z}^{N}$.
+The image $\nu(H^{0}(G/P,nM_{\lambda})\backslash\{0\})=\big{\{}v_{1},\dots,v_{l}\big{\}}$ is in bijection (the string parametrization) with the so-called dual crystal basis $\big{\{}e_{v_{1}}^{(n\lambda)},\cdots,e_{v_{l}}^{(n\lambda)}\big{\}}$ of $H^{0}(G/P,nM_{\lambda})$ (see [16, §3] for a detailed exposition).
+Moreover, the dual crystal basis is compatible with the weight decomposition
+$H^{0}(G/P,nM_{\lambda})=\bigoplus_{\mu\in X(T)}H^{0}(G/P,nM_{\lambda})[\mu],$
+that is, for each $i=1,\dots,l$, $e_{v_{i}}^{(n\lambda)}\in H^{0}(G/P,nM_{\lambda})[\mu]$ for some $\mu$.
+The string polytope $\Delta_{\lambda}$ is the Okounkov body $\Delta_{\nu}(G/P,M_{\lambda})$, i.e.
+it is the closure of the set $\bigcup_{n\geq 1}\tfrac{1}{n}\nu(H^{0}(G/P,nM_{\lambda})\backslash\{0\})$.
+Note that $\Delta_{\lambda}\cap\tfrac{1}{n}\mathbb{Z}^{N}$ is in bijection with the dual crystal basis of $H^{0}(G/P,nM_{\lambda})$.
+The weight polytope $\mathcal{P}_{\lambda}\subseteq X(T)_{\mathbb{R}}$ is the convex hull of the Weyl orbit $W\lambda$ in $X(T)_{\mathbb{R}}$.
+Note that $\mathcal{P}_{\lambda}\cap\tfrac{1}{n}X(T)$ is in bijection with weights apprearing in $H^{0}(G/P,nM_{\lambda})$.
+By taking weights, one has weight map $p:\Delta_{\lambda}\longrightarrow\mathcal{P}_{\lambda}$ which is affine and sends $\Delta_{\lambda}\cap\tfrac{1}{n}\mathbb{Z}^{N}$ to $\mathcal{P}_{\lambda}\cap\tfrac{1}{n}X(T)$
+3.3. Proof of Theorem 3.1
+Let $V_{n}=H^{0}(G/P,nM_{\lambda})$ and $V_{\bullet}=\bigoplus_{n\geq 0}V_{n}$.
+For $t\in\mathbb{R}$, let $V_{n}^{t}=H^{0}(G/P,nM_{\lambda})_{t,{\mathrm{deg}}(F_{Q})}$ as in Proposition 2.12, and $V_{\bullet}^{t}=\bigoplus_{n\geq 0}V_{n}^{nt}$.
+Let $\nu:V_{\bullet}\backslash\{0\}\longrightarrow\mathbb{Z}^{N}$ be the string valuation.
+Let $\varphi:\Delta_{\nu}(G/P,M_{\lambda})=\Delta_{\lambda}\longrightarrow\mathbb{R}$ be the Boucksom-Chen concave transform associated to $\left(V_{\bullet},\nu,\big{\{}V_{\bullet}^{t}\big{\}}\right)$.
+Theorem 3.3.
+$\varphi(x)=\langle{\mathrm{deg}}(F_{Q}),p(x)\rangle$.
 Proof.
-By Ballaÿ’s theorem [2, Theorem 1.2],
-$\quad\quad\quad\zeta_{1}(h_{\mathcal{L}_{\lambda}},X_{w})$ $=\displaystyle\lim_{n\rightarrow\infty}\frac{\mu_{\max}(\pi_{*}\mathcal{L}_{n\lambda}|_{\mathcal{X}_{w}})}{n}$
-$=\lim_{n\rightarrow\infty}\frac{n\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle}{n}=\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle.$
-Theorem 2.11.
-The height filtration of $h_{\mathcal{L}_{\lambda}}$ on $X$ is given by successively deleting Schubert cells $C_{w}$ for $w\in W_{Q}\backslash W/W_{P}$, i.e.
-$Z_{t}=X\Bigg{\backslash}\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\geq t}C_{w}=\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle<t}C_{w}.$
-In particular, successive minima are $\zeta_{w}=\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ and Zhang’s successive minima are $e_{i}=\min\big{\{}\zeta_{w}:\ell(w)=d-i+1\big{\}}$.
-Proof.
-On $X_{w}$, $\zeta_{1}(h_{\mathcal{L}_{\lambda}},X_{w})=\zeta_{w}$ and for any $x\in C_{w}$, $h(x)\geq\zeta_{w}$.
-Thus $Z_{\zeta_{w}}(X_{w})\subseteq X_{w}\backslash C_{w}=\bigcup_{w^{\prime}<w}X_{w^{\prime}}$.
-For a closed subvariety $Y$ of $X$, let $Z_{t}(Y)$ be the Zariski closure in $Y$ of $\big{\{}y\in Y:h(y)<t\big{\}}$.
-We claim that $Z_{\zeta_{w}}(X_{w})=X_{w}\backslash C_{w}=\bigcup_{w^{\prime}<w}X_{w^{\prime}}$.
-In fact, suppose conversely that $Z_{\zeta_{w}}(X_{w})\subsetneq X_{w}\backslash C_{w}$.
-Then there exists $X_{w^{\prime}}\subsetneq X_{w}$ such that $Z_{\zeta_{w}}(X_{w})\cap X_{w^{\prime}}\subsetneq X_{w^{\prime}}$.
-This forces $Z_{\zeta_{w}}(X_{w^{\prime}})\subseteq Z_{\zeta_{w}}(X_{w})\cap X_{w^{\prime}}\subsetneq X_{w^{\prime}}$.
-But on $X_{w^{\prime}}$, the essential minimum is $\zeta_{w^{\prime}}<\zeta_{w}$, so $Z_{\zeta_{w}}(X_{w^{\prime}})=X_{w^{\prime}}$.
-We get contradiction and we have thus proved that on $X_{w}$, the set $Z_{t}(X_{w})=X_{w}$ when $t>\zeta_{w}$ and $Z_{t}(X_{w})=X_{w}\backslash C_{w}$ when $t\leq\zeta_{w}$.
-Let $w_{0}\in W_{Q}\backslash W/W_{P}$ be the longest element.
-On $X=X_{w_{0}}$, when $t>\zeta_{w_{0}}$, $Z_{t}=X_{w_{0}}$ and when $t\leq\zeta_{w_{0}}$, $Z_{t}=X_{w_{0}}\backslash C_{w_{0}}=\bigcup_{w^{\prime}}X_{w^{\prime}}$ with $w^{\prime}<w_{0}$ and $\ell(w^{\prime})=\ell(w_{0})-1$.
-Repeating this procedure on each smaller $X_{w^{\prime}}$, we can complete the proof by induction.
-∎
-3. Reinforcement of Zhang’s inequality into an equality
-In this section, we shall establish the following Zhang-type equality on flag varieties over function fields: the height of the variety in fact equals to a weighted average of the successive minima.
-Theorem 3.1.
-Let $q:W/W_{P}\longrightarrow W_{Q}\backslash W/W_{P}$ be the canoniacl projection. For $w\in W_{Q}\backslash W/W_{P}$, define $m_{w}=\#q^{-1}(w)$. Then
-$h_{\mathcal{L}_{\lambda}}(X)=\frac{1}{\#W/W_{P}}\sum_{w\in W_{Q}\backslash W/W_{P}}m_{w}\zeta_{w}.$
-We shall establish this theorem by convex geometry.
-Let $E$ be a vector bundle on $C$ with HN filtration $0=E_{0}\subseteq E_{1}\subseteq\cdots\subseteq E_{r}=E$ with successive slopes $\mu_{i}=\mu(E_{i}/E_{i-1})$.
-Chen defines a measure $\nu_{E}$ on $\mathbb{R}$ by
-$\nu_{E}=\frac{1}{\operatorname{rank}(E)}\sum_{i=1}^{r}\operatorname{rank}(E_{i}/E_{i-1})\delta_{\mu_{i}},$
-where $\delta_{a}$ is the Dirac measure on $\mathbb{R}$ supported at $a\in\mathbb{R}$.
-For any $\varepsilon>0$ and any measure $\mu$, let $T_{\varepsilon}\mu$ be the unique measure such that $\int_{\mathbb{R}}f(x)\mathrm{d}T_{\varepsilon}\mu(x)=\int_{\mathbb{R}}f(\varepsilon x)\mathrm{d}\mu(x)$ for any compactly-supported continuous function $f:\mathbb{R}\longrightarrow\mathbb{R}$.
-Theorem 3.2 (Chen; [11], Theorem 1.1).
-Let $\pi:\mathcal{X}\longrightarrow C$ be a projective flat morphism and $\mathcal{L}$ be a line bundle on $\mathcal{X}$. Assume $\mathcal{L}$ is generically big. Then
-  1. (1)
-    $\big{\{}T_{\frac{1}{n}}\nu_{\pi_{*}\mathcal{L}^{n}}\big{\}}_{n}$ converges weakly. We denote the limit measure by $\nu^{\pi}_{\mathcal{L}}$.
-  2. (2)
-    $\operatorname{Vol}(\mathcal{L})=\dim\mathcal{X}\cdot\operatorname{Vol}(\mathcal{L}_{K})\cdot\int_{\mathbb{R}}x_{+}\;\mathrm{d}\nu^{\pi}_{\mathcal{L}}$, where $x_{+}=\max\{0,x\}$.
-Let $w_{0}\in W$ be the longest element and fix a reduced decomposition $w_{0}=s_{1}\cdot s_{2}\cdots s_{{\ell(w_{0})}}$.
-Recall that $\lambda:P\longrightarrow\mathbb{G}_{m}$ is a strictly anti-dominant character and $N=\dim G/P$.
-Let $\Delta_{\lambda}=\Delta_{\underline{w}_{0},\lambda}\subseteq\mathbb{R}^{N}$ be the string polytope associated to $\lambda$ with respect to $\underline{w}_{0}=(s_{1},\cdots,s_{\ell(w_{0})})$.
-It is equipped with a weight map $p:\Delta_{\lambda}\longrightarrow\mathcal{P}_{\lambda}$ where $\mathcal{P}_{\lambda}\subseteq X(T)_{\mathbb{R}}$ is the weight polytope associated to $\lambda$ (see [1, §1] for details).
-The following properties will be useful:
-  1. (1)
-    The volume $\operatorname{Vol}(\Delta_{\lambda})$ of the convex polytope equals to the volume $\operatorname{Vol}(M_{\lambda})$ of the line bundle.
-  2. (2)
-    Integral points $\Delta_{\lambda}\cap\mathbb{Z}^{N}$ parametrize a basis of $H^{0}(G/P,M_{\lambda})$. In particular, #$\#\Delta_{\lambda}\cap\mathbb{Z}^{N}=\dim H^{0}(G/P,M_{\lambda})$.
-  3. (3)
-    Lattice points $\mathcal{P}_{\lambda}\cap X(T)$ parameterize weights occuring in $H^{0}(G/P,M_{\lambda})$ and $p$ sends an integral point to the underlying weight of the corresponding basis vector.
-  4. (4)
-    $\Delta_{\lambda_{1}}+\Delta_{\lambda_{2}}=\Delta_{\lambda_{1}+\lambda_{2}}$ and $p_{\lambda_{1}}+p_{\lambda_{2}}=p_{\lambda_{1}+\lambda_{2}}$.
-Here the sum is the Minkowski sum.
-In particular, the diagram
-    ${\Delta_{\lambda}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p_{\lambda}}$$\scriptstyle{\cdot n}$${\Delta_{n\lambda}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p_{n\lambda}}$${\mathcal{P}_{\lambda}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\cdot n}$${\mathcal{P}_{n\lambda}}$
-    is commutative for any $n\in\mathbb{N}$, where $\cdot n$ is the scalar multiplication by $n$.
-Let $\mu$ be the Lebesgue measure on $\mathbb{R}^{N}$ normalized with respect to $\mathbb{Z}^{N}$.
-Lemma 3.3.
-The following equalities hold:
-$\nu_{\pi_{*}\mathcal{L}^{n}_{\lambda}}=\frac{1}{\#(\Delta_{n\lambda}\cap\mathbb{Z}^{N})}\sum_{x\in\Delta_{n\lambda}\cap\mathbb{Z}^{N}}\delta_{{\mathrm{deg}}F_{Q}\circ p(x)},$
-$T_{\frac{1}{n}}\nu_{\pi_{*}\mathcal{L}_{\lambda}^{n}}=\frac{1}{\#(\Delta_{\lambda}\cap\frac{1}{n}\mathbb{Z}^{N})}\sum_{x\in\Delta_{\lambda}\cap\frac{1}{n}\mathbb{Z}^{N}}\delta_{{\mathrm{deg}}F_{Q}\circ p(x)},$
-$\nu^{\pi}_{\mathcal{L}_{\lambda}}=\frac{1}{\operatorname{Vol}(\Delta_{\lambda})}\cdot({\mathrm{deg}}F_{Q}\circ p)_{*}\mu.$
-Proof.
-Bearing Lemma 2.8, Proposition 2.9 and the aforementioned properties of string polytopes at hand, we can carry over the argument in [11, Proposition 3.5] to show the first equality.
-Appling $T_{1/n}$ and taking limits, we obtain the second and third equalities.
-∎
-Theorem 3.4.
-The variety height $h_{\mathcal{L}_{\lambda}}(X)$ of $X$ is $\frac{1}{\operatorname{Vol}(\Delta_{\lambda})}\int_{\Delta_{\lambda}}{\mathrm{deg}}(F_{Q})\circ p\;\mathrm{d}\mu$.
-Proof.
-We may consider $\mathcal{L}_{\lambda}+tf$ for large $t$ so that $\mathcal{L}_{\lambda}+tf$ is ample, where $f$ is the line bundle given by a vertical fiber.
-The measure $\mu^{\pi}_{\mathcal{L}_{\lambda}+tf}$ is supported on $\mathbb{R}_{+}$ and is a right shift of $\mu^{\pi}_{\mathcal{L}_{\lambda}}$ by $t$, so Theorem 3.2 says
-$\quad\quad\quad(\mathcal{L}_{\lambda}+tf)^{N+1}$ $=(N+1)\operatorname{Vol}(\mathcal{L}_{\lambda,K})\int_{\mathbb{R}}x\;\mathrm{d}\nu^{\pi}_{\mathcal{L}_{\lambda}+tf}$
-$=(N+1)\int_{\Delta_{\lambda}}{\mathrm{deg}}(F_{Q})\circ p+t\;\mathrm{d}\mu$
-where the second equality is due to Lemma 3.3 and $\operatorname{Vol}(\mathcal{L}_{\lambda,K})=\operatorname{Vol}(\Delta_{\lambda})$. Thus
-$h_{\mathcal{L}_{\lambda}+tf}(X)=\frac{1}{\operatorname{Vol}(\Delta_{\lambda})}\int_{\Delta_{\lambda}}{\mathrm{deg}}(F_{Q})\circ p\;\mathrm{d}\mu+t.$
-Note that $h_{\mathcal{L}_{\lambda}+tf}(X)=h_{\mathcal{L}_{\lambda}}(X)+t$. So substracting $t$, we get
-$h_{\mathcal{L}_{\lambda}}(X)=\frac{1}{\operatorname{Vol}(\Delta_{\lambda})}\int_{\Delta_{\lambda}}{\mathrm{deg}}(F_{Q})\circ p\;\mathrm{d}\mu.$
-To prove Theorem 3.1, we need the following general result.
+$V_{n}^{nt}=\bigoplus_{\mu}V_{n}[n\mu]$ for $\mu\in\mathcal{P}_{\lambda}\cap\tfrac{1}{n}X(T)$ with $\langle{\mathrm{deg}}(F_{Q}),n\mu\rangle\geq nt$, and hence $V_{n}^{nt}=\bigoplus_{x}k\cdot e_{nx}^{(n\lambda)}$ for $x\in\Delta_{\lambda}\cap\tfrac{1}{n}\mathbb{Z}^{N}$ with $\langle{\mathrm{deg}}(F_{Q}),p(nx)\rangle\geq nt$.
+Therefore by definition,
+$\quad\quad\quad\Delta(V_{\bullet}^{t})$ $=\text{closure of }\bigcup\nolimits_{n}\big{\{}x\in\Delta_{\lambda}\cap\tfrac{1}{n}\mathbb{Z}^{N}:\langle{\mathrm{deg}}(F_{Q}),p(x)\rangle\geq t\big{\}}$
+=$=\big{\{}x\in\Delta_{\lambda}:\langle{\mathrm{deg}}(F_{Q}),p(x)\rangle\geq t\big{\}}$
+which concludes the proof. ∎
+Consider the graded linear series $\widetilde{V}_{\bullet}=\bigoplus_{n\geq 0}H^{0}(X,nL)$. Fix an isomorphism $F_{Q,K}\simeq Q_{K}$. Then
+(3.1) $\widetilde{V}_{\bullet}=\bigoplus_{n\geq 0}\Big{(}F_{Q}\times_{Q}H^{0}(G/P,nM_{\lambda})\Big{)}_{K}\simeq\bigoplus_{n\geq 1}H^{0}(G/P,nM_{\lambda})_{K}.$
+Consider the filtrations of $\widetilde{V}_{n}=\Big{(}F_{Q}\times_{Q}H^{0}(G/P,nM_{\lambda})\Big{)}_{K}$ induced by the Harder-Narasimhan filtration of $\pi_{*}\mathcal{L}_{n\lambda}=F_{Q}\times_{Q}H^{0}(G/P,nM_{\lambda})$ as in Proposition 2.12
+$\widetilde{V}_{\bullet}^{t}=\bigoplus_{n\geq 0}\Big{(}F_{Q}\times_{Q}H^{0}(G/P,nM_{\lambda})_{nt,{\mathrm{deg}}(F_{Q})}\Big{)}_{K}\simeq\bigoplus_{n\geq 0}\Big{(}H^{0}(G/P,nM_{\lambda})_{nt,{\mathrm{deg}}(F_{Q})}\Big{)}_{K}.$
+Via the string parametrization of the dual crystal basis, the string valuation $\nu$ on $H^{0}(G/P,nM_{\lambda})$ extends canonically to a valuation $\widetilde{\nu}$ on $H^{0}(G/P,nM_{\lambda})_{K}$.
+This induces a valuation on $\widetilde{V}_{n}\simeq H^{0}(G/P,nM_{\lambda})_{K}$ via the isomorphism (3.1).
+We shall abuse notation and also call it $\widetilde{\nu}$.
+The Boucksom-Chen concave transform of ${\mathcal{L}_{\lambda}}$ with respect to $\widetilde{\nu}$ is the one associated to $\left(\widetilde{V}_{\bullet},\widetilde{\nu},\big{\{}\widetilde{V}_{\bullet}^{t}\big{\}}\right)$, and obviously it coincides with the one associated to $\left(V_{\bullet},\nu,\big{\{}V_{\bullet}^{t}\big{\}}\right)$.
+Theorem 3.1 follows.
+3.4. Proof of Theorem 3.2
+Let $\mu$ be the Lebesgue measure on the real span of $\Delta_{\lambda}$, normalized with respect to $\mathbb{Z}^{N}$. By [6], we have
+$h_{\mathcal{L}_{\lambda}}(X)=\frac{\mathcal{L}_{\lambda}^{\dim\mathcal{X}}}{\dim\mathcal{X}\cdot L^{\dim X}}=\frac{1}{\operatorname{Vol}(\Delta_{\lambda})}\int_{\Delta_{\lambda}}\langle{\mathrm{deg}}(F_{Q}),p(x)\rangle\,\mathrm{d}\mu(x).$
 Proposition 3.4.
-For any linear function $f:X(T)_{\mathbb{R}}\longrightarrow\mathbb{R}$,
+For any linear function $f:X(T)_{\mathbb{R}}\longrightarrow\mathbb{R}$, we have
 $\frac{1}{\operatorname{Vol}(\Delta_{\lambda})}\int_{\Delta_{\lambda}}f\circ p\;\mathrm{d}\mu=\frac{1}{\#W/W_{P}}\sum_{w\in W/W_{P}}f(w\lambda).$
-We start with two lemmata.
+Proof of Theorem 3.2.
+Take $f={\mathrm{deg}}(F_{Q})$ in Proposition 3.4. ∎
+The rest of this subsection is devoted to the proof of Proposition 3.4. We start with two lemmata.
 Lemma 3.5.
@@
 Proof.
-Note that the Lebesgue measure $\mu$ on $\Delta_{\lambda}$ is the limit of the Dirac measure $\delta_{\Delta_{\lambda}\cap\frac{1}{n}\mathbb{Z}^{N}}$, i.e.
+Note that the measure $\mu$ on $\Delta_{\lambda}$ is the limit of the Dirac measures $\delta_{\Delta_{\lambda}\cap\frac{1}{n}\mathbb{Z}^{N}}$, i.e.
 $\mu=\lim_{n}\delta_{\Delta_{\lambda}\cap\frac{1}{n}\mathbb{Z}^{N}}$.
@@
 where $\operatorname{mult}_{n}$ is the multiplicity function
-$\operatorname{mult}_{n}:\mathcal{P}_{\lambda}\cap\tfrac{1}{n}X(T)\longrightarrow\mathbb{N},\quad\mu\longmapsto\dim H^{0}(G/P,L_{n\lambda})[n\mu].$
+$\operatorname{mult}_{n}:\mathcal{P}_{\lambda}\cap\tfrac{1}{n}X(T)\longrightarrow\mathbb{N},\quad\mu\longmapsto\dim H^{0}(G/P,M_{n\lambda})[n\mu].$
 Since both $\operatorname{mult}_{n}$ and $\delta_{\mathcal{P}_{\lambda}\cap\tfrac{1}{n}X(T)}$ are $W$-invariant, we deduce that
@@
 In this case, $T^{W}=Z(G)$ and the desired equality follows directly.
-$H^{0}(G/P,L_{\lambda})$ is a simple $G$-module so all weights $\mu$ occuring in $H^{0}(G/P,L_{\lambda})$ sharing the same restriction to $Z(G)$.
+$H^{0}(G/P,M_{\lambda})$ is a simple $G$-module so all weights $\mu$ occuring in $H^{0}(G/P,M_{\lambda})$ share a same restriction to $Z(G)$.
 So by the above claim, the lemma is true for lattice points $\mathcal{P}_{\lambda}\cap X(T)$.
@@
 Taking ratios, we obtain the desired equality. ∎
-Proof of Theorem 3.1.
-Take $f={\mathrm{deg}}(F_{Q})$ in Proposition 3.5 and apply Theorem 3.4. ∎
-Remark 3.8.
-The fact that the variety height and successive minima are given by
-$\displaystyle h_{\mathcal{L}_{\lambda}}(X)=\frac{1}{\operatorname{Vol(\Delta_{\lambda})}}\int_{\mathcal{P}_{\lambda}}{\mathrm{deg}}(F_{Q})\;\mathrm{d}p_{*}\mu,\quad\zeta_{w}=\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$
-shows that the Arakelov geometry of $(F/P,\mathcal{L}_{\lambda})$ depends only on the function ${\mathrm{deg}}(F_{Q})$ on $X(T)_{\mathbb{Q}}$, which might be surprising at the first glance.
-This is in fact because the Arakelov geometry of $(F/P,\mathcal{L}_{\lambda})$ is determined by the Harder-Narasimhan filtration of $\pi_{*}\mathcal{L}_{n\lambda}=F_{Q}\times_{Q}H^{0}(G/P,M_{n\lambda})$ (Lemma 2.8) for all $n\geq 1$, and the unipotent part of $Q$ acts trivially [23, Lemma 5.1].
+We end this section by the case $\operatorname{Gr}(2,4)$.
 Example 3.7.
 Let $G=\operatorname{GL}_{4}$, $E$ be a vector bundle on $C$ of rank $4$ and consider the $\operatorname{Gr}(2,4)$-bundle $\operatorname{Gr}_{2}(E)$.
-In this case the string polytopes (with respect to a certain reduced decomposition) is just the Gelfand-Zetlin polytope [18, Remark 2.4].
+In this case the string polytopes (with respect to a certain reduced decomposition) is just the Gelfand-Zetlin polytope [15, Remark 2.4].
 Identify $X(T)_{\mathbb{Q}}$ with $\bigoplus_{i}\mathbb{Q}\lambda_{i}$ as before.
-The line bundle $\mathcal{O}(1)$ correspond to the character $\lambda=(0,0,1,1)$.
+The line bundle $\mathcal{O}(1)$ corresponds to the character $\lambda=(0,0,1,1)$.
 The Gelfand-Zetlin polytope $\operatorname{GZ}_{\lambda}$ is defined by $0\leq x_{2}\leq x_{1}\leq x_{3}\leq 1$ and $x_{2}\leq x_{4}\leq x_{3}$ in $\mathbb{R}^{4}$, with vertices $(0,0,0,0)$, $(0,0,1,0)$, $(0,0,1,1)$, $(1,0,1,0)$, $(1,0,1,1)$ and $(1,1,1,1)$.
@@
 $\sum_{w\in W/W_{P}}\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle=\sum_{1\leq i<j\leq 4}\mu_{i}+\mu_{j}.$
-We see that $\frac{1}{\#W/W_{P}}\sum_{w\in W/W_{P}}\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ is also $\frac{1}{2}(\mu_{1}+\mu_{2}+\mu_{3}+\mu_{4})$.
+Thus $\frac{1}{\#W/W_{P}}\sum_{w\in W/W_{P}}\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle=\frac{1}{2}(\mu_{1}+\mu_{2}+\mu_{3}+\mu_{4})=h(X)$, which is predicted by Theorem 3.2.
 4. Base loci and movable cones
-4.1. Preliminaries on algebraic geometry
-We briefly recall some general theory of base loci.
-Let $\mathbf{k}$ be a field.
-In our case, it would be either the base field $k$ or the function field $K=K(C)$.
-Let $Y\rightarrow\operatorname{Spec}\mathbf{k}$ be a normal projective variety over $\mathbf{k}$, and $L$ be a line bundle over $X$.
-Let $V$ be a $\mathbf{k}$-subspace of $H^{0}(Y,L)$.
-Consider the evaluation map
-$H^{0}(Y,L)\otimes_{\mathbf{k}}L^{\vee}\rightarrow\mathcal{O}_{Y}.$
-The base locus $\mathrm{Bs}(V)$ of $V$ is the reduced scheme associated to the closed subscheme defined by the ideal sheaf given by the image of $\pi^{*}V\otimes{L^{\vee}}\rightarrow\mathcal{O}_{Y}$.
-Definition 4.1.
+Keep the notations in §2.
+We fix a closed point $p_{0}\in C$ and set $f=\pi^{*}(\mathcal{O}_{C}(p_{0}))$.
+For any $w\in W_{Q}\backslash W/W_{P}$, let $\mathcal{C}_{w}=F_{Q}\times_{Q}QwP/P$.
+Let $t\in\mathbb{Q}$.
+The main theorem of this section is a computation of base loci.
+Theorem 4.1.
+Let $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$ be the augmented base locus of $\mathcal{L}_{\lambda}-tf$. Then
+$\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)=\mathcal{X}\Bigg{\backslash}\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle>t}\mathcal{C}_{w}=\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\leq t}\mathcal{C}_{w}.$
+Comparing Theorem 2.1 and Theorem 4.1, we find that $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$ agrees with the Zariski closure of $Z_{t}$ in $\mathcal{X}$ for every $t\not=\zeta_{w}$.
+At critical points, $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$ is upper continuous while $Z_{t}$ is lower continuous.
+As a corollary, we compute the movable cones on flag bundles. (Please see §4.3 for the notations.)
+Theorem 4.2.
+The $k$-th movable cone $\operatorname{Mov}^{k}(F/P)$ is the cone defined by
+  1. (1)
+    $\langle\alpha^{\vee},\cdot\rangle<0$ for any $\alpha\in\Delta\backslash\Delta_{P}$, and
+  2. (2)
+    $\langle{\mathrm{deg}}(F_{Q}),w\cdot\rangle>0$ for any $w\in W/W_{P}$ with $\ell(w)\geq n-k+1$.
+In particular, let $\mathcal{L}$ be a line bundle on $F/P$ and $e_{k}(h_{\mathcal{L}})$ be the $k$-th minimum of Zhang of the height function $h_{\mathcal{L}}$, we have
+Corollary 4.3.
+$\mathcal{L}$ is $k$-movable if and only if $e_{k}(h_{\mathcal{L}})>0$.
+4.1. Augmented base loci and the restricted volume
+Let $\mathbf{k}$ be a field and $Y$ be a normal projective variety over $\mathbf{k}$.
+Let $L$ be a line bundle on $Y$ and $V\subseteq H^{0}(Y,L)$ be a subspace.
+We shall denote by $\mathrm{Bs}(V)$ the base locus of $V$.
 The stable base locus of $L$ is defined as $\mathrm{B}(L)=\bigcap_{n\geq 1}\mathrm{Bs}(H^{0}(Y,nL))$.
-Take any ample line bundle $A$.
-The augmented base locus of $L$ is defined as $\mathrm{B}_{+}(L)=\bigcap_{n\geq 1}\mathrm{B}(nL-A)$, which is independent of the choice of $A$.
-The following characterization of augmented base loci in terms of restricted volumes is well known, which we include here for the reader’s convenience.
+The augmented base locus of $L$ is defined as $\mathrm{B}_{+}(L)=\bigcap_{n\geq 1}\mathrm{B}(nL-A)$ for any ample line bundle $A$.
+This is independent of the choice of $A$.
+Let $Z$ be a closed subvariety of $Y$ of dimension $d$. The restricted volume is defined as
+$\operatorname{Vol}(Y|Z,L)=\limsup_{n\rightarrow\infty}\frac{\dim_{\mathbf{k}}\mathrm{Im}\Big{(}H^{0}(Y,nL)\rightarrow H^{0}(Z,nL|_{Z})\Big{)}}{n^{d}/d!}.$
+The following characterization of augmented base loci in terms of restricted volumes is well known, which we include here for readers’ convenience.
 Theorem 4.4 ([11], Theorem C).
-Let $Z$ be a closed subvariety of $Y$. The restricted volume is defined as
-$\operatorname{Vol}(Y|Z,L)=\limsup_{n\rightarrow+\infty}\frac{\dim_{\mathbf{k}}\mathrm{Im}(H^{0}(Y,nL)\rightarrow H^{0}(Z,nL|_{Z}))}{n^{\dim Z}/(\dim Z)!}$
-where Im()normal-$\mathrm{Im}(\cdot)$ denotes the image. Then $\mathrm{B}_{+}(L)=\bigcup_{\operatorname{Vol}(Y|Z,L)=0}Z$.
-Let $\mathrm{Pic}(Y)$ be the picard group of $Y$.
-Since $\mathrm{B}(L)=\mathrm{B}(mL)$ for any $m>1$, we can extend the definitions of $\mathrm{B}(\cdot)$ and $\mathrm{B}_{+}(\cdot)$ to $\mathrm{Pic}(Y)_{\mathbb{Q}}:=\mathrm{Pic}(Y)\otimes\mathbb{Q}$.
-Moreover, let $N^{1}(Y):=\mathrm{Div}(X)/\equiv_{num}$ be the numerical group of divisors.
-It is proved in [13] that the augmented base loci are well defined on $N^{1}(Y)_{\mathbb{R}}:=N^{1}(Y)\otimes_{\mathbb{Z}}\mathbb{R}$.
-Let $\mathcal{Y}$ be a projective variety with a surjective morphism $\mathcal{Y}\rightarrow C$, and $\mathcal{L}$ be a line bundle over $\mathcal{Y}$.
-For any scheme-theoretical point $p\in C$, we denote by $\kappa(p)$ the residue field of $p$.
-Let $\mathcal{Z}$ be a subvariety of $\mathcal{Y}$.
-The fiber $\mathcal{Z}\times_{C}\mathrm{Spec}(\kappa(p))$ of $\mathcal{Z}$ over $p$, is denoted as $\mathcal{Z}_{p}$.
-We also denote by $\mathcal{L}_{p}$ the restriction $\mathcal{L}|_{\mathcal{Y}_{p}}$ of $\mathcal{L}$ to $\mathcal{Y}_{p}$.
-We have the following lemma, which will be useful for our computation of augmented base loci in the next subsection.
+$\mathrm{B}_{+}(L)$ is the union of all closed subvarieties $Z$ with $\operatorname{Vol}(Y|Z,L)=0$.
+Let $N^{1}(Y)$ be the Néron-Severi group of $Y$. It is proved in [10] that augmented base loci are well defined on $N^{1}(Y)_{\mathbb{R}}:=N^{1}(Y)\otimes_{\mathbb{Z}}\mathbb{R}$.
+Let $\mathcal{Y}\longrightarrow C$ be a projective flat family over the curve $C$, and $\mathcal{L}$ be a line bundle over $\mathcal{Y}$.
+For any scheme theoretical point $p\in C$, we denote by $\kappa(p)$ the residue field of $p$.
+Let $\mathcal{Z}$ be a closed subvariety of $\mathcal{Y}$, we denote by $\mathcal{Z}_{p}$ the fiber $\mathcal{Z}\times_{C}\mathrm{Spec}(\kappa(p))$ of $\mathcal{Z}$ over $p$.
+We denote by $\mathcal{L}_{p}$ the restriction $\mathcal{L}|_{\mathcal{Y}_{p}}$ of $\mathcal{L}$ to $\mathcal{Y}_{p}$.
 Lemma 4.5.
-The restriction of $\mathrm{B}(\mathcal{L})$ to $\mathcal{Y}_{p}$ can be computed by
-$B(\mathcal{L})_{p}=\bigcap\nolimits_{n\geq 1}\mathrm{Bs}(\mathrm{Im}(H^{0}(\mathcal{X},n\mathcal{L})\otimes_{k}\kappa(p)\rightarrow H^{0}(\mathcal{Y}_{p},n\mathcal{L}_{p}))).$
-4.2. Augmented base loci on the flag bundle $\mathcal{X}$
-Let $\lambda\in X(P)$ be a strictly anti-dominant character.
-We fix a closed point $p_{0}\in C$, and set $f=\pi^{*}(\mathcal{O}_{C}(p_{0}))$.
-In this subsection, we are going to compute the augmented base locus $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$.
-Remind that we have the decomposition
-$H^{0}(G/P,M_{\lambda})=\bigoplus_{\nu\in X(T)}H^{0}(G/P,M_{\lambda})[\nu]$
+$\mathrm{B}(\mathcal{L})_{p}=\bigcap\nolimits_{n\geq 1}\mathrm{Bs}\left(\mathrm{Im}\Big{(}H^{0}(\mathcal{Y},n\mathcal{L})\otimes_{k}\kappa(p)\rightarrow H^{0}(\mathcal{Y}_{p},n\mathcal{L}_{p})\Big{)}\right)$.
+4.2. Proof of Theorem 4.1
+Let $M_{\lambda}=G\times_{P}k_{\lambda}$, which is an ample line bundle on $G/P$. Recall that we have the decomposition
+$H^{0}(G/P,M_{\lambda})=\bigoplus_{\mu\in X(T)}H^{0}(G/P,M_{\lambda})[\mu]$
 and the filtration
-$H^{0}(G/P,M_{\lambda})_{t,{\mathrm{deg}}(F_{Q})}=\bigoplus_{\langle{\mathrm{deg}}(F_{Q}),\nu\rangle\geq t}H^{0}(G/P,M_{\lambda})[\nu].$
-We start with a lemma describing the vanishing behaviour of sections in $H^{0}(G/P,M_{\lambda})[\nu]$.
+$H^{0}(G/P,M_{\lambda})_{t,{\mathrm{deg}}(F_{Q})}=\bigoplus_{\langle{\mathrm{deg}}(F_{Q}),\mu\rangle\geq t}H^{0}(G/P,M_{\lambda})[\mu].$
+We start with a lemma describing the vanishing behaviour of sections in $H^{0}(G/P,M_{\lambda})[\mu]$.
 Lemma 4.6.
@@
   2. (2)
-    $s\in H^{0}(G/P,M_{\lambda})[\nu]$ for $\nu\not\leq w\lambda$Longrightarrow $s(x)=0$ for any $x\in\overline{BwP}/P$.
+    $s\in H^{0}(G/P,M_{\lambda})[\mu]$ for $\mu\not\leq w\lambda$Longrightarrow $s(x)=0$ for any $x\in\overline{BwP}/P$.
 Proof.
@@
 Consequently, $\mathrm{Res}(s)(x)\neq 0$ for any $x\in BwP/P$.
-As for $(2)$, for any $\nu\not\leq w\lambda$, $H^{0}(\overline{BwP}/P,M_{\lambda})[\nu]=0$. Consequently $\mathrm{Res}(s)=0$ for any $s\in H^{0}(G/P,M_{\lambda})[\nu]$.
+As for $(2)$, for any $\mu\not\leq w\lambda$, $H^{0}(\overline{BwP}/P,M_{\lambda})[\mu]=0$. Consequently, $\mathrm{Res}(s)=0$ for any $s\in H^{0}(G/P,M_{\lambda})[\mu]$. ∎
+Let $p$ be a scheme theoretical point of $C$ and consider the filtration $H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})_{\bullet,{\mathrm{deg}}(F_{Q})}$ of $H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})$ given by
+$H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})_{t,{\mathrm{deg}}(F_{Q})}=F_{Q,\kappa(p)}\times_{Q}H^{0}(G/P,M_{\lambda})_{t,{\mathrm{deg}}(F_{Q})}.$
+We have
 Lemma 4.7.
-Let $p$ be a scheme-theoretical point of $C$.
-Consider the filtration $H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})_{\bullet,{\mathrm{deg}}(F_{Q})}$ of $H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})$ given by
-$H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})_{t,{\mathrm{deg}}(F_{Q})}=F_{Q,\kappa(p)}\times_{Q}H^{0}(G/P,M_{\lambda})_{t,{\mathrm{deg}}(F_{Q})}$
-for $t\in\mathbb{R}$, we have $\operatorname{Bs}\Big{(}H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})_{t,{\mathrm{deg}}(F_{Q})}\Big{)}=\mathcal{X}_{p}\backslash\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\geq t}\mathcal{C}_{w,p}$.
+$\operatorname{Bs}\Big{(}H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})_{t,{\mathrm{deg}}(F_{Q})}\Big{)}=\mathcal{X}_{p}\backslash\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\geq t}\mathcal{C}_{w,p}$.
 Proof.
 Let $F_{T,\kappa(p)}$ be a reduction of $F_{Q,\kappa(p)}$ to $T$.
-For any $\nu\in X(T)$, consider the $\kappa(p)$-subspace $H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})[\nu]=F_{T,\kappa(p)}\times_{T}H^{0}(G/P,M_{\lambda})[\nu]$ of $H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})=F_{T,\kappa(p)}\times_{T}H^{0}(G/P,M_{\lambda})$.
+For any $\mu\in X(T)$, consider the $\kappa(p)$-subspace $H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})[\mu]=F_{T,\kappa(p)}\times_{T}H^{0}(G/P,M_{\lambda})[\mu]$ of $H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})=F_{T,\kappa(p)}\times_{T}H^{0}(G/P,M_{\lambda})$.
 For $w^{\prime}\in W/W_{P}$, set $C_{w^{\prime}}=F_{T,\kappa(p)}\times_{T}Bw^{\prime}P/P$ and $X_{w^{\prime}}=F_{T,\kappa(p)}\times_{T}\overline{Bw^{\prime}P}/P$.
@@
 On the other hand, we claim that if $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle<t$, then for any $s\in H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})_{t,{\mathrm{deg}}(F_{Q})}$, the restriction of $s$ to $X_{w}$ is zero.
-In fact, we may assume $s\in H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})[\nu]$ for some $\nu$ and $\langle{\mathrm{deg}}(F_{Q}),\nu\rangle\geq t$ implies that $\nu\not\leq w^{\prime}\lambda$ for any $w^{\prime}\in\pi^{-1}(w)$ (because if $\nu\leq w^{\prime}\lambda$ for some $w^{\prime}$, then $\langle{\mathrm{deg}}(F_{Q}),\nu\rangle\leq\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle<t$).
+In fact, we may assume $s\in H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})[\mu]$ for some $\mu$ and $\langle{\mathrm{deg}}(F_{Q}),\mu\rangle\geq t$ implies that $\mu\not\leq w^{\prime}\lambda$ for any $w^{\prime}\in\pi^{-1}(w)$ (because if $\mu\leq w^{\prime}\lambda$ for some $w^{\prime}$, then $\langle{\mathrm{deg}}(F_{Q}),\mu\rangle\leq\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle<t$).
 By Lemma 4.6(2), the restriction of $s$ to $X_{w^{\prime}}$ is zero for each $w^{\prime}$.
@@
 ∎
-Now we are ready to compute the augmented base locus of $\mathcal{L}_{\lambda}-tf$.
 Theorem 4.1.
-The augmented base locus of $\mathcal{L}_{\lambda}-tf$ is given by
+Let $B_{+}(\mathcal{L}_{\lambda}-tf)$ be the augmented base locus of $\mathcal{L}_{\lambda}-tf$. Then
 $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)=\mathcal{X}\Bigg{\backslash}\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle>t}\mathcal{C}_{w}=\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\leq t}\mathcal{C}_{w}.$
@@
 We may assume that $t\in\mathbb{N}$ by taking a sufficiently large tensor power of $\mathcal{L}_{\lambda}-tf$.
-Let $t_{0}=\min\{\langle{\mathrm{deg}}(F_{Q}),\nu\rangle:H^{0}(G/P,M_{\lambda})[\nu]\not=0\text{ and }\langle{\mathrm{deg}}(F_{Q}),\nu\rangle>t\}$.
+Let $t_{0}=\min\{\langle{\mathrm{deg}}(F_{Q}),\mu\rangle:H^{0}(G/P,M_{\lambda})[\mu]\not=0\text{ and }\langle{\mathrm{deg}}(F_{Q}),\mu\rangle>t\}$.
 Take $E=F_{Q}\times_{Q}H^{0}(G/P,M_{\lambda})_{t_{0},{\mathrm{deg}}(F_{Q})}\otimes\mathcal{O}_{C}(-tx)$, which is a subbundle of $\pi_{*}(\mathcal{L}_{\lambda}-tf).$ Moreover, $\mu_{\min}(E)=t_{0}-t>0$, i.e.
@@
 (4.1) $\mathrm{B}(\mathcal{L}_{\lambda}-tf)\subseteq\mathcal{X}\backslash\coprod\nolimits_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle>t}\mathcal{C}_{w}.$
-Now we take an ample line bundle $\mathcal{A}$ over $\mathcal{X}$ of form $\mathcal{L}_{\lambda^{\prime}}-t^{\prime}f$ where $\lambda^{\prime}\in X(T)$ is a strictly anti-dominant character and $\langle{\mathrm{deg}}(F_{Q}),\nu\rangle-t^{\prime}>0$ for any $\nu\in X(T)$ such that $H^{0}(G/P,M_{\lambda^{\prime}})[\nu]\not=0.$
+Now we take an ample line bundle $\mathcal{A}$ over $\mathcal{X}$ of form $\mathcal{L}_{\lambda^{\prime}}-t^{\prime}f$ where $\lambda^{\prime}\in X(T)$ is a strictly anti-dominant character and $\langle{\mathrm{deg}}(F_{Q}),\mu\rangle-t^{\prime}>0$ for any $\mu\in X(T)$ such that $H^{0}(G/P,M_{\lambda^{\prime}})[\mu]\not=0.$
 Take sufficiently large $n$ such that $\langle{\mathrm{deg}}(F_{Q}),w\lambda^{\prime}\rangle-t^{\prime}<n(t_{0}-t)$ for any $w\in W_{Q}\backslash W/W_{P}$.
@@
 $\limsup_{m\rightarrow+\infty}\frac{\mu_{\max}(\pi_{*}(m(\mathcal{L}_{\lambda}-tf)|_{\mathcal{X}_{w}})}{n}=\zeta_{1}(\mathcal{L}_{\lambda}-tf|_{\mathcal{X}_{w}})=\zeta_{w}-t\leq 0,$
-we have $\operatorname{Vol}(\mathcal{L}_{\lambda}-tf|_{\mathcal{X}_{w}})=0$ by a comparison result [12, Theorem 2.4] deduced from the Riemann-Roch theorem.
+we have $\operatorname{Vol}(\mathcal{L}_{\lambda}-tf|_{\mathcal{X}_{w}})=0$ by a comparison result [9, Theorem 2.4] deduced from the Riemann-Roch theorem.
 Therefore $\operatorname{Vol}(\mathcal{X}|\mathcal{X}_{w},\mathcal{L}_{\lambda}-tf)=0$, which implies that $\mathcal{X}_{w}\subseteq B_{+}(\mathcal{L}_{\lambda}-tf)$ due to Theorem 4.4.
 In conclusion, $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)=\mathcal{X}\backslash\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle>t}\mathcal{C}_{w}.$
-Remark 4.7.
-Comparing Theorem 2.11 and Theorem 4.6, we find that $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)_{K}$ agrees with the height filtration $Z_{t}$ for every $t\not=\zeta_{w}$.
-At critical points, $\mathrm{B}_{+}$ is upper continuous while $Z_{t}$ is lower continuous.
-4.3. Positive cones of flag bundles over curves
-We first give a reminder on the definitions of various positive cones of a projective variety.
-Let $Y$ be a smooth projective variety over a field $\mathbf{k}$ (of characteristic zero).
-We recall the following:
-  1. (1)
+4.3. Movable cones and proof of Theorem 4.2
+Let $Y$ be a projective smooth variety over $\mathbf{k}$. Recall
   • The big cone $\mathrm{Big}(Y)\subseteq N^{1}(Y)_{\mathbb{R}}$ is the cone of numerical classes of big divisors. Its closure is called Pseudo-effective cone, denoted as $\mathrm{Psef}(Y)$.
-  2. (2)
   • The ample cone $\mathrm{Amp}(Y)\subseteq N^{1}(Y)_{\mathbb{R}}$ is the cone of numerical classes of ample divisors. Its closure is called nef cone, denoted as $\mathrm{Nef}(Y)$.
-  3. (3)
-    For each $k=1,\cdots,\dim Y$, the $k$-th movable cone is defined as $\operatorname{Mov}^{k}(Y)=\big{\{}D\in N^{1}(Y)_{\mathbb{R}}:\mathrm{codim}(\mathrm{B}_{+}(D))\geq k\big{\}}$ which is open in $N^{1}(Y)_{\mathbb{R}}$.
-They are open cones in $N^{1}(Y)_{\mathbb{R}}$ and we have obvious inclusions $\operatorname{Mov}^{d}\subseteq\cdots\subseteq\operatorname{Mov}^{1}$.
-Note that $\operatorname{Mov}^{1}$ is the big cone and $\operatorname{Mov}^{d}$ is the ample cone.
-Now we come back to the case of flag varieties. Denote by
+  • For each $k=1,\cdots,\dim Y$, the $k$-th movable cone is defined as $\operatorname{Mov}^{k}(Y)=\big{\{}D\in N^{1}(Y)_{\mathbb{R}}:\mathrm{codim}(\mathrm{B}_{+}(D))\geq k\big{\}}$ which is a open in $N^{1}(Y)_{\mathbb{R}}$.
+We have obvious inclusions $\operatorname{Mov}^{d}(Y)\subseteq\cdots\subseteq\operatorname{Mov}^{1}(Y)$.
+Note that $\operatorname{Mov}^{1}(Y)$ is just the big cone and $\operatorname{Mov}^{d}(Y)$ is just the ample cone.
+Now we go back to the setting of §2. Denote by
 $c:X(P)\longrightarrow\mathrm{Pic}(G/P),\quad\lambda\longmapsto M_{\lambda}$
@@
 Since $G$ and $P$ are connected linear algebraic groups over a characteristic zero field, both $\mathrm{Pic}(G)$ and $\operatorname{Pic}(P)$ are finite.
-Notice that the numerical equivalence and the linear equivalence coincide on $G/P$ since $G/P$ is smooth fano.
+Notice that the numerical equivalence and the linear equivalence coincide on $G/P$ since it is smooth and Fano.
 Therefore the linear map $X(P)_{\mathbb{R}}\rightarrow N^{1}(G/P)_{\mathbb{R}}$ is surjective.
@@
 […]
-We now compute $\mathrm{Mov}^{i}(F/P)$.
+Then we compute $\mathrm{Mov}^{i}(F/P)$.
 Lemma 4.8.
@@
 ∎
-By the above description, we define two functions $\langle\alpha^{\vee},\cdot\rangle$ and $\langle{\mathrm{deg}}(F_{Q}),w\cdot\rangle$ on $N^{1}(F/P)_{\mathbb{R}}$ as
+Thanks to the above description, we may define two functions $\langle\alpha^{\vee},\cdot\rangle$ and $\langle{\mathrm{deg}}(F_{Q}),w\cdot\rangle$ on $N^{1}(F/P)_{\mathbb{R}}$ as
   1. (1)
@@
 ∎
-Corollary 4.10.
-$\mathcal{L}_{\lambda}-tf\in\operatorname{Mov}^{k}(F/P)$ normal-$\Longleftrightarrow$ $e_{k}(h_{\mathcal{L}_{\lambda}})>t$ where $e_{k}$ is the $k$-th minimum of Zhang.